### v1

# Degenerate symplectic fixed points and Gromov-Witten invariants

Abstract
In this paper we establish a connection between Gromov-Witten invariants and the number of degenerate fixed points of Hamiltonian diffeomorphisms on a closed rational symplectic manifold via deformed Hamiltonian spectral invariants. We prove a new cuplength estimate, in particular including Arnol’d conjecture over complex numbers, for fixed points of Hamiltonian diffeomorphisms on closed rational symplectic manifolds admitting nonzero Gromov-Witten invariants with two point insertions. We extend Schwarz’s quantum cuplength to the notion of deformed quantum cuplength for symplectic periods and employ it to estimate the number of fixed points of Hamiltonian diffeomorphisms on monotone symplectic manifolds with nonzero mixed Gromov-Witten invariants. We generalize Givental’s symplectic fixed point theorem for toric manifolds to closed rational symplectic manifolds admitting nonzero Gromov-Witten invariants with fixed marked points and one point insertion.
1 Introduction
The Arnol’d conjecture [2, 3] asserts that every Hamiltonian diffeomorphism $\varphi$ of a closed symplectic manifold $(M,\omega)$ possesses at least as many fixed points as a smooth function $f:M\to{\mathbb{R}}$ on $M$ possesses critical points. This conjecture has been known for a long time, and we refer to Hofer and Zehnder [34, Chapter 6] for its detailed history. The main motivation of this paper is the following weaker version, i.e., the homological Arnol’d conjecture:
#{fixed points $\#\{\hbox{fixed points of}\;\varphi\}\geq{\rm cuplength}(M;R).$
Here ${\rm cuplength}(M;R)$ denotes the cuplength of $M$ over a coefficient ring $R$ which is defined as the maximal number $k+1\in{\mathbb{N}}$ such that there exist $k$ cohomology classes $a_{i}\in H^{*}(M;R)$ with $\deg(a_{i})>0$ satisfying $a_{1}\cup\cdots\cup a_{k}\neq 0$. If all the fixed points of $\varphi$ are a priori nondegenerate, then the nondegenerate homological Arnol’d conjecture asserts that the number of the fixed points of $\varphi$ is not less than the Betti sum of $M$ over any coefficient ring. Since the advent of Floer homology [15, 16, 17], there has been a huge progress in the nondegenerate version, for instance, in the 1990’s Fukaya-Ono [20], Liu-Tian [41] and Ruan [59] independently settled Arnol’d conjecture over rational numbers, and recently Abouzaid-Blumberg [1], and Bai-Xu [5] and Rezchikov [57] extended their results to the coefficients over any finite field and intergers, respectively. For earlier results on the nondegenerate Arnol’d conjecture, we refer to [11, 17, 32, 53], etc. It is worth to mention that Pardon [54] and Filippenko-Wehrheim [14] employed different versions of the virtual technique to reprove Arnol’d conjecture over rational numbers.
However, compared to the nondegenerate case, progress on the degenerate homological Arnold conjecture has not kept pace and, in fact, remains far from being solved except for some special cases, see for instance a standard torus $(\mathbb{T}^{2n},\omega_{0})$ by Conley and Zehnder [11], symplectic manifolds $(M,\omega)$ with $\pi_{2}(M)=0$ by Floer [18] and Hofer [31], $(\mathbb{C}P^{n},\omega_{\rm FS})$ by Fortune [19] and negatively monotone symplectic manifolds $(M^{2n},\omega)$ with minimal Chern number at least $n$ by Le and Ono [40], etc. As far as the author knows, the projective space ${\mathbb{C}}P^{n}$ with the Fubini-Study form $\omega_{\rm FS}$ is the only monotone symplectic manifold, i.e., $\omega|_{\pi_{2}(M)}=\lambda c_{1}|_{\pi_{2}(M)}\neq 0$ with $\lambda>0$, for which the degenerate homological Arnol’d conjecture has been proven until now.
As the title of [17] indicated, Floer was aware of the significance of $J$-homolophic spheres toward the degenerate Arnol’d conjecture more than 35 years ago. At first glance, the presence of those $J$-homolophic spheres seems to obstruct the occurrence of symplectic fixed points of Hamiltonian diffeomorphisms on symplectic manifolds. Meanwhile, it was commented by McDuff and Salamon [48, Remark 9.4.12] that “The existence of a nontrivial Gromov–Witten invariant is more significant, though as yet the geometric and dynamical consequences of this condition are far from being fully understood. So far, most applications have come via properties of the small quantum cohomology ring …” Our first theorem gives a concrete connection between the number of fixed points of Hamiltonian diffeomorphisms and nonzero Gromov-Witten invariants with two point insertions on a rational symplectic manifold $(M,\omega)$, i.e., $\omega(\pi_{2}(M))=p\cdot{\mathbb{Z}}$ for some positive number $p$, called the minimal symplectic period of $(M,\omega)$.
Theorem 1.1.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$, which admits a nonzero Gromov-Witten invariant of the form:
${\rm GW}_{0,k+3,A}\big{(}[pt],a_{0},[pt],a_{1}\ldots,a_{k}\big{)}$ (1.1)
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, and $a_{0},\ldots,a_{k}$ are rational homology classes of even degree. Then, for any $\varphi\in\operatorname{Ham}(M,\omega)$,
#$\#\operatorname{Fix}(\varphi)\geq\bigg{\lceil}\frac{p\cdot\operatorname{cuplength}(M;{\mathbb{C}})}{\Theta(M,\omega)}\bigg{\rceil}$
[…]
Remark 1.1.
Clearly, we have $\Theta\geq p$ in the above theorem because $\operatorname{GW}_{0,k+3,A}\neq 0$ with two point insertions implies that the homology class $A$ can be represented by a $J$-holomorphic stable map of genus zero which has a positive symplectic area. Note also that if $\Theta(M,\omega)<\infty$, then the infimum in the definition of $\Theta(M,\omega)$ can be attained by the Gromov compactness theorem (cf. [48]). The question of whether there are Gromov-invariants with two point insertions is related to the notion of rationally connectedness in algebraic geometry. Voisin [70] and Tian [64, 65] studied some important examples including rationally connected $3$-folds and $4$-folds on this kind of question. Actually, a folklore conjecture asserts that every rationally connected projective manifold carries a nonzero genus zero Gromov-Witten invariant with two point insertions, see for instance [37, Conjecture 3.4].
Remark 1.2.
Gromov-Witten invariants have some important applications in contact/ symplectic dynamics such as: one was given by Liu-Tian [42] and Taubes [63] to prove Weinstein conjecture by using the nonvanishing property of these invariants (prior to these two results, Hofer and Viterbo [33] and Chen [10] had been aware of the important role of the appearance of $J$-holomorphic spheres toward Weinstein conjecture); another one was first discovered by Lu [44] and latter reproved by Usher [68] for bounding Hofer-Zehnder capacity from above for a large class of symplectic manifolds, etc.
Remark 1.3.
Under the same assumptions in Theorem 1.1, one can prove that there exists an open dense set of deformation parameter $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that the $\zeta$-spectral norm $\gamma_{\zeta}$ (cf. Definition 3.5) satisfies $\gamma_{\zeta}(\varphi)<\langle[\omega],A\rangle$ for all $\varphi\in\operatorname{Ham}(M,\omega)$ with finite #$\#{\rm Fix}(\varphi)$, and hence $\gamma_{\zeta}$ is bounded on $\operatorname{Ham}(M,\omega)$ by $\langle[\omega],A\rangle$, see Remark 4.3. Here a very interesting question is whether this strict inequality holds for all Hamiltonian diffeomorphisms on $(M,\omega)$.
The following theorem is a straightforward generalization of [28, Theorem 8].
Theorem 1.2.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$. If there exists a class $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M,\Lambda_{0})$ such that the $\zeta$-spectral norm of $\varphi\in\operatorname{Ham}(M,\omega)$ satisfies $\gamma_{\zeta}(\varphi)<p$, then the number of fixed points of $\varphi$ has a lower bound given by $\operatorname{cuplength}(M;{\mathbb{C}})$. Hence, if the Hofer norm of $\varphi$ satisfies $\|\varphi\|<p$, then, by  (3.1), Arnol’d conjecture over complex numbers holds for $\varphi$.
Remark 1.4.
Prior to Theorem 1.2, Schwarz has proved that Arnol’d conjecture over ${\mathbb{Z}}_{2}$ holds for $\|\varphi\|<p$, see [61, Theorem 1.1].
Through introducing the $\zeta$-deformed quantum cuplength ${{\rm qcl}_{\zeta}}$ for symplectic periods (cf. Definition 2.7), our approach developed in this paper gives the following extension of Schwarz’s estimate [61, Theorem 1.2] for degenerate symplectic fixed points to all closed rational symplectic manifolds.
Theorem 1.3.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with period $p>0$, $\omega(\pi_{2}(M))=p\cdot{\mathbb{Z}}$. Then, for each $\varphi\in\operatorname{Ham}(M,\omega)$, we have the estimate
#$\#{\rm Fix}(\varphi)\geq\sup\limits_{\zeta\in H_{*}(M;\Lambda_{0})}\sup\limits_{g\in\Gamma_{\omega}}\bigg{\lceil}\frac{p\cdot{{\rm qcl}_{\zeta}}(g)}{p+g+\|\varphi\|}\bigg{\rceil}.$
To the best of the author’s knowledge, for closed monotone symplectic manifold, Schwarz [61] has extended all previously known estimates for degenerate symplectic fixed points via his approach of quantum cuplength estimate. Our next theorem reveals that his result [61, Corollary 1.3] can be reinterpreted in terms of nonzero mixed Gromov-Witten invariants (cf. Ruan and Tian [58, Section 2]).
Theorem 1.4.
Let $(M^{2n},\omega)$ be a closed monotone symplectic manifold with minimal Chern number $N\geq 1$, and let $a\in H_{*<2n}(M;{\mathbb{Q}})$. Suppose that for each $2n<l\in{\mathbb{N}}$, there exists a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A_{l}}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b_{0}^{l},\ldots,b_{k}^{l})$ (1.2)
for some integer $k=k(l)\in{\mathbb{N}}\cup\{0\}$, where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, and $b_{0}^{l},\ldots,b_{k}^{l}$ are rational homology classes of even degree. Then, for any $\varphi\in\operatorname{Ham}(M,\omega)$,
#$\#{\rm Fix}(\varphi)\geq\frac{2N}{2n-\deg(a)}$
provided that the function $l\mapsto k(l)$ is bounded.
For $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M,\Lambda_{0})$, we call a class $a\in H_{*<2n}(M;{\mathbb{C}})$ a $\zeta$-nonnilpotent element in the $\zeta$-deformed quatum homology ring $QH_{*}(M,\zeta)_{\zeta}$ if $\underbrace{a*_{\zeta}\cdots*_{\zeta}a}_{l}\neq 0$ for all $l\in{\mathbb{N}}$. Note that the existence of a $\zeta$-nonnilpotent element implies that for each $2n\leq l\in{\mathbb{N}}$, one has a nonzero mixed Gromov-Witten invariant as in (1.2). Furthermore, if $\zeta=0$, the function $k(l)$ as in Theorem 1.4 can be taken to be identically zero. Hence, we obtain
Corollary 1.5 ([61, Corollary 1.3]).
Let $(M,\omega)$ be a monotone symplectic manifold with minimal Chern number $N$ determined by $c_{1}(\pi_{2}(M))=N\cdot{\mathbb{Z}}$. Suppose that $a\in H_{k}(M;{\mathbb{C}})$ ($0\leq k<2n$) is a 00-nonnilpotent element in $QH_{*}(M,\zeta)_{0}$. Then, for all $\varphi\in\operatorname{Ham}(M,\omega)$,
#$\#{\rm Fix}(\varphi)\geq\frac{2N}{2n-k}.$
Compared to Theorem 1.1, we have following theorem involved in nonzero mixed Gromov-Witten invariants with one point insertion in favor of the author’s belief that the existence of nontrivial Gromov–Witten invariants matters much for counting symplectic fixed points of Hamiltonian diffeomorphisms on a large class of sympectic manifolds.
Theorem 1.5.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],b_{1},\ldots,b_{k})$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{Q}})$ with $2\leq l\in{\mathbb{N}}$, and the classes $b_{1},\ldots,b_{k}$ are rational homology classes of even degree. Then, for each $\varphi\in\operatorname{Ham}(M,\omega)$,
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{pl}{\langle[\omega],A\rangle}\bigg{\rceil}.$ (1.3)
Assume in addition that for some $2\leq l\in{\mathbb{N}}$, $(M^{2n},\omega)$ has a nonzero Gromov-Witten invariant with fixed marked points of the form
${\rm GW}_{0,l+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt])$
with $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$ and $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{C}})$, then
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{2p}{\max_{i}\{2n-a_{i}\}}\cdot\frac{\langle c_{1}(TM),A\rangle}{\langle[\omega],A\rangle}\bigg{\rceil}$ (1.4)
for all $\varphi\in\operatorname{Ham}(M,\omega)$.
Inspired by Schwarz’s quantum cuplength estimate, a different way for estimating degenerate symplectic fixed points was obtained by the author via introducing the so-called fundamental quantum factorization, see [28, Definition 23]. Using the big quantum homolgy ring, one can extend this notion as follows.
Definition 1.6.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$. We say that $M$ has a principal fundamental quantum factorization associated to $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M,\Lambda_{0})$ (denoted by PFQF for short) of length $l\in{\mathbb{N}}$ with order $g\in{\mathbb{N}}$ if there exist $u_{1},\ldots,u_{l}\in H_{*<2n}(M;{\mathbb{C}})$ such that
$zT^{-gp}[M]+\beta=u_{1}*_{\zeta}\cdots*_{\zeta}u_{l}$
where $0\neq z\in{\mathbb{C}}$ and $\beta\in H_{*<2n}(M;\Lambda)$.
Since the existence of PFQF yields a nonzero mixed Gromov-Witten invariant with one point insertion, the first assertion of Theorem 1.5 immediately leads to
Corollary 1.7 ([28, Theorem 9]).
Let $(M,\omega)$ be a closed rational symplectic manifold admitting a PFQF of length $l$ with order $g$. Then each $\varphi\in\operatorname{Ham}(M,\omega)$ has at least $\lceil\frac{l}{g}\rceil$ fixed points.
1.1 Applications and examples
1.1.1 Symplectic manifolds with $k$-points blow-up
The process of blowing up in complex geometry is well-known. Blowing up replaces a point $x$ in a complex manifold $X$ by the set $\Sigma$ of all lines through this point, which is called an exceptional divisor biholomorphic to ${\mathbb{C}}P^{n-1}$. In the symplectic setting, the role of a point is played by a symplectically embedded standard ball, and blowing up amounts to removing the interior of a symplectic ball and collapsing the bounding sphere via the Hopf map to the exceptional divisor.
Let $(M,\omega)$ be a closed symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant of the form (1.1). Recall that the Gromov width $c_{G}$ of a symplectic manifold $(M^{2n},\omega)$ is defined as
$c_{G}:=\sup\{\pi r^{2}|\exists\;\hbox{a symplectic embedding}\;B^{2n}(r)\hookrightarrow M\}$
where $B^{2n}(r)$ stands for the open ball of radius $r>0$ with center at the origin in ${\mathbb{R}}^{2n}$. It is well known that $c_{G}\leq\Theta(M,\omega)$, see [29, 44, 45].
Given $k,m\in{\mathbb{N}}$ with $\frac{kp}{m}<c_{G}$, we now take a symplectic embedding $\psi:B^{2n}(\lambda)\hookrightarrow M$ such that $\pi\lambda^{2}=\frac{kp}{m}$ and a positive number $\epsilon\in{\mathbb{R}}$ such that $\pi(\lambda+\epsilon)^{2}<c_{G}$. Associated to $\psi$, we blow up $(M,\omega)$ at some point $q_{0}\in M$ of weight $\lambda$, and denote by $(\widetilde{M},\widetilde{\omega}_{\psi,\lambda,\epsilon})$ the resulting symplectic manifold. For the details about this symplectic surgery, we refer the readers to McDuff-Salamon [47, Section 7.1] and the references therein. We abbreviate $\widetilde{\omega}=\widetilde{\omega}_{\psi,\lambda,\epsilon}$. We denote by $Z\subset\widetilde{M}$ the exceptional divisor which can be identified with ${\mathbb{C}}P^{n-1}$, and $\pi_{M}:\widetilde{M}\to M$ the projection map given by $\pi_{M}|_{M\setminus\{q_{0}\}}=id$ and $\pi_{M}(\widetilde{q})=q_{0}$ for all $\widetilde{q}\in Z$. For simplicity, for any cohomology classes $a\in H^{*}(M;{\mathbb{Q}})$, $b\in H^{*}(M;{\mathbb{Z}})$ we denote the pulling back classes $a\in H^{*}(\widetilde{M};{\mathbb{Q}})$, $b\in H^{*}(\widetilde{M};{\mathbb{Z}})$ as
$\widetilde{a}=\pi_{M}^{*}a,\quad\widetilde{b}=\pi_{M}^{*}b.$
Let $e$ denote the Poincaré dual of the divisor $Z$ in $H^{2}(\widetilde{M};{\mathbb{Z}})$, i.e., $e={PD}_{\widetilde{M}}[Z]$. It is well known that
$[\widetilde{\omega}]=\widetilde{[\omega]}-\pi\lambda^{2}e,\quad c_{1}(T\widetilde{M})=\widetilde{c_{1}(TM)}-(n-1)e$ (1.5)
(see page 311 in [47]). From (1.5) and the fact that $\widetilde{M}$ is diffeomorphic to the oriented connected sum $M\sharp\overline{{\mathbb{C}}P^{n}}$ (here $\overline{{\mathbb{C}}P^{n}}$ is the manifold ${\mathbb{C}}P^{n}$ with the opposite orientation), it is not hard to see that the symplectic period of $(\widetilde{M},\widetilde{\omega})$ is $p\cdot{\rm gcd}(m,k)/{m}$, and that $(\widetilde{M},\widetilde{\omega})$ is monotone if $(M,\omega)$ is monotone with $\omega|_{\pi_{2}(M)}=\kappa c_{1}(TM)|_{\pi_{2}(M)}$ and $\pi\lambda^{2}=\kappa(n-1)$.
According to the existence of a nonzero Gromov-Witten invariant of the form (1.1), one may assume that $A\in H_{2}(M;{\mathbb{Z}})/{\rm torsion}$ is a class which realizes the quantity $\Theta(M,\omega)$, see Remark 1.1. Hence, we get a nonzero Gromov-Witten invariant of the form:
${\rm GW}^{M}_{0,k+3,A}\big{(}[pt],a_{0},[pt],a_{1}\ldots,a_{k}\big{)}$
with $\Theta(M,\omega)=\langle[\omega],A\rangle$, where $a_{0},\ldots,a_{k}$ are rational homology classes of even degree. Hu [36] obtained an explicit relation of Gromov-Witten invariants between a symplectic manifold and its one point blow-up. In particular, using [36, Theorem 1.2], one can see that
${\rm GW}^{\widetilde{M}}_{0,k+3,\widetilde{A}}\big{(}\widetilde{[pt]}_{M},\widetilde{a}_{0},\widetilde{[pt]}_{M},\widetilde{a}_{1}\ldots,\widetilde{a}_{k}\big{)}={\rm GW}^{M}_{0,k+3,A}\big{(}[pt]_{M},a_{0},[pt]_{M},a_{1}\ldots,a_{k}\big{)}\neq 0,$
where $[pt]_{M}$ is the point class in $M$, $\widetilde{A}={\rm PD}_{\widetilde{M}}\widetilde{{\rm PD}_{M}(A)}$ and $\widetilde{a}_{i}={\rm PD}_{\widetilde{M}}\widetilde{{\rm PD}_{M}(a_{i})}$ for all $i\in\{1,\ldots,k\}$. Since $\widetilde{[pt]}_{M}$ is indeed the point class $[pt]_{\widetilde{M}}$ on $\widetilde{M}$, the nonzero Gromov-Witten invariant yields
$\Theta(\widetilde{M},\widetilde{\omega})\leq\langle[\widetilde{\omega}],\widetilde{A}\rangle=\langle[\omega],A\rangle.$
This, together with Theorem 1.1, leads to
Theorem 1.6.
Let $(M,\omega)$ be a closed symplectic manifold with symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant of the form (1.1). Assume that $(\widetilde{M},\widetilde{\omega})$ is the symplectic blow-up of $(M,\omega)$ of weight $\lambda$ with the property that $\pi\lambda^{2}=kp/m<c_{G}$ for some $k,m\in{\mathbb{N}}$. Then
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{p\cdot{\rm gcd}(m,k)\cdot\operatorname{cuplength}(\widetilde{M};{\mathbb{C}})}{m\cdot\Theta(M,\omega)}\bigg{\rceil}$
for all $\varphi\in\operatorname{Ham}(\widetilde{M},\widetilde{\omega})$.
Remark 1.8.
In case that $\Theta(M,\omega)=p$, the above inequality reduces to
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{{\rm gcd}(m,k)\cdot\operatorname{cuplength}(\widetilde{M};{\mathbb{C}})}{m}\bigg{\rceil}$
for all $\varphi\in\operatorname{Ham}(\widetilde{M},\widetilde{\omega})$. In particular, for $({\mathbb{C}}P^{n},\omega_{\rm FS})$, one can choose $k=1,m=2$, and hence gets
#$\#{\rm Fix}(\varphi)\geq\big{\lceil}\operatorname{cuplength}(\widetilde{{\mathbb{C}}P^{n}};{\mathbb{C}})/2\big{\rceil}=\lceil(n+1)/2\rceil$
for any $\varphi\in\operatorname{Ham}(\widetilde{{\mathbb{C}}P^{n}},\widetilde{\omega}_{\rm FS})$. We remark here that the blow-up $(\widetilde{{\mathbb{C}}P^{n}},\widetilde{\omega}_{\rm FS})$ in this case is not monotone except for $n=3$ since, otherwise, $\frac{\kappa(n+1)}{2}=\frac{p}{2}=\pi\lambda^{2}=\kappa(n-1)$ yields $n=3$.
Following McDuff-Polterovich [46], one can blow up an almost Kähler manifold at finite many points, and this process of blowing up is closely related to the symplectic packing problem. Let $(M,\omega,J)$ be a symplectic manifold such that $\omega$ tames $J$. Assume that $J$ is integrable near the points $x_{1},\ldots,x_{k}\in M$ and let $(\widetilde{M},\widetilde{J})$ be its complex blow-up at these points. Then there is a holomorphic map $\pi_{M}:(\widetilde{M},\widetilde{J})\to(M,J)$ which is a bijection over $M\setminus\{x_{1},\ldots,x_{k}\}$ and maps the exceptional divisor to the points $x_{j}$, $j\in\{1,\ldots,k\}$. As before, we denote by $e_{1},\ldots,e_{k}$ the cohomology classes on $\widetilde{M}$ which are Poincaré dual to the exceptional divisors. If
$\psi=\coprod_{j=1}^{k}\psi_{j}:\coprod_{j=1}^{k}(B(\lambda_{j}+\varepsilon),\omega,i)\longrightarrow(M,\omega,J)$
is a symplectic and holomorphic embedding of $k$ disjoint balls of radii $\lambda_{j}+\varepsilon$ into $M$ whose centers are mapped to the points $x_{1},\ldots,x_{k}$, then by [46, Proposition 2.1B] one can construct a symplectic form $\widetilde{\omega}$ on $\widetilde{M}$ which is tamed by $\widetilde{J}$ and satisfies
$[\widetilde{\omega}]=[\pi_{M}^{*}\omega]-\sum_{j=1}^{k}\pi\lambda_{j}^{2}e_{j}.$
Additionally, we assume that $(M,\omega)$ is a closed symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant of the form (1.1). If one can choose $m,n_{1},\ldots,n_{k}\in{\mathbb{N}}$ such that $\pi\lambda_{j}^{2}=pn_{j}/m$, then the minimal symplectic period of $(\widetilde{M},\widetilde{\omega})$ is $p\cdot{\rm gcd}(m,n_{1},\ldots,n_{k})/m$. Using [36, Theorem 1.2] repeatedly and arguing as before, one can see that $(\widetilde{M},\widetilde{\omega})$ admits a nonzero Gromov-Witten invariant of the form (1.1) with $\langle[\widetilde{\omega}],\widetilde{A}\rangle\leq\Theta(M,\omega)$. Hence, by Theorem 1.1, we obtain
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{p\cdot{\rm gcd}(m,n_{1},\ldots,n_{k})}{m\cdot\Theta(M,\omega)}\cdot\operatorname{cuplength}(\widetilde{M};{\mathbb{C}})\bigg{\rceil}$
for all $\varphi\in\operatorname{Ham}(\widetilde{M},\widetilde{\omega})$.
1.1.2 Coadjoint orbits of $U(n)$
We shall give symplectic fixed point estimates for the monotone coadjoint orbits of the unitary group $U(n)$ for which its Lie algebra is of type $A_{n-1}$ in the theory abstract root systems. For general monotone coadjoint orbits of compact simple Lie groups, one can obtain similar lower bound estimates for symplectic fixed points, see Theorem A.1 in the appendix. Our approach is based on Caviedes Castro’s computation [8] of the upper bound for the quantity $\Theta(M,\omega)$.
The Lie algebra of unitary group $U(n)$ is the space $\mathfrak{u}(n)$ of all $n\times n$ skew-Hermitian matrices. Take the maximal torus $T=U(1)^{n}$ of diagonal matrices in $U(n)$. Its Lie algebra can be identified as
${\mathfrak{t}}=\{(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}\}.$
Let $\{e_{1},\ldots,e_{n}\}$ be the standard basis of ${\mathbb{R}}^{n}$. We choose the following adjoint invariant metric on $\mathfrak{u}(n)$
$\mathfrak{u}(n)\times\mathfrak{u}(n)\longrightarrow{\mathbb{R}},\quad(X,Y)\longmapsto-{\rm Trace}(X,Y).$
It turns out that the roots are of the form
$\alpha_{i,j}:{\mathfrak{t}}\longrightarrow{\mathbb{R}},\quad(x_{1},\ldots,x_{n})\longmapsto x_{i}-x_{j}$
where $1\leq i\neq j\leq n$. Hence, under the identification ${\mathfrak{t}}\cong{\mathfrak{t}}^{*}$, we have $\alpha_{i,j}=e_{i}-e_{j}$, and the coroots are
$\check{\alpha}_{i,j}=\frac{2(e_{i}-e_{j})}{\langle e_{i}-e_{j},e_{i}-e_{j}\rangle}=e_{i}-e_{j}.$
The Weyl group is the symmetric group $S_{n}$ generated by the reflections
$s_{\alpha_{i,j}}:{\mathbb{R}}^{n}\longrightarrow{\mathbb{R}}^{n},\quad(x_{1},\ldots,x_{i},\ldots,x_{j},\ldots,x_{n})\mapsto(x_{1},\ldots,x_{j},\ldots,x_{i},\ldots,x_{n}).$
The longest element $w_{0}$ in $S_{n}$ relative to the set of simple roots $\{\alpha_{i,i+1}|1\leq i\leq n-1\}$ is the permutation
$w_{0}:{\mathbb{R}}^{n}\longrightarrow{\mathbb{R}}^{n},\quad(x_{1}\ldots,x_{n})\to(x_{n},\ldots,x_{1}).$
Now we consider the coadjoint orbit ${\mathcal{O}}_{\lambda}$ of $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in{\mathbb{R}}^{n}\setminus\{0\}$ with $\lambda_{1}\geq,\ldots,\geq\lambda_{n}$, i.e.,
${\mathcal{O}}_{\lambda}:=\big{\{}A\in\mathfrak{u}(n)|A^{*}=-A,\;\hbox{eigenvalues of}\;A\;\hbox{are}\;\sqrt{-1}\lambda_{1},\ldots,\sqrt{-1}\lambda_{n}\big{\}}$
which is endowed with the Kostant-Kirillov-Souriau form $\omega_{\lambda}$. We choose $\lambda$ as above such that $(\omega_{\lambda},\omega_{\lambda})$ is monotone. Note that
$w_{0}=s_{\alpha_{1,n}}s_{\alpha_{2,n-1}}\cdots s_{\alpha_{[n/2],n-[n/2]+1}}.$
Combining this with Theorem A.1, one finds that
$\Theta(M,\omega)\leq\sum_{k=1}^{[n/2]}\langle\lambda,\check{\alpha}_{k,n-k+1}\rangle=\sum_{k=1}^{[n/2]}\lambda_{k}-\lambda_{n-k+1}.$
Note also that the symplectic period of $({\mathcal{O}}_{\lambda},\omega_{\lambda})$ is
$p=\min_{1\leq i<j\leq n}\{\ell_{ij}\langle\lambda,\check{\alpha}_{i,j}\rangle>0|\ell_{ij}\in{\mathbb{Z}}\}=\min_{2\leq i\leq n}\{\ell_{i}(\lambda_{i-1}-\lambda_{i})>0|\ell_{i}\in{\mathbb{Z}}\}.$
Consequently, by Theorem 1.1, for any $\varphi\in\operatorname{Ham}({\mathcal{O}}_{\lambda},\omega_{\lambda})$, we have
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{\min_{2\leq i\leq n}\{\ell_{i}(\lambda_{i-1}-\lambda_{i})>0|\ell_{i}\in{\mathbb{Z}}\}\cdot{\rm cuplength}(M;{\mathbb{C}})}{\sum_{k=1}^{[n/2]}\lambda_{k}-\lambda_{n-k+1}}\bigg{\rceil}.$ (1.6)
Remark 1.9.
According to the theory of abstract root systems, there are only finite many types of root systems: type $A_{n-1}$, type $B_{n}$, type $C_{n}$, type $D_{n}$, types $E_{6}$, $E_{7}$, $E_{8}$, type $F_{4}$ and type $G_{2}$, see for instance [24]. These types of abstract root systems precisely correspond to the Dynkin diagrams. The symplectic fixed point estimate for remaining types can be done similarly, we omit them and refer the readers to [8, Section 9] and Appendix A.
1.1.3 Monotone partial flag manifolds
Given a sequence $0=n_{0}<n_{1}<,\ldots,<n_{r}<n_{r+1}=n$ of integers, the partial flag manifold $F=F(n_{1},\ldots,n_{r},n)$ is a complex manifold consisting of all sequences
$0\subset V_{1}\subset\ldots\subset V_{r}\subset{\mathbb{C}}^{n}$
where each $V_{i}$ is a complex linear space with $\dim_{\mathbb{C}}(V_{i})=n_{i}$. The complex dimension of $F(n_{1},\ldots,n_{r},n)$ is $\sum_{i=1}^{r}(n_{i}-n_{i-1})(n_{i+1}-n_{i})$. Let $F_{n}$ denote the complete flag manifold $F(1,2,\ldots,n)$ for short. Let $\{e_{1},\ldots,e_{n}\}$ be the standard basis of ${\mathbb{C}}^{n}$. Denote by $P=P(n_{1},\ldots,n_{r},n)\subset GL_{n}({\mathbb{C}})$ the isotropic subgroup of the standard flag
$V_{i}={\rm span}\{e_{1},\ldots,e_{n_{i}}\}.$
Since the intersection of $P$ and $U(n)$ is $U(k_{1})\times\ldots\times U(k_{r+1})$ with $k_{i}=n_{i}-n_{i-1}$ for $i=1,\ldots,r+1$, the partial flag manifold reads as
$F(n_{1},\ldots,n_{r},n)=GL_{n}({\mathbb{C}})/P=U(n)/U(k_{1})\times\ldots\times U(k_{r+1}).$
In particular, the complete flag manifold can be written as $F_{n}=GL_{n}({\mathbb{C}})/B=U(n)/T$, where $B\subset GL_{n}({\mathbb{C}})$ is a Borel subgroup consisting of upper triangular invertible matrices, and $T$ is a maximal torus in $U(n)$ consisting of diagonal matrices.
Next we give a coadjoint orbit description of flag manifolds. As before, we identify the Lie algebra $\mathfrak{u}(n)$ of the unitary group $U(n)$ with $\sqrt{-1}\mathfrak{u}(n)^{*}$ via a $U(n)$-invariant scalar product. Fix $\lambda={\rm diag}(\lambda_{1},\ldots,\lambda_{n})\in-\sqrt{-1}\mathfrak{u}(n)$ with
$\underbrace{\lambda_{1}=\ldots=\lambda_{n_{1}}}_{k_{1}}>\underbrace{\lambda_{n_{1}+1}=\ldots=\lambda_{n_{2}}}_{k_{2}}>\ldots>\underbrace{\lambda_{n_{r}+1}=\ldots=\lambda_{n}}_{k_{r+1}}.$
Then one can identify the coadjoint orbit ${\mathcal{O}}_{\lambda}\subset-\sqrt{-1}\mathfrak{u}(n)$ of $\lambda$ by
$F=U(n)/U(k_{1})\times\ldots\times U(k_{r+1})\longrightarrow{\mathcal{O}}_{\lambda},\quad[A]\longmapsto A\lambda A^{*}.$
This gives $F$ a Kähler form $\omega$ via pulling back by the Kostant-Kirillov-Souriau form $\omega_{\lambda}$ on ${\mathcal{O}}_{\lambda}$. It is shown in [49] that $c_{1}(T{\mathcal{O}}_{\lambda})=[\omega_{\lambda}]$ if and only if
$\lambda=(\underbrace{n-n_{1},\ldots}_{k_{1}},\underbrace{n-n_{1}-n_{2},\ldots}_{k_{2}},\ldots,\underbrace{n-n_{r-1}-n_{r},\ldots}_{k_{r}},\underbrace{-n_{r},\ldots,-n_{r}}_{k_{r+1}})+(\underbrace{c,\ldots,c}_{n})$
for some $c\in{\mathbb{R}}$.
Now we give symplectic fixed point estimates for monotone flag manifolds in the following cases:
  • 1.
    The complex projective space ${\mathbb{C}}P^{n-1}$longleftrightarrow $\lambda=(n-1,\underbrace{-1,\ldots,-1}_{n-1})$.
  • 2.
    The Grassmannian $Gr_{\mathbb{C}}(k,n)$longleftrightarrow $\lambda=(\underbrace{n-k,\ldots,n-k}_{k},\underbrace{-k,\ldots,-k}_{n-k})$.
  • 3.
    The partial flag manifold $F(1,n-1,n)$longleftrightarrow $\lambda=(n-1,\underbrace{0,\ldots,0}_{n-2},-n+1)$.
  • 4.
    The partial flag manifold $U(n)/U(k)^{m}$ with $n=mk$ corresponds to
    $\small\lambda=(\underbrace{(m-1)k,\ldots,(m-1)k}_{k},\underbrace{(m-3)k,\ldots,(m-3)k}_{k},\ldots,\underbrace{-(m-1)k\ldots,-(m-1)k}_{k}).$
  • 5.
    The complete flag manifold $F_{n}$longleftrightarrow $\lambda=(\underbrace{n-1,n-3,\ldots,-n+1}_{n})$.
Since ${\operatorname{cuplength}}(M;{\mathbb{C}})={\dim}_{\mathbb{C}}(M)+1$ for any simply connected algebraic manifold $M$ (see [6]), we obtain
${\rm cuplength}(F(n_{1},\ldots,n_{r},n);{\mathbb{C}})=\sum_{i=1}^{r}(n_{i}-n_{i-1})(n_{i+1}-n_{i})+1.$
Using (1.6), we finds the following:
Case 1. For any $\varphi\in{\rm Ham}({\mathbb{C}}P^{n-1},\omega)$ with $n\geq 2$, #$\#{\rm Fix}(\varphi)\geq n$.
Case 2. For any $\varphi\in{\rm Ham}(Gr_{\mathbb{C}}(k,n),\omega)$, we have
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{n\cdot(k(n-k)+1)}{n\cdot\min\{k,n-k\}}\bigg{\rceil}=\max\{k,n-k\}+1.$
Case 3. For any $\varphi\in{\rm Ham}(F(1,n-1,n),\omega)$ with $n\geq 3$,
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{(n-1)\cdot 2(n-2)}{2(n-1)}\bigg{\rceil}=n-2.$
Case 4. For any $\varphi\in{\rm Ham}(U(n)/U(k)^{m},\omega)$ with $n=mk$ and $m\geq 2$,
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{2k\cdot(k^{2}m(m-1)/2+1)}{2k(m-\lfloor m/2\rfloor)\cdot\lfloor m/2\rfloor}\bigg{\rceil}=2k^{2}-\bigg{\lfloor}\frac{k^{2}\lfloor m/2\rfloor-1}{\lfloor m/2\rfloor(m-\lfloor m/2\rfloor)}\bigg{\rfloor}.$
Case 5. For any $\varphi\in{\rm Ham}(F_{n},\omega)$,
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{2\cdot(n(n-1)/2+1)}{2(n-\lfloor n/2\rfloor)\cdot\lfloor n/2\rfloor}\bigg{\rceil}=2.$
Remark 1.10.
Note that the complex Grassmanian $Gr_{\mathbb{C}}(k,n)$ with minimal Chern number $n$ has a 00-nonnilpotent element of degree $\dim_{\mathbb{R}}(Gr_{\mathbb{C}}(k,n))-2$ given by the Poincaré dual of $c_{1}(TGr_{\mathbb{C}}(k,n))$ , see [61]. By Corollary 1.5, we have #$\#{\rm Fix}(\varphi)\geq n$ for all $\varphi\in{\rm Ham}(Gr_{\mathbb{C}}(k,n),\omega)$. This is a better lower bound estimate than that given in case 2. Similarly, for the complete flag manifold $F_{n}$ with minimal Chern number $2$, there is a 00-nonnilpotent element of degree $\dim_{\mathbb{R}}(F_{n})-2$, and hence we obtain the same estimate as in case 5, see also [61], which seems to be far from the expected result.
1.1.4 Fano toric manifolds
Let $(M^{2m},\omega)$ be a toric manifold, namely a symplectic quotient ${\mathbb{C}}^{n}/\mkern-5.0mu/{\mathbb{T}}^{k}$ of ${\mathbb{C}}^{n}$ by a subtorus ${\mathbb{T}}^{k}\subset{\mathbb{T}}^{n}$, see Definition B.1. The homology class $H_{2m-2}(M;{\mathbb{Z}})/{\rm torsion}$ is spanned by the classes $[X_{1}],\ldots,[X_{n}]$ which are given by $n$ smooth divisors $X_{1},\ldots,X_{n}$ of $M$. For an $n$-tuple $d=(d_{1},\dots,d_{n})$ of nonnegative integers and a homology class $\alpha\in H_{*}(M;{\mathbb{Q}})$, we abbreviate the following genus zero Gromov-Witten invariants of $M$ with fixed marked points in class $A\in H_{2}(M;{\mathbb{Z}})$
${\rm GW}_{A}([X]^{*d},\alpha):={\rm GW}^{\{1,\ldots,\sum_{i}d_{i}+1\}}_{0,1+\sum_{i}d_{i},A}([X_{1}],\ldots,[X_{1}],\ldots,[X_{n}],\ldots,[X_{n}],\alpha)$ (1.7)
where each class $[X_{i}]$ occurs $d_{i}$ times. For each integer $\ell\geq 2$ and each nonzero vector $d=(d_{1},\ldots,d_{n})$ with each integer $d_{i}\geq 0$ in $\mathcal{D}^{\rm eff}(\tau)$ (this is a cone in ${\mathbb{Z}}^{n}$ corresponding to the set of effective classes in $H_{2}(M;{\mathbb{Z}})/{\rm torsion}$, see (A.2)), according to Theorem B.1, if $M$ is Fano then we have
${\rm GW}_{A_{\ell\xi_{d}}}([X]^{*\ell d^{+}},[pt])={\rm GW}_{0,1,0}([pt])\neq 0,$
where the homology classes $A_{\ell\xi_{d}}\in H_{2}(M;{\mathbb{Z}})/{\rm torsion}$ are effective classes given by (A.1) and (A.3), and the $n$-tuple $d^{+}=(d_{1}^{+},\ldots,d_{n}^{+})$ is given by $d_{i}^{+}:=\max\{d_{i},0\}$. Since each divisor $X_{i}$ has codimension two, using Theorem 1.5 and letting $\ell\to\infty$ in the above nonzero Gromov-Witten invariants, we recover the following Givental’s result.
Theorem 1.7 ([27, Theorem 1.1]).
Let $(M,\omega)$ be a Fano toric manifold with period $p>0$, $\omega(\pi_{2}(M))=p\cdot{\mathbb{Z}}$. Let $\mathcal{E}$ be the set of all nonzero effective homology classes of $M$. Then
#$\#{\rm Fix}(\varphi)\geq\sup_{A\in\mathcal{E}}\bigg{\lceil}\frac{p\cdot\langle c_{1}(TM),A\rangle}{\langle[\omega],A\rangle}\bigg{\rceil}.$
for any $\varphi\in\operatorname{Ham}(M,\omega)$. In particular, if $(M,\omega)$ is a monotone toric manifold with minimal Chern number $N\geq 1$, then each Hamiltonian diffeomorphism $\varphi$ has at least $N$ fixed points.
Remark 1.11.
In [27, Theorem 1.1], Givental assumed that the symplectic form is primitive, i.e., $\omega(\pi_{2}(M))={\mathbb{Z}}$. That’s why the symplectic period $p$ does not occur in his lower bound.
1.2 Organization of the paper
The paper is structured as follows: Section 2 covers preliminaries, including the Novikov ring, non-Archimedean normed spaces, big quantum homology, deformed Floer homology and PSS isomorphisms. Sectionetc 3 develops deformed Hamiltonian spectral invariants and proves the key Ljusternik-Schnirelman inequality. Section 4 establishes critical links between nonvanishing Gromov-Witten invariants and spectral invariant behavior. Section 5 provides the proofs for the main Theorems 1.11.5 on degenetate symplectic fixed point estimates. Appendices A and B apply the theory to coadjoint orbits of compact simple Lie groups and symplectic toric manifolds, respectively.
Acknowledgements
I am deeply indebted to Prof. Jinxin Xue for his help and encouragement and for his grants support NSFC 12271285 and the New Cornerstone investigator program. I thank Professors Guangcun Lu and Zhiyu Tian for valuable comments on Theorem 1.1 of this paper. I thank Prof. Tian-Jun Li and Shuo Zhang for useful remarks on Gromov-Witten invariants with two point insertions. I am grateful to Professors Kaoru Ono and Michael Usher for helpful discussions about spectral invariants with bulk studied in K. Fukaya, Y.-G. Oh, H. Ohta and K. Ono [22] and M. Usher [68], respectively.
2 Preliminaries
2.1 Starting notations
The set of spherical symplectic periods of $(M,\omega)$ is denoted by
$\Gamma_{\omega}=\{\langle[\omega],A\rangle\;|\;A\in\pi_{2}(M)\}.$
We say that a symplectic manifold $(M,\omega)$ is rational if there is a number $p>0$ such that $\Gamma_{\omega}=p\cdot{\mathbb{Z}}$.
Definition 2.1.
  1. (1)
    The Novikov ring associated to $\Gamma_{\omega}$ is the ring of formal Laurent series given by
    $\Lambda=\bigg{\{}\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\;\big{|}\;a_{i}\in{\mathbb{C}},\;\lambda_{i}\in\Gamma_{\omega},\;\forall C>0,\;\#\{i\in{\mathbb{N}}|a_{i}\neq 0,\lambda_{i}<C\}<\infty\bigg{\}}$
    where $T$ is a formal variable.
  2. (2)
    The valuation map on $\Lambda$ is the map
    $\nu:\Lambda\to\Gamma_{\omega},\quad\nu\big{(}\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\big{)}=\max\{-\lambda_{i}\;|\;a_{i}\neq 0\}.$
Denote
$\Lambda_{0}:=\bigg{\{}\sum_{i}a_{i}T^{g_{i}}\in\Lambda\;\big{|}\;g_{i}\geq 0\;\forall i\bigg{\}}.$
Note that $\Lambda$ is the field of fractions of the ring $\Lambda_{0}$. We define the map $\exp:\Lambda_{0}\to\Lambda_{0}$ by the usual Taylor series
$\exp(a)=\sum_{k=1}^{\infty}\frac{a^{k}}{k!}.$
It is not difficult to see that this is a well-defined map.
2.2 Non-Archimedean normed vector space
Definition 2.2.
Let $V$ be a vector space over $\Lambda$, and $\ell:V\to{\mathbb{R}}\cup\{-\infty\}$ a function. The pair $(V,\ell)$ is called a non-Archimedean normed vector space if
  • (i)
    $\ell(0)=-\infty$, and $\ell(c)\in{\mathbb{R}}$ for all nonzero $c\in V$;
  • (ii)
    $\ell(\alpha\cdot c)=\ell(c)+\nu(\alpha)$ for all $\alpha\in\Lambda$ and $c\in V$;
  • (iii)
    $\ell(c_{1}+c_{2})\leq\max\{\ell(c_{1}),\ell(c_{2})\}$ for all $c_{1},c_{2}\in V$.
Remark 2.3.
There is a symbolic difference between our property (ii) in the above definition and the property (F2) in [69, Definition 2.2] because of the symbolic difference for the valuation map $\nu$. It is shown that the equality in the (iii) holds if $\ell(c_{1})\neq\ell(c_{2})$, see [13, Proposition 2.1] or [69].
Definition 2.4.
A finite-dimensional non-Archimedean normed vector space $(V,\ell)$ over a Novikov field $\Lambda$ is called an orthogonalizable $\Lambda$-space if there exists an orthogonal basis for $C$, meaning that a finite ordered collection $(v_{1},\ldots,v_{n})$ of elements of $V$ with the property
$\ell\big{(}\sum_{i=1}^{n}\alpha_{i}v_{i}\big{)}=\max_{1\leq i\leq n}\ell(\alpha_{i}v_{i})$
for all $\alpha_{1},\ldots,\alpha_{n}\in\Lambda$.
Definition 2.5.
Let $(V,\ell)$ be a non-Archimedean normed vector space over a Novikov field $\Lambda$. We say that two subspaces $E$ and $F$ of $V$ are orthogonal if
$\ell(e+f)=\max\big{\{}\ell(e),\ell(f)\big{\}}$
for all $e\in E$ and $f\in F$.
Through a non-Archimedean Gram-Schmidt process, one can show that any subspace $U$ of an orthogonalizable $\Lambda$-space $(V,\ell)$ has an orthogonal complement $W$, i.e., $U\oplus W=V$, and $U$ and $W$ are orthogonal, see [69, Corollary 2.19].
Next we record the following standard result, i.e., the singular value decomposition for linear maps between orthogonalizable spaces, which will be used in the proof of Lemma 3.8.
Lemma 2.6 ([69, Theorem 3.5]).
Let $(V,\ell_{V})$ and $(W,\ell_{W})$ be two orthogonalizable $\Lambda$-spaces. Let $L:V\to W$ be a $\Lambda$-linear map, and $\xi$ a nonzero vector in ${\rm ker}(L)$. Then there is orthogonal bases $(\xi_{1},\ldots,\xi_{n})$ for $V$ and $(\eta_{1},\ldots,\eta_{m})$ for $W$ such that the following holds:
  • $L\xi_{i}=\eta_{i}$ for $i=1,\ldots,r$;
  • $\xi_{r+1}=\xi$;
  • $(\xi_{r+1},\dots,\xi_{n})$ is an orthogonal ordered basis for ${\rm ker}(L)$.
2.3 Big quantum homology ring
In this subsection we will review the big quantum homology which is closely related to the deformed spectral invarints.
Given $A\in H_{2}(M;{\mathbb{Z}})$, for homogeneous classes $a_{i}\in H_{*}(M;{\mathbb{Z}})/{\rm torsion},i=1\ldots,k$ with $k\geq 3$ which satisfy
$\sum_{i=1}^{k}(2n-{\rm deg}(a_{i}))=2n+2\langle c_{1}(TM),A\rangle+2k-6,$ (2.1)
we denote ${\rm GW}_{0,k,A}(a_{1},\ldots,a_{k})$ the Gromov-Witten invariant by counting (in the appropriate virtual sense) the number of $J$-holomorphic spheres in the class $A$ passing through the submanifolds $N_{i}$, where $J$ is a generic almost complex structure compactible with $\omega$ and the $N_{i}$ are generic cycles representing the classes $a_{i}$, see Ruan and Tian [58], Fukaya and Ono [20] and Ruan [59]. We put $\operatorname{GW}_{0,k,A}(a_{1},\ldots,a_{k})=0$ unless (2.1) holds, and define
${\rm GW}_{k}(a_{1},\ldots,a_{k})=\sum_{A\in H_{2}(M,{\mathbb{Z}})}{\rm GW}_{0,k,A}(a_{1},\ldots,a_{k})T^{\langle[\omega],A\rangle}.$
Then, by Gromov’s compactness theorem, one can extend the above definition linearly over $\Lambda$ to a $\Lambda$-module homomorphism
${\rm GW}_{k}:H_{*}(M;\Lambda)^{\otimes k}\longrightarrow\Lambda.$
For each $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, we define the $\zeta$-deformed quantum multiplication on $H_{*}(M;\Lambda)$ as follows: for any $a,b\in H_{*}(M;\Lambda)$, the product
$*_{\zeta}:H_{*}(M;\Lambda)\otimes H_{*}(M;\Lambda)\to H_{*}(M;\Lambda)$
is given by the formula
$\langle a*_{\zeta}b,c\rangle_{\rm PD_{M}}=\sum_{k=0}^{\infty}\frac{1}{k!}{\rm GW}_{k}(a,b,c,\underbrace{\zeta,\ldots,\zeta}_{k})$
for any $c\in H_{*}(M;\Lambda)$ where $\langle\cdot,\cdot\rangle_{\rm PD_{M}}$ denotes the Poincare duality. Alternately, one may express this product in terms of a homogeneous basis as follows. Let $\{c_{j}\}_{j=1}^{K}$ be a homogeneous basis for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}$, namely, $c_{i}\circ c^{j}=\delta_{i}^{j}$ where $\circ$ denotes the Poincaré intersection pairing. Then the product $*_{\zeta}$ can be obtained by extending linearly from the following formula
$a*_{\zeta}b=\sum_{A\in H_{2}(M;{\mathbb{Z}})}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}{\rm GW}_{0,k+3,A}(a,b,c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\langle[\omega],A\rangle}c^{j}.$
It is well-known that the product $*_{\zeta}$ is associative and graded commutative, see [39]. So we obtain a ${\mathbb{Z}}_{2}$-graded commutative ring
$QH_{*}(M,\omega)_{\zeta}=\big{(}H_{*}(M;\Lambda),*_{\zeta}\big{)}$
which obviously has the fundamental class $[M]$ as its unity, and which is called the $\zeta$-deformed homology ring of $(M,\omega)$.
Recall that the mixed Gromov-Witten invariant ${\rm GW}_{0,k,A}^{\{1,\ldots,l\}}(a_{1},\ldots,a_{l},b_{1},\ldots,b_{k})$ from Ruan and Tian [58] formally counts (with appropriate signs) the tuples $(u,\vec{w})=(u,\{w_{j}\}_{j=1}^{k})$ consisting of a $J$-holomorphic sphere $u:S^{2}\to M$ representing the class $A$ with $u(z_{i})\in\alpha_{i}$ and $u(w_{j})\in\beta_{j}$ for all $i,j$, where $\{z_{i}\}_{i=1}^{l}$ is a tuple of fixed pairwise distinct points on $S^{2}$ and $w_{j}\in S^{2}\setminus\{z_{i}\}_{i=1}^{l}$ for all $j$, and each $\alpha_{i}$ and $\beta_{j}$ are generic representatives of the classes $a_{i}$ and $b_{j}$ respectively. This quantity is zero unless
$\sum_{i=1}^{l}(2n-\deg(a_{i}))+\sum_{j=1}^{k}(2n-\deg(b_{j}))=2n+2\langle c_{1}(TM),A\rangle+2k.$ (2.2)
Using the splitting, multilinearity and symmetry properties of the mixed Gromov-invariants (cf. [58] or [48, Chapter 11]), one can see that an $l$-fold product is then obtained by extending linearly from the formula
$a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}=\sum_{A\in H_{2}(M;{\mathbb{Z}})}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}{\rm GW}_{0,l+k+1,A}^{I}(a_{1}\ldots,a_{l},c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\langle[\omega],A\rangle}c^{j}$ (2.3)
where $I=\{1,\ldots,l+1\}$, and $\{c_{j}\}_{j=1}^{K}$ is a homogeneous basis for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}$.
As a vector space over ${\mathbb{C}}$, the big quantum homology $H_{*}(M;\Lambda)$ is isomorphic to $H_{*}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda$. And each element $a\in H_{*}(M;\Lambda)$ can be written as $a=\sum_{g\in\Gamma_{\omega}}a_{g}T^{g}$ with $a_{g}\in H_{*}(M;{\mathbb{C}})$ and the property that for any $C>0$,
#$\#\{g\in\Gamma_{\omega}\;|\;a_{g}\neq 0,g<C\}<\infty.$
This latter finiteness condition gives rise to a valuation function $I_{\nu}:H_{*}(M;\Lambda)\to\Gamma_{\omega}$ defined by
$I_{\nu}(a)=\max\{-g|a_{g}\neq 0\}.$ (2.4)
There exists a pairing $\Delta:H_{*}(M;\Lambda)\times H_{*}(M;\Lambda)\to\Lambda$ defined by
$\Delta\bigg{(}\sum a_{g}T^{g},\sum b_{h}T^{h}\bigg{)}=\sum a_{g}\circ b_{h}\cdot T^{g+h}$
where $\circ$ stands for the Poincaré intersection pairing. It is easy to verify that the pairing is well-defined, and nondegenerate in the sense that $\Delta(a,b)=0$ for all $b$ then $a=0$. Then, by taking coefficients of the zero-order term in $\Delta(a,b)$ with respect to the formal variant $T$, one can define a ${\mathbb{C}}$-value nondegenerate paring ${\textstyle\prod}:H_{*}(M;\Lambda)\times H_{*}(M;\Lambda)\to{\mathbb{C}}$ by
$\prod\bigg{(}\sum a_{g}T^{g},\sum b_{h}T^{h}\bigg{)}=\sum a_{g}\circ b_{-g}$ (2.5)
(see page 1359 in [68] or [22, Chapter 15]).
2.3.1 Deformed quantum cuplength
Schwarz [61, Section 2.2] generalized the notion of classical cuplength to small quantum cohomology, called quantum cuplength, and employed it to obtain some estimates for the number of 1-periodic solutions of Hamiltonian equations. Now we further generalize this invariant to our deformed quantum homology as follows.
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, for homology classes $\alpha_{1},\ldots,\alpha_{k}\in H_{*}(M;\Lambda)$, using the splitting property of the Gromov invariants, the $k$-fold big quantum product can be represented as a formal series in $H_{*}(M;\Lambda)$
$\alpha_{1}*_{\zeta}\ldots*_{\zeta}\alpha_{k}=\sum(\alpha_{1}*_{\zeta}\ldots*_{\zeta}\alpha_{k})_{g}T^{g}.$
Definition 2.7.
The $\zeta$-deformed quantum cuplength for $g\in\Gamma_{\omega}$ is defined as
${{\rm qcl}_{\zeta}}(g)=\max\big{\{}k+1\;|\;\exists\;a_{i}\in H_{*<2n}(M;{\mathbb{C}})\;\hbox{such that}\;(a_{1}*_{\zeta}\ldots*_{\zeta}a_{k})_{g}\neq 0\big{\}}$
if there exist homology classes $a_{i}\in H_{*<2n}(M;{\mathbb{C}})$ such that $(a_{1}*_{\zeta}\ldots*_{\zeta}a_{k})_{g}\neq 0$, and we let ${{\rm qcl}_{\zeta}}(g)=0$ if $(a_{1}*_{\zeta}\ldots*_{\zeta}a_{k})_{g}=0$ for arbitrary nonzero classes $a_{1},\ldots,a_{k}\in H_{*}(M;{\mathbb{C}})$.
Remark 2.8.
Obviously, $g>0$ if ${{\rm qcl}_{\zeta}}(g)>0$. From the properties of Gromov-Witten invariants (especially, corresponding to the term for which there are no $\zeta$ as inserted classes, and no $J$-holomorphic spheres representing $A=0\in H_{*}(M;{\mathbb{Z}})$ in the relavent Gromov-Witten invariants), one can see that
$a_{1}*_{\zeta}\ldots*_{\zeta}a_{k}=a_{1}\cap\ldots\cap a_{k}+\sum_{0<g\in\Gamma_{\omega}}b_{g}T^{g}$
for any $a_{1},\ldots,a_{k}\in H_{*}(M;{\mathbb{C}})$ where $\cap$ stands for the usual homology intersection product. Therefore, the classical cuplength coincides with ${{\rm qcl}_{\zeta}}(0)$, and ${{\rm qcl}_{\zeta}}(g)=0$ for all $g\in\{\gamma\in\Gamma_{\omega}|\gamma<0\}$.
2.3.2 Morse theoretical big quantum homology
The $\zeta$-deformed quantum ring $H_{*}(M;\Lambda)$ has a Morse theoretical interpretation which we describe as follows. Let $\{\beta_{i}\}_{i=1}^{m}$ be a basis of $\oplus_{i=0}^{n-2}H_{2i}(M,{\mathbb{Q}})$. Thom’s theorem [66, Théorème II.29] implies that one may choose an embedding
$f_{i}:\Delta_{i}\longrightarrow M$ (2.6)
as a representative for each $\beta_{i}$ (i.e., $f_{i*}[\Delta_{i}]=\beta_{i}$), where each $\Delta_{i}$, $i=1,\ldots,m$, is a smooth closed oriented manifold of dimension $2d(i)$. For simplicity, hereafter we denote
$\delta(I):=\sum^{k}_{j=1}(2n-2d(i_{j})-2).$
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M,\Lambda_{0})$, we let $\theta\in\Omega^{2}(M;\Lambda_{0})$ be a closed Novikov-ring valued two-form which is Poincaré dual to the degree-$(2n-2)$ component of $\zeta$. Then we have
$\zeta-{\rm PD_{M}}[\theta]=\sum_{i=1}^{m}a_{i}\beta_{i}\in\bigoplus_{i=0}^{n-2}H_{2i}(M,\Lambda_{0})$
where $a_{i}\in\Lambda_{0}$ for all $i\in\{1,\ldots,m\}$. For any $I=(i_{1},\ldots,i_{k})\in\{1,\ldots,m\}^{k}$, we set
$a_{I}=a_{i_{1}}\ldots a_{i_{k}}.$
Let $f\in C^{\infty}(M,{\mathbb{R}})$ be a Morse function. Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, we choose a generic Riemannian metric $\rho$ on $M$ such that the stable and unstable manifolds $W^{s}(f),W^{u}(f)$ for $x,y\in{\rm Crit}(f)$ transverse both to each other and to the maps $f_{i}:\Delta_{i}\to M$. A grading
${\rm ind}_{f}:{\rm Crit}(f)\longrightarrow\{0,\ldots,2n\}$
is given by the Morse index. Denote
$\mathcal{M}_{x,y}(f,\rho)=\big{\{}\gamma\in C^{\infty}({\mathbb{R}},M)\;\big{|}\;\dot{\gamma}+{\rm grad}_{\rho}f(\gamma)=0,\;\gamma(-\infty)=x,\;\gamma(+\infty)=y\big{\}}.$
Since $\dim\widehat{\mathcal{M}}_{x,y}(f,\rho)={\rm ind}_{f}(x)-{\rm ind}_{f}(y)$ and ^$\widehat{\mathcal{M}}_{x,y}(f,\rho)$ carries a free ${\mathbb{R}}$-action, if ${\rm ind}_{f}(x)={\rm ind}_{f}(y)+1$ then the moduli space
$\mathcal{M}_{x,y}(f,\rho)=\widehat{\mathcal{M}}_{x,y}(f,\rho)/{\mathbb{R}}$
is a compact 0-dimensional manifold. Equipping unstable manifolds with arbitrary orientations, together with the orientation of $M$, gives rise to orientations on ^$\widehat{\mathcal{M}}_{x,y}(f,\rho)$, which, compared with the canonical orientation by the flow, assign a sign to every element $[\gamma]\in\mathcal{M}_{x,y}(f,\rho)$ provided that ${\rm ind}_{f}(x)={\rm ind}_{f}(y)+1$. Then we obtain a Morse complex
$CM_{k}(f;{\mathbb{Z}})=\sum_{\begin{subarray}{c}p\in{\rm Crit}(f),\\ {\rm ind}_{f}(p)=k\end{subarray}}{\mathbb{Z}}\langle p\rangle$
with differential
$\partial:CM_{k+1}(f;{\mathbb{Z}})\to CM_{k}(f;{\mathbb{Z}}),\quad\partial x=\sum_{{\rm ind}_{f}(x)={\rm ind}_{f}(y)+1}\sharp_{\rm alg}\mathcal{M}_{x,y}(f,\rho)\cdot y$
where $\sharp_{\rm alg}\mathcal{M}_{x,y}(f,\rho)\in{\mathbb{Z}}$ is the number of the negative gradient flowlines counted with sign. The homology of the complex $(CM(f;{\mathbb{Z}}),\partial)$ is called Morse homology of $(f,\rho)$ with integer coefficients, naturally isomorphic to the singular homology $H_{*}(M;{\mathbb{Z}})$.
We now consider a complex with $\Lambda$-coefficients
$CM_{*}(f;\Lambda)=CM_{*}(f;{\mathbb{Z}})\otimes_{{\mathbb{Z}}}\Lambda$
with differential $\partial^{f}:CM(f;\Lambda)\to CM(f;\Lambda)$ given by $\partial\otimes_{{\mathbb{Z}}}{\rm id}$. The homology of this complex is independent of the choice of $\rho$, and naturally isomorphic to $H_{*}(M;\Lambda)$. We denote it by $HM_{*}(f;\Lambda)$.
The $\zeta$-deformed quantum multiplication on $H_{*}(M;\Lambda)$ has the following Morse theoretical description. Let $(f,\rho^{0}),(g,\rho^{1}),(h,\rho^{2})$ be three Morse-Smale pairs as above. Let $J$ be an almost complex structure on $M$ which smoothly depends on $z\in{\mathbb{C}}P^{1}$. Given $A\in\pi_{2}(M)$, we consider the space of $J$-holomorphic spheres
$\mathcal{M}_{A}(J)=\{u:{\mathbb{C}}P^{1}\to M\;|\;du(z)\circ j=J(z,u(z))du(z),\;[u]=A\}.$
Now we define
$\mathcal{M}_{k}(A,J)=\big{\{}(u,\vec{z})\in\mathcal{M}_{A}(J)\times({\mathbb{C}}P^{1})^{k}\big{\}},$
and denote by ¯$\overline{\mathcal{M}}_{k}(A,J)$ the compactification of this space with evaluation maps $ev_{i}\;(1\leq i\leq k)$ at the $i$-th marked point. For the marked space $({\mathbb{C}}P^{1},0,-1,\infty)$, each element $(u,\vec{z})$ in $\mathcal{M}_{k}(A,J)$ carries the natural evaluation maps $ev_{0},ev_{-1},ev_{\infty}$ given by $ev_{0}(u,\vec{z})=u(0)$, $ev_{-1}(u,\vec{z})=u(-1)$ and $ev_{\infty}(u,\vec{z})=u(\infty)$. For $p\in\operatorname{Crit}(f),q\in\operatorname{Crit}(g),r\in\operatorname{Crit}(h)$, let $\iota^{0}:W^{u}(p)\to M$, $\iota^{1}:W^{u}(q)\to M$, $\iota^{2}:W^{s}(r)\to M$ denote the inclusion maps. The fiber product
¯$\overline{\mathcal{M}}_{k}(A,J)_{(ev_{0},ev_{-1},ev_{\infty},ev_{1},\ldots,ev_{k})}\times_{(\iota^{0},\iota^{1},\iota^{2},f_{i_{1}},\ldots,f_{i_{k}})}W^{u}(p)\times W^{u}(q)\times W^{s}(r)\times\Delta_{i_{1}}\times\ldots\times\Delta_{i_{k}}$
has a Kuranishi structure (cf. [20, 21, 23]) with corners of dimension
$\operatorname{ind}_{f}(p)+\operatorname{ind}_{g}(q)-\operatorname{ind}_{h}(r)+2\langle c_{1}(TM),A\rangle-2n-\delta(I).$
Then where the associated multisections are denoted by $\mathsf{s}_{P,I}$, the $\zeta$-deformed quantum product
$*_{\zeta}^{\rm Morse}:CM_{*}(f;\Lambda)\times CM_{*}(g;\Lambda)\longrightarrow CM_{*}(h;\Lambda)$
is defined by extending linearly from $p*_{\zeta}^{\rm Morse}q$ for $p\in\operatorname{Crit}(f)$ and $q\in\operatorname{Crit}(g)$ given by
$p*_{\zeta}^{\rm Morse}q=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}A\in\pi_{2}(M),\\ r\in\operatorname{Crit}(h)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \operatorname{ind}_{f}(p)+\operatorname{ind}_{g}(q)-\operatorname{ind}_{h}(r)=2n+\delta(I)-2c_{1}(A)\end{subarray}}|\mathsf{s}_{P,I}^{-1}(0)|\exp\big{(}\int_{A}\theta\big{)}a_{I}T^{\int_{A}\omega}r$
where $|\mathsf{s}_{P,I}^{-1}(0)|$ denotes the sum of the rational multiplicities of the points of the zero-dimensional vanishing locus of $\mathsf{s}_{P,I}$. On homology level (after identifying the Morse homology with the $\zeta$-deformed homology), the map $*_{\zeta}^{\rm Morse}$ induces a map from $H_{*}(M;\Lambda)\otimes H_{*}(M;\Lambda)$ to $H_{*}(M;\Lambda)$ by sending $a\otimes b$ to
$\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}\sum_{A\in\pi_{2}(M)}\sum_{I\in\{1,\ldots,m\}^{k}}{\rm GW}_{0,k+3,A}(a,b,c_{j},a_{i_{1}}\beta_{i_{1}},\ldots,a_{i_{k}}\beta_{i_{k}})\exp\big{(}\int_{A}\theta\big{)}T^{\int_{A}\omega}c^{j}$
where$\{c_{j}\}_{j=1}^{K}$ is a homogeneous basis for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}$. Since $\zeta={\rm PD_{M}}[\theta]+\sum^{m}_{i=1}a_{i}\beta_{i}$, it follows from the divisor axiom of Gromov-Witten invariants [48, Proposition 7.5.7] that the last formula is
$\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}\sum_{A\in\pi_{2}(M)}{\rm GW}_{0,k+3,A}(a,b,c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\int_{A}\omega}c^{j}$
which coincides with the $\zeta$-deformed quantum multiplication $*_{\zeta}$ on $H_{*}(M;\Lambda)$.
Remark 2.9.
In the construction of big quantum homology ring, one can avoid using Kuranishi structures by imposing the assumption that $(M,\omega)$ is strongly semipositive (i.e., for any $A\in\pi_{2}(M)$ with $02-n\leq\langle c_{1}(TM),A\rangle<$ 0 we have $\int_{A}\omega\leq 0$) and instead follow the constructions of Hofer and Salamon [32], see [68, Section 3.2.2]. This will not effect our applications in Section 1.1 since those examples listed there meet the latter condition. The generalization to all closed symplectic manifolds requires more subtle technical arguments, but less than the technical hurdles overcome in the construction of $A_{\infty}$–algebras associated to appropriate Lagrangian submanifolds [21], and hence one could just appeal to that work or [23].
2.4 Deformed Hamiltonian Floer homology
Following Usher [68] (or cf. [22]), we review the construction of deformed Hamiltonian Floer homology with additional algebraic structures.
2.4.1 Deformed Floer complexes
Let ${\mathcal{H}}=C^{\infty}(S^{1}\times M,{\mathbb{R}})$. For $H\in{\mathcal{H}}$, we denote by $\{\varphi_{H}^{t}\}_{t\in[0,1]}$ the Hamiltonian isotopy of $H$ which is given by integrating the time-dependent vector field $X_{H_{t}}$, where $H_{t}=H(t,\cdot)$ and $X_{H_{t}}$ is determined uniquely by $-dH_{t}=\omega(X_{H_{t}},\cdot)$. Denote by ${\rm Per}(H)$ the set of contractible $1$-periodic orbits of the Hamiltonian flow of $H$. We say that $H$ is normalized if $\int^{1}_{0}H_{t}\omega^{n}=0$ for all $t\in S^{1}$.
We consider the set $\mathcal{L}_{pt}(M)$ of all pairs $(x,\bar{x})$ where $x:S^{1}\to M$ is a loop and $\bar{x}:D^{2}\to M$ a disc with $\bar{x}|_{\partial D^{2}}=x$, called a capping of $x$. We denote by $\pi_{2}(x)$ the set of homotopy classes of such $\bar{x}^{\prime}s$. We identify $(x,\bar{x})$ and $(x^{\prime},\bar{x}^{\prime})$ if $x=x^{\prime}$ and $\bar{x}$ is homotopic to $\bar{x}^{\prime}$ relative to the boundary $x$.
Given $H\in{\mathcal{H}}$, we consider the action functional $\mathcal{A}_{H}:\mathcal{L}_{pt}(M)\longrightarrow{\mathbb{R}}$ given by
[…]
Let ${\rm Crit}(\mathcal{A}_{H})$ denote the set of critical points of $\mathcal{A}_{H}$. It is easy to see that
${\rm Crit}(\mathcal{A}_{H})=\big{\{}(x,\bar{x})\in\mathcal{L}_{pt}(M)\;\big{|}\;\dot{x}(t)=X_{H_{t}}(x(t))\big{\}}.$
For each $H\in{\mathcal{H}}$, we define the action spectrum of $H$ as
${\rm Spec}(H)=\big{\{}\mathcal{A}_{H}(x,\bar{x})\;\big{|}\;(x,\bar{x})\in{\rm Crit}(\mathcal{A}_{H})\big{\}}.$
It is shown that the set ${\rm Spec}(H)$ has measure zero in ${\mathbb{R}}$ for any $H\in{\mathcal{H}}$ for arbitrary $(M,\omega)$ by Oh [50] and for aspherical $(M,\omega)$ by Schwarz [60], and that for any one-parameter smooth family of normalized periodic Hamiltonians $H^{s}:S^{1}\times M\to{\mathbb{R}},s\in[0,1]$ with
$\varphi_{H^{s}}^{0}\equiv id,\quad\varphi_{H^{s}}^{1}\equiv\varphi_{H^{0}}^{1}\quad\hbox{for all}\;s\in[0,1],$
we have ${\rm Spec}(H^{s})={\rm Spec}(H^{0})$, see [22, Corollary 10.3].
We say that $H$ is nondegenerate if for each $p\in{\rm Fix}(\varphi_{H}^{1})$, $1$ is not the eigenvalue of the linearization $d_{p}\varphi_{H}^{1}:T_{p}M\to T_{p}M$. Clearly, for a nondegenerate Hamiltonian $H$, the cardinality of the set ${\rm Per}(H)$ of one-periodic solutions of $\dot{x}(t)=X_{H_{t}}(x(t))$ is finite. In this case, we have an integer grading
$\mu_{H}:{\rm Crit}(\mathcal{A}_{H})\longrightarrow{\mathbb{Z}}$
induced by the Conley-Zehnder index. More specifically, in this paper, the Conley-Zehnder index is the negative of that in [11], and has the property that for a $C^{2}$-small Morse function $f$, each constant Hamiltonian one-periodic orbit $q$ with its constant capping $\bar{q}$ satisfies $\mu_{H}(q,\bar{q})=n-{\rm ind}_{q}(f)$, where ${\rm ind}_{q}(f)$ denotes the Morse index of $f$ at the critical point $q$.
We introduce an equivalence relation $\sim$ on ${\rm Crit}(\mathcal{A}_{H})$ such that $(x_{1},\bar{x}_{1})\sim(x_{2},\bar{x}_{2})$ if and only if $x_{1}=x_{2}$ and
$\int_{D^{2}}\bar{x}_{1}^{*}\omega=\int_{D^{2}}\bar{x}_{2}^{*}\omega.$
Denote by $\mathcal{O}(H)$ the set of such equivalence classes. The equivalence class of $(x,\bar{x})\in{\rm Crit}(\mathcal{A}_{H})$ is denoted by $[x,\bar{x}]$. Then we have the maps
${\rm Crit}(\mathcal{A}_{H})\longrightarrow\mathcal{O}(H)\longrightarrow{\rm Per}(H)$
by $(x,\bar{x})\mapsto[x,\bar{x}]$ and $[x,\bar{x}]\mapsto x$.
We define the Floer complex $CF_{*}(H)$ of a nondegenerate Hamiltonian $H\in{\mathcal{H}}$ as the set of elements having the formal sum
$\sum_{[x,\bar{x}]\in\mathcal{O}(H)}\alpha_{[x,\bar{x}]}[x,\bar{x}],\;\alpha_{[x,\bar{x}]}\in{\mathbb{C}}$ (2.7)
with the property that for any $C>0$,
#$\#\big{\{}[x,\bar{x}]|\alpha_{[x,\bar{x}]}\neq 0,\mathcal{A}_{H}([x,\bar{x}])>C\big{\}}<\infty.$
The Novikov field $\Lambda$ acts on $CF_{*}(H)$ as follows: for each $\lambda\in\Gamma_{\omega}$, we pick an arbitrary class $A_{\lambda}\in\pi_{2}(M)$ satisfying $\langle[\omega],A_{\lambda}\rangle=\lambda$ and define the action
$\big{(}\sum a_{i}T^{\lambda_{i}}\big{)}\cdot\big{(}\sum\beta_{[x,\bar{x}]}[x,\bar{x}]\big{)}=\sum a_{i}\beta_{[x,\bar{x}]}[x,\bar{x}\sharp A_{\lambda_{i}}].$
It is easy to verify that this action is well-defined. Consequently, $CF_{*}(H)$ is a vector space over $\Lambda$ of finite dimension equal to $\sharp{\rm Per}(H)$. More precisely, for each $x\in{\rm Per}(H)$, if we pick a lifting $[x,\bar{x}]\in\mathcal{O}(H)$, then any element $c$ of $CF_{*}(H)$ is uniquely written as a finite sum
$c=\sum_{x\in{\rm Per}(H)}a_{x}[x,\bar{x}],\;a_{x}\in\Lambda.$ (2.8)
Definition 2.10.
The filtration function $\ell_{H}$ on $CF_{*}(H)$ is defined by
$\ell_{H}(c)=\max\big{\{}\mathcal{A}_{H}([x,\bar{x}])\;\big{|}\;\alpha_{[x,\bar{x}]}\neq 0\big{\}}$
for any $c=\sum\alpha_{[x,\bar{x}]}[x,\bar{x}]$ of the form (2.7).
Equivalently, after fixing a lifting $x\to[x,\bar{x}]$ for each $x\in{\rm Per}(H)$, one can define the value of $\ell_{H}$ at $c$ given in the form (2.8) by
$\ell_{H}(c)=\max\{\mathcal{A}_{H}([x,\bar{x}])+\nu(a_{x})|a_{x}\neq 0\}$
where $\nu$ is the valuation map on $\Lambda$ (cf. Definition 2.1(2)). It is easy to verify that $(CF_{*}(H),\ell_{H})$ is an orthogonalizable $\Lambda$-space. Moreover, $\ell_{H}$ gives rise to a filtration on $CF_{*}(H)$ by
$CF_{*}^{\lambda}(H)=\{c\in CF_{*}(H)\;|\;\ell_{H}(c)\leq\lambda\}\quad\hbox{for any}\;\lambda\in{\mathbb{R}}.$
2.4.2 Gradings
As a vector space over $\Lambda$, $CF_{*}(H)$ carries a natural ${\mathbb{Z}}_{2}$-grading which we define as follows.
[…]
Note that the parity of $\mu_{H}(x,\bar{x})$ depends only on $x\in{\rm Per}(H)$ but not on its capping $\bar{x}$, since for any two cappings $w$ and $u$ of $x$, we have
$\mu_{H}(x,w)-\mu_{H}(x,u)=2\langle c_{1}(TM),[w\sharp u^{\prime}]\rangle.$
Here $w\sharp u^{\prime}:S^{2}\to M$ is a $2$-sphere given by gluing $w$ and $u^{\prime}$ along $x$ where $u^{\prime}$ is the $u$ with opposite orientation.
2.4.3 Maslov index of homotopy classes of cylinders
Let $H_{\pm}\in{\mathcal{H}}$ and $x_{\pm}\in{\rm Per}(H_{\pm})$. We denote by $\pi_{2}(x_{-},x_{+})$ the set of path components of the space of continuous maps $u:[-\infty,+\infty]\times S^{1}\to M$ satisfying $u(\{\pm\infty\}\times S^{1})=x_{\pm}$.
Given $H_{i}\in{\mathcal{H}}$ and $x_{i}\in{\rm Per}(H_{i}),i=0,1,2$, there is the natural gluing map
$\pi_{2}(x_{0},x_{1})\times\pi_{2}(x_{1},x_{2})\longrightarrow\pi_{2}(x_{0},x_{2})$
$(C_{0},C_{1})\longmapsto C_{0}\sharp C_{1}$
where $C_{0}\sharp C_{1}$ is the equivalence class of a gluing map $w$ of any two representatives $u,v$ of $C_{0}$ and $C_{1}$. More precisely, let $u,v:[-\infty,+\infty]\times S^{1}\to M$ be two continuous maps which represent $C_{0}$ and $C_{1}$ respectively, and define $w:[-\infty,+\infty]\times S^{1}\to M$ by putting $w(s,t)=u(-\log(-s),t)$ for $s\leq 0$ and $w(s,t)=v(\log(s),t)$ for $s\geq 0$. Here $\log:[0,\infty]\to[-\infty,\infty]$ is an extension of the natural logarithm.
Similarly, one can define a concatenation map
$\pi_{2}(x_{0})\times\pi_{2}(x_{0},x_{1})\longrightarrow\pi_{2}(x_{1}),\quad(\omega,C)\longmapsto\omega\sharp C.$
For $x_{\pm}\in{\rm Per}(H_{\pm})$, we now use the Conley-Zehnder index to define a map $\mu$ on $\pi_{2}(x_{-},x_{+})$ as follows. For $C\in\pi_{2}(x_{-},x_{+})$, we pick an arbitrary representative $u:[-\infty,+\infty]\times S^{1}\to M$ of $C$, and pick any disc $w$ such that $w|_{\partial D^{2}}=x_{-}$ (note that $x_{-}$ is contractible by our assumption). We then glue the negative end of $u$ to the boundary of $w$ along $x_{-}$ to obtain a disc $w\sharp u:D^{2}\to M$ with $w\sharp u|_{\partial D^{2}}=x_{+}$. Then we define
$\mu(C)=\mu_{H_{+}}(x_{+},w\sharp u)-\mu_{H_{-}}(x_{-},w).$
Clearly, $\mu(C)$ is independent of the choice of $u$ and $w$, and has the property
$\mu(C_{0}\sharp C_{1})=\mu(C_{0})+\mu(C_{1})\quad\hbox{for all}\;C_{0}\in\pi_{2}(x_{0},x_{1}),C_{1}\in\pi_{2}(x_{1},x_{2}).$
2.4.4 Deformed Floer boundary maps
Definition 2.11.
We say that $H\in{\mathcal{H}}$ is strongly nondegenerate if for each $p\in{\rm Fix}(\varphi_{H}^{1})$, $1$ is not an eigenvalue of the linearization $d_{p}\varphi_{H}^{1}:T_{p}M\to T_{p}M$, and each submanifold $f_{i}(\Delta_{i})$, $i=1,\ldots,m$, does not intersect the set $\{\varphi_{H}^{t}(p)|t\in[0,1]\}$.
By the Sard-Smale theorem, one can show that the space of strongly nondegenerate Hamiltonians is residual in ${\mathcal{H}}$ in the $C^{l}$-topology for all $2\leq l\leq\infty$, see Appendix A in [68].
Let $J=\{J_{t}\}_{t\in S^{1}}$ be an $S^{1}$-family of $\omega$-compatible almost complex structures on $M$, and $H\in{\mathcal{H}}$ a strongly nondegenerate Hamiltonian. Given $x_{\pm}\in{\rm Per}(H)$, we consider the solution $u:{\mathbb{R}}\times S^{1}\to M$ of the Floer equation:
¯$\overline{\partial}_{J,H}u:=\partial_{s}u+J(t,u(s,t))\big{(}\partial_{t}u-X_{H}(t,u(s,t))\big{)}=0$
with the asymptotic boundary condition $u(s,t)=x_{\pm}(t)$ for $s\to\pm\infty$. Then $u$ extends continuously to a map $u:[-\infty,\infty]\times S^{1}\to M$ with $u(\pm\infty,t)=x_{\pm}(t)$, which represents some class $C\in\pi_{2}(x_{-},x_{+})$. We denote by ^$\widehat{\mathcal{M}}(x_{-},x_{+},C;H,J)$ the moduli space of such solutions with
[…]
where $|\cdot|_{J_{t}}$ is the norm associated to an $S^{1}$-family of metrics induced by $J$ for all $t$.
Consider the product space
^$\widehat{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)=\widehat{\mathcal{M}}(x_{-},x_{+},C;H,J)\times({\mathbb{R}}\times S^{1})^{k}.$
Note that ${\mathbb{R}}$ acts on each factor by $s$-translation, we have a quotient
$\mathcal{M}_{k}(x_{-},x_{+},C;H,J)=\widehat{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)/{\mathbb{R}}.$
Using the constructions of Kuranishi structures from [20, Theorem 19.44], one can show that the compactification of $\mathcal{M}_{k}(x_{-},x_{+},C;H,J)$, denoted by ¯$\overline{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)$, admits an oriented Kuranishi structure with corners, see [68, Section 3.2.2]. Futhermore, one has evaluation maps
$\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}:\overline{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)\longrightarrow M$
given by evluating at the $i$-th marked point, i.e. $\operatorname{ev}_{i}([u,z_{1},\ldots,z_{k}])=u(z_{i})$ for all $[u,z_{1},\ldots,z_{k}]$ in the dense subset $\mathcal{M}_{k}(x_{-},x_{+},C;H,J)$ and $i\in\{1,\ldots,k\}$. Given an ordered set $I=(i_{1},\ldots,i_{k})$ with each $i_{j}\in\{1,\ldots,m\}$, we consider the fiber product
~$\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I}):=\overline{\mathcal{M}}(x_{-},x_{+},C;H,J)_{\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}}\times_{f_{i_{1}},\ldots,f_{i_{k}}}(\Delta_{i_{1}}\times\cdots\times\Delta_{i_{k}}).$
It follows from [21, Appendix A1] or [23, Section 3] that there is an oriented Kuranishi structure with corners on ~$\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})$ which has the (virtual) dimension
$\dim\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})=\mu(C)-\sum^{k}_{j=1}(2n-2d(i_{j})-2)-1.$
When $\mu(C)=\delta(I)+1$, we let $\mathsf{s}_{C,I}$ denote the multisection associated to Kuranishi structure on ~$\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})$, and denote $|\mathsf{s}_{C,I}^{-1}(0)|$ the sum of the rational multiplicities of the points of the zero-dimensional vanishing locus of $\mathsf{s}_{C,I}$.
Following [68, Proposition 3.5] (see also [22, Chapter 6]), we will deform the standard Floer differential by a general homology class $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M,\Lambda_{0})$. For any $[x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\in\mathcal{O}(H)$, we denote
$\mathsf{n}_{(H,J);k}\big{(}[x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\big{)}=\sum_{\begin{subarray}{c}C\in\pi_{2}(x_{-},x_{+})\\ [x_{+},\bar{x}_{+}]=[x_{+},\bar{x}_{-}\sharp C]\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k}\\ \mu(C)=\delta(I)+1\end{subarray}}\frac{1}{k!}|\mathsf{s}_{C,I}^{-1}(0)|\exp\big{(}\int_{C}\theta\big{)}a_{I}.$
Here $\bar{x}_{-}\sharp C$ stands for the concatenation of $\bar{x}$ and any representative $u$ of the class $C$. Clearly, the equivalent class $[x_{+},\bar{x}_{-}\sharp C]$ does not depend on the choice of $u$. Now we define the map $\partial_{\zeta}^{H,J}:CF_{*}(H)\to CF_{*}(H)$ by extending $\Lambda$-linearly from
$\partial_{\zeta}^{H,J}[x_{-}\bar{x}_{-}]=\sum_{[x_{+},\bar{x}_{+}]\in\mathcal{O}(H)}\sum_{k=0}^{\infty}\mathsf{n}_{(H,J);k}\big{(}[x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\big{)}[x_{+},\bar{x}_{+}].$
Then Propositions 3.5 and 3.6 in [68] imply the following.
Proposition 2.12.
For any $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M,\Lambda_{0})$, the $\partial_{\zeta}^{H,J}$ as above is a well-defined endomorphism and satisfies
$\big{(}\partial_{\zeta}^{H,J}\big{)}^{2}=0.$
With this preparation regarding the complex, now we can define the $\zeta$-deformed Floer homology of $(CF_{*}(H),\partial_{\zeta}^{H,J})$ as follows.
Definition 2.13.
The homology of the complex $(CF_{*}(H),\partial_{\zeta}^{H,J})$ is defined by
$HF_{\zeta}(H)=\frac{{\rm ker}\;\partial_{\zeta}^{H,J}}{{\rm im}\;\partial_{\zeta}^{H,J}}.$
It is shown in [68, Proposition 3.8] that for a strongly nondegenerate $H\in{\mathcal{H}}$, $HF_{\zeta}(H)$ is independent of the choice of the almost complex structure $J$. This is the reason why we suppress the $J$ from the notation of the deformed Floer homology. Note that the differential $\partial_{\zeta}^{H,J}$ preserves the ${\mathbb{C}}$-space $CF_{*}^{\lambda}(H)$, the corresponding homology is denoted by $HF_{\zeta}^{\lambda}(H)$, called filtered Floer homology of $H$.
2.4.5 Pair of pants products
For any Hamiltonian $H:S^{1}\times M\to{\mathbb{R}}$, one can modify $H$ in a standard way such that the associated Hamiltonian isotopy is constant near $[1]\in S^{1}$, namely, letting $H^{\prime}(t,x)=\chi^{\prime}(t)H(\chi(t),x)$ for some smooth function $\chi:[0,\frac{1}{2}]\to[0,1]$ with the property that $\chi=0$ near $t=0$ and $\chi=1$ near $t=\frac{1}{2}$. In what follows, we always assume that for some small $\epsilon>0$, $H(t,0)\equiv 0$ for $|t|<\epsilon$. Given two such Hamiltonians $H,K$, we define a new Hamiltonian $H*K:S^{1}\times M\to{\mathbb{R}}$ by
$H*K(t,x)=\begin{cases}2H(2t,x)&0\leq t\leq 1/2,\\ 2K(2t-1,x)&1/2\leq t\leq 1.\\ \end{cases}$
We now assume that $H,K$ and $H*K$ are strongly nondegenerate Hamiltonians. Then there exists a deformed version of pair of pants product
¯$\bar{*}_{\zeta}^{\rm PP}:CF_{*}(H)\otimes CF(K)\to CF(H*K)$
which resembles the construction of the product on Hamiltonian Floer homology in [55] (see also [43]). The precise definition of the product ¯$\bar{*}_{\zeta}^{\rm PP}$ is not used in this paper, and hence we will not go into details about it. Here we only mention that ¯$\bar{*}_{\zeta}^{\rm PP}$ is a chain map and restricts to a map
¯$\bar{*}_{\zeta}^{\rm PP}:CF^{\lambda}(H)\otimes CF^{\mu}(K)\to CF^{\lambda+\mu}(H*K)$
for all $\lambda,\mu\in{\mathbb{R}}$, which,on homology level, induces the map
$*_{\zeta}^{\rm PP}:HF^{\lambda}(H)\otimes HF^{\mu}(K)\to HF^{\lambda+\mu}(H*K),$
see [68, Proposition 3.11].
2.5 PSS isomorphisms
Following [68, Proposition 3.3.2], we will define a deformed Piunikhin-Salamon-Schwarz isomorphism from the Morse complex $(CM(f;\Lambda),\partial^{f})$ to the $\zeta$-deformed Floer complexes $(CF_{*}(H),\partial_{\zeta}^{H,J})$. Such kind of isomorphism has been studied in [55, 43].
For a generic Morse-Smale pair $(f,\rho)$ and a strongly nondegenerate Hamiltonian $H$ on $M$, the deformed PSS map
$\psi^{\operatorname{PSS}}_{\zeta}:(CM(f;\Lambda),\partial^{f})\longrightarrow(CF_{*}(H),\partial_{\zeta}^{H,J})$
is defined by emumerating “spiked discs” with incidence conditions corresponding to $I=(i_{1},\ldots,i_{k})$ as follows.
Let $\chi:{\mathbb{R}}\to[0,1]$ be a smooth cutoff function satisfying $\chi(s)=0$ for $s\leq 0$, $\chi(s)=1$ for $s\geq 1$ and $\chi^{\prime}(s)\geq 0$ for all $s$. Given an $S^{1}$-dependent family of almost complex structures $J$, let $\{J_{s}\}_{s\in[0,1]}$ be a two-parameters family of almost complex structures such that $J_{0,t}=J_{0}$, $J_{1,t}=J_{t}$ and $J_{s,0}=J_{0}$ for all $s,t\in[0,1]$. Here $J_{0}$ is a time-independent almost complex structure compatible with $\omega$. We define the ×${\mathbb{R}}\times S^{1}$-dependent pair $(H_{\chi},J_{\chi})$ by
$H_{\chi}(s,t)=\chi(s)H_{t},\quad J_{\chi}(s,t)=J_{\chi(s),t}.$
Given $\gamma\in{\rm Per}(H)$ and $C\in\pi_{2}(\gamma)$, we choose any disc $u$ representing $C$ and denote $\mu(C)=\mu_{H}(\gamma,u)$. Let $\mathcal{M}^{\operatorname{PSS}}(\gamma)$ denote the space of the solutions $u:{\mathbb{R}}\times S^{1}\to M$ of the Floer equation
$\partial_{s}u+J_{\chi}\big{(}\partial_{t}u-X_{H_{\chi}}(u)\big{)}=0$ (2.9)
with asymptotic boundary condition $u(+\infty,t)=\gamma(t)$ for all $t$ and finite energy
[…]
Note that since $H_{\chi}\equiv 0$ and $J_{\chi}\equiv J_{0}$ for $s<0$, the equation (2.9) becomes $J_{0}$-homomorphic curve on $(-\infty,0)\times S^{1}$ with finite energy, the removable singularity theorem implies that $\lim_{s\to-\infty}u(s,t)$ exists and does not depend on $t$. So one can define the map $ev_{-\infty}:\mathcal{M}^{\operatorname{PSS}}(\gamma)\to M$ by
$\operatorname{ev}_{-\infty}(u)=\lim_{s\to-\infty}u(s,t).$
We now define
$\mathcal{M}^{\operatorname{PSS}}_{k}(\gamma,C)=\big{\{}(u,\vec{z})\in\mathcal{M}^{\operatorname{PSS}}(\gamma)\times({\mathbb{R}}\times S^{1})^{k}\;\big{|}\;[u]=C\big{\}}$
and write its standard compactification as ¯$\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(\gamma,C)$ with evaluation maps $ev_{i},i\in\{1,\ldots,k\}$ at the $i$-th marked point and the evaluation map $ev_{-\infty}$ by extending the map $ev_{-\infty}$ on $\mathcal{M}^{\operatorname{PSS}}(\gamma)$ as above to the compactification.
Given $p\in\operatorname{Crit}(f)$ and $I=(i_{1},\ldots,i_{k})\in\{1,\ldots,m\}^{k}$, let $\iota_{p}:W^{u}(p)\hookrightarrow M$ denote the inclusion of the unstable manifold of $p$. Consider the fiber product
$\mathcal{M}^{\operatorname{PSS}}_{k}(\gamma,C,p;\Delta_{I})=\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(\gamma,C)_{\operatorname{ev}_{-\infty},\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}}\times_{\iota_{p},f_{i_{1}},\dots,f_{i_{k}}}\big{(}W^{u}(p)\times\Delta_{i_{i}}\times\cdots\times\Delta_{i_{k}}\big{)}$
which has a Kuranishi structure with corners of dimension
$\mu(C)+\operatorname{ind}_{f}(p)-n-\delta(I).$
Let $\mathsf{s}_{\operatorname{PSS}}$ denote the corresponding multisection. Then the PSS map
$\psi^{\operatorname{PSS}}_{\zeta,H}:(CM(f;\Lambda),\partial^{f})\longrightarrow(CF_{*}(H),\partial_{\zeta}^{H,J})$
is defined by extending $\Lambda$-linearly from
$\psi^{\operatorname{PSS}}_{\zeta,H}p=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}\gamma\in{\rm Per}(H),\\ C=[u]\in\pi_{2}(\gamma)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \mu(C)+\operatorname{ind}_{f}(p)=n+\delta(I)\end{subarray}}|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|\exp\big{(}\int u^{*}\theta\big{)}a_{I}[\gamma,u].$
It is shown in [68, Proposition 3.9] that the deformed PSS map $\psi^{\operatorname{PSS}}_{\zeta,H}$ is a chain map which induces a ring isomorphism (over $\Lambda$)
$\Psi^{\rm PSS}_{\zeta,H}:HM(f;\Lambda)\longrightarrow HF_{\zeta}(H).$
Moreover, this map respects the ${\mathbb{Z}}_{2}$-gradings of the respective complexes. Here we remark that the map $\Psi^{\operatorname{PSS}}_{\zeta,H}$ is independent of the choice of the function $\chi$, the metric $\rho$ and the almost complex structures $\{J_{s}\}_{s\in[0,1]}$.
3 Spectral invariants
3.1 Minmax critical value selector
Fix a ground field ${\mathbb{F}}$. Let $X$ be a closed smooth manifold of positive dimension. We will denote by $H_{*}(X)$ the singular homology of $X$ with coefficient field ${\mathbb{F}}$. Let $f\in C^{\infty}(X,{\mathbb{R}})$ be a smooth function. For any $\lambda\in{\mathbb{R}}$, we define
$X^{\lambda}:=\{x\in X\;|\;f(x)<\lambda\}.$
To a non-zero singular homology class $a\in H_{*}(X)\setminus\{0\}$, we associate a numerical invariant defined by
$c_{LS}(a,f)=\inf\{\lambda\in{\mathbb{R}}\;|\;a\in{\rm im}(i^{\lambda}_{*})\},$
where $i^{\lambda}_{*}:H_{*}(X^{\lambda})\to H_{*}(X)$ is the map induced by the inclusion $i^{\lambda}:X^{\lambda}\to X$. The function
$c_{LS}:H_{*}(X)\setminus\{0\}\times C^{\infty}(X,{\mathbb{R}})\to{\mathbb{R}}$
is called the minmax critical value selector of $f$. The following properties are well-known from classical Ljusternik-Schnirelman theory; see, e.g., [34, 9, 12, 26].
Proposition 3.1.
The minmax critical value selector $c_{LS}$ satisfies the following properties.
  1. 1.
    $c_{LS}(a,f)$ is a critical value of $f$, and $c_{LS}(ka,f)=c_{LS}(a,f)$ for any nonzero $k\in{\mathbb{F}}$.
  2. 2.
    $c_{LS}(a,f)$ is Lipschitz continuous in $f$ with respect to the $C^{0}$-distance.
  3. 3.
    Let $[pt]$ and $[X]$ denote the point class and the fundamental class, respectively. Then
    $c_{LS}([pt],f)=\min f\leq c_{LS}(a,f)\leq\max f=c_{LS}([X],f).$
  4. 4.
    If $b\neq k[X]$ and $c_{LS}(a\cap b,f)=c_{LS}(a,f)$, then the set
    $\Sigma=\{x\in\operatorname{Crit}(f)\;|\;f(x)=c_{LS}(a,f)\}$
    is homologically non-trivial, i.e. for every open neighborhood $V$ of $\Sigma$ the map $i_{*}:H_{k}(V)\to H_{k}(X)$ induced by the inclusion $i:V\hookrightarrow X$ is non-zero for some $k>0$. Clearly, homologically nontrivial sets are infinite sets.
3.2 Deformed Hamiltonian spectral invariants
In this section, we follow the works of Schwarz [60] and Oh [52] to define the spectral invariants and review their properties which will be useful in this paper.
For $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$ and a strongly non-degenerate Hamiltonian $H\in{\mathcal{H}}$, as in the previous section we have the PSS isomorphism
$\Psi^{\rm PSS}_{\zeta,H}:H_{*}(M;\Lambda)\longrightarrow HF_{\zeta}(H).$
Given $a\in H_{*}(M;\Lambda)\setminus\{0\}$, we put
$\rho_{\zeta}(a,H)=\inf\big{\{}\lambda\;|\;\Psi^{\rm PSS}_{\zeta,H}(a)\in\operatorname{im}(i^{\lambda})\big{\}}$
where $i^{\lambda}:HF_{\zeta}^{\lambda}(H)\to HF_{\zeta}(H)$ is the natural map induced by the inclusion of $CF_{*}^{\lambda}(H)$ to $CF_{*}(H)$. Then $-\infty<\rho_{\zeta}(a,H)<+\infty$, and for any two strongly nondegenerate Hamiltonians $H,K\in{\mathcal{H}}$, $\rho_{\zeta}(a,\cdot)$ satisfies
[…]
This has been proved in [68], see also Oh [52] for the undeformed case. Recall that the Hofer norm on the space of continuous functions $H:S^{1}\times M\to{\mathbb{R}}$ is given by
[…]
So if $H,K$ are both normalized then we get
$|\rho_{\zeta}(a,H)-\rho_{\zeta}(a,K)|\leq\|H\|$
[…]
for all nonzero classes $a\in H(M,\Lambda)$. Then one can further extend $\rho_{\zeta}(a,\cdot)$ to a map on $C^{0}(S^{1}\times M,{\mathbb{R}})$. The numbers $\rho_{\zeta}(a,H)$ are called deformed spectral invariants of an element $H\in C^{0}(S^{1}\times M,{\mathbb{R}})$ (if $a=0$, we put $\rho_{\zeta}(0,H)=-\infty$).
Proposition 3.2 ([68, Proposition 3.13] or [22, Theorem 7.8]).
The deformed spectral invariant
$\rho_{\zeta}:H_{*}(M;\Lambda)\times C^{\infty}(S^{1}\times M,{\mathbb{R}})\to[-\infty,\infty)$
has the following properties:
  • (P1)
[…]
  • (P2)
    Spectrality: If $H$ is strongly nondegenerate, then there is $c\in CF_{*}(H)$ such that $[c]=\Psi^{\rm PSS}_{\zeta,H}(a)$ and $\rho_{\zeta}(a,H)=\ell_{H}(c)$, and hence $\rho_{\zeta}(a,H)\in\operatorname{Spec}(H)$. If $M$ is rational, for any $H\in{\mathcal{H}}$ one also has $\rho_{\zeta}(a,H)\in\operatorname{Spec}(H)$.
  • (P3)
    Homotopy invariance: $\rho_{\zeta}(a,H)=\rho_{\zeta}(a,K)$, when $\varphi_{H}=\varphi_{K}$ in the universal covering of the group of Hamiltonian diffeomorphisms $\widetilde{{\rm Ham}}(M,\omega)$, and $H,K$ are normalized.
  • (P4)
    Quantum shift: $\rho_{\zeta}(\lambda a,H)=\rho_{\zeta}(a,H)+\nu(\lambda)$ for all $\lambda\in\Lambda$.
  • (P5)
    Projective invariance: $\rho_{\zeta}(wa,H)=\rho_{\zeta}(a,H)$ for any $0\neq w\in{\mathbb{C}}$.
  • (P6)
    Triangle inequality: $\rho_{\zeta}(a*_{\zeta}b,H*K)\leq\rho_{\zeta}(a,H)+\rho_{\zeta}(a,K)$ for any $H,K\in{\mathcal{H}}$.
  • (P7)
    Valuation inequality: $\rho_{\zeta}(a+b,H)\leq\max\{\rho_{\zeta}(a,H),\rho_{\zeta}(b,H)\}$. And if $\rho_{\zeta}(a,H)\neq\rho_{\zeta}(b,H)$, then this equality holds.
  • (P8)
    Duality: Set $\overline{H}(t,x)=-H(t,\varphi_{H}^{t}(x))$. Then
    $\rho_{\zeta}(a,\overline{H})=-\inf\big{\{}\rho_{\zeta}(b,H)\;|\;b\in H_{*}(M;\Lambda),\;{\textstyle\prod}(b,a)\neq 0\big{\}}.$
Remark 3.3.
The properties in Proposition 3.2 follow by adapting arguments in the undeformed case, see for instance [52]. Here we only mention that the spectrality property for a strongly nondegenerate $H$ follows from [67, Corollary 1.3] straightforwardly, and then for arbitrary $H$ on a rational symplectic manifold follows from [51, Lemma 7.2], and that the second statement in the (P7) follows from Remark 2.3.
Remark 3.4.
If $(M,\omega)$ is strongly semipositive, we do not need any Kuranishi structure in the constructions of the deformed spectral invariants $\rho_{\zeta}$ and just appeal to [68, Section 3.2] without missing any property listed as above. In this case Proposition 3.2 is enough in the following content and for our applications in Section 1.1. However, if one would like to obtain Theorems 1.1-1.5 without the strongly semipositive condition constraint, then one of course will need Kuranishi structures or something similar to deal with multiple covers of spheres of negative Chern number, but at least in principle one could use constructions just at the level of Fukaya and Ono [20] rather than [21], see [68, Section 3].
Definition 3.5.
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, for any $H\in{\mathcal{H}}$, we define
$\gamma_{\zeta}(H):=\rho_{\zeta}([M],H)+\rho_{\zeta}([M],\overline{H})$
and
$\gamma_{\zeta}(\varphi):=\inf\{\gamma_{\zeta}(H)\;|\;\varphi=\varphi_{H}^{1},\;H\in{\mathcal{H}}\}.$
Remark 3.6.
According to the (P1), (P3) and (P6) as above, we see that $\gamma_{\zeta}$ satisfies
$0\leq\gamma_{\zeta}(\varphi)\leq\|\varphi\|:=\inf\{\|H\|\;|\;\varphi=\varphi_{H}^{1},\;H\in{\mathcal{H}}\}.$ (3.1)
In fact, $\gamma_{\zeta}$ gives rise to a norm on $\operatorname{Ham}(M,\omega)$, called a $\zeta$-spectral norm. This can be proven by straightforwardly adapting arguments that are well-known in the $\zeta=0$ case, see [52]. Since this fact is not explicitly used in the present paper, we leave its proof to the reader.
Spectral invariants are notoriously difficult to compute in general. However, for a $C^{2}$-small autonomous Hamiltonian $H:M\to{\mathbb{R}}$, the deformed Hamiltonian spectral invariants $\rho_{\zeta}$ of $H$ can be computed in terms of the minmax critical sector $c_{LS}$ for $H$.
Lemma 3.7.
Let $H:M\to{\mathbb{R}}$ be a $C^{2}$-small smooth function (i.e., the $C^{2}$-norm $\|H\|_{C^{2}}$ is sufficiently small). Then $\rho_{\zeta}(a,H)=c_{LS}(a,H)$ for all $a\in H_{*}(M,{\mathbb{C}})\setminus\{0\}$.
Proof.
First we notice that for a closed manifold $M$ with positive dimension, the set of smooth functions on $M$ is dense in the space of Morse functions in $C^{2}$-topology. One may perturb $H$ to a $C^{2}$-small smooth Morse function in $C^{2}$-norm, and then further to an autonomous Hamiltonian whose critical points miss the images of the maps $f_{i}:\Delta_{i}\to M$ as in (2.6), i.e., a strongly nondegenerate autonomous Hamiltonian $H$. Since $c_{LS}(a,H)$ and $\rho_{\zeta}(a,H)$ are both Lipschitz continuous in $H$, it suffices to prove the lemma for a $C^{2}$-small Morse and strongly nondegenerate $H$.
Since $H$ is a $C^{2}$-small Morse function, the set ${\rm Per}(H)$ indeed consists of critical points of $H$ (cf. [62, Section 7]). Note that for each $q\in\operatorname{Crit}(H)$, the set $\pi_{2}(q)$ consisting of the homotopy classes of discs with boundary mapping to $q$ is isomorphic to $\pi_{2}(M)$. Given $A\in\pi_{2}(M)$, let $A_{q}\in\pi_{2}(q)$ be the element corresponding to $A$. According to our convention for Conley-Zehnder index of Hamiltonian periodic orbits, we have $\mu(A_{q})=n-\operatorname{ind}_{H}(q)+2\langle c_{1}(TM),A\rangle$. For $q\in\operatorname{Crit}(H)$, we denote by $[q,0]$ the class of the constant periodic orbit $q$ with constant capping. Clearly,
$\mathcal{A}_{H}([q,0])=H(q).$ (3.2)
Given $q\in\operatorname{Crit}(H)$, recall that the PSS map $\psi_{\zeta,H}^{\rm PSS}:CM(H;\Lambda)\to CF_{*}(H)$ defined as in Section 2.5 is
$\psi^{\operatorname{PSS}}_{\zeta,H}q=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}p\in{\rm Crit}(H),\\ A\in\pi_{2}(M)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \operatorname{ind}_{f}(q)-\operatorname{ind}_{f}(p)+2c_{1}(A)=\delta(I)\end{subarray}}|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|\exp\big{(}\int_{A}\theta\big{)}a_{I}[p,A_{p}].$ (3.3)
Here $\mathsf{s}_{\operatorname{PSS}}$ is the multisection of the Kuranishi structure with corners on the fiber product
$\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})=\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)_{ev_{-\infty},ev_{1},\ldots,ev_{k}}\times_{\iota_{q},f_{i_{1}},\dots,f_{i_{k}}}\big{(}W^{u}(q)\times\Delta_{i_{i}}\times\cdots\times\Delta_{i_{k}}\big{)}$
where ¯$\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)$ is the compactification of the space of perturbed holomorphic planes with $k$ marked points representing $A$ and asymptotic to $p$.
Now we may choose a time-independent almost complex structure $J$ compatible with $\omega$ such that the pair $(H,\rho_{J})$ is a Morse-Smale pair where $\rho_{J}(\cdot,\cdot)=\omega(\cdot,J\cdot)$, see Theorem 8.1 in [62]. Then we choose the two-parameter family $\{J_{s}\}_{s\in[0,1]}$ of almost complex structures as in Section 2.5 such that $J_{s}\equiv J$. In this case, we will show that the corresponding PSS map
$\psi_{\zeta,H}^{\rm PSS}:CM(H;\Lambda)\to CF_{*}(H)$
sends $q$ to $[q,0]$.
A typical element of ¯$\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)$ consists of a disc component and possibly some other components, each of which is cylindrical solution to the Floer equation for $H$ or a $J$-holomorphic sphere.
The space $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})$ has an extra $S^{1}$-symmetry by the $S^{1}$ action of disc, cylinder components of the domain ×${\mathbb{R}}\times S^{1}$, and by rotating the spherical components. Following Fukaya and Ono [20, Mainlemma 22.4], we see that the fixed locus $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ of this action is isolated, and away from the fixed locus the action is locally free. So, after taking a quotient by this $S^{1}$ action, one can construct a Kuranishi structure of dimension $-1$ on the quotient of the complement of the fixed locus and choose a multisection on it which is transversal to 00. Then we lift this multisection to the complement $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})\setminus\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ such that the quantity $|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|$ is given by contributions only coming from $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$. Note that the $S^{1}$-action on the spherical components is locally free except those corresponding to a topological trivial sphere, and that each element in $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ is independent of the $S^{1}$-coordinate, and hence, a solution $\gamma:[0,\infty)\to M$ of
$\dot{\gamma}(s)=-\chi(s)\;{\rm grad}_{\rho_{J}}H(\gamma(s))\quad\hbox{with}\;\gamma(0)\in W^{u}(q;H)$ (3.4)
(corresponding to the disc component), and possibly some number of negative gradient flowlines of $H$ (corresponding to the cylindrical components), or constant maps $u:S^{2}\to M$ with image $\gamma(0)$, where all subject to suitable incidence conditions. This implies that $A=0$ since the nonspherical components are $S^{1}$-independent, and that there will be no cylindrical components since their images in $M$ are $1$-dimensional objects which do not satisfy nontrivial incidence conditions. So the only contributions to $\psi_{\zeta,H}^{\rm PSS}(q)$ come from the solutions $\gamma$ of (3.4) with $\gamma(\infty)=p\in\operatorname{Crit}(H)$ with incidence conditions, corresponding to the points $\gamma(0)\in W^{u}(q)\cap W^{s}(p)$ which meet the cycles $\beta_{1}^{I},\dots,\beta_{k}^{I}$. Furthermore, we may ask that these later intersections are mutually transverse to each other and to the $W^{u}(q)$ and $W^{s}(p)$, then a dimension counting yields
$\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=\dim\big{(}W^{u}(q)\cap W^{s}(p)\big{)}=\sum_{j=1}^{k}(2n-2d(i_{j})).$
Consequently, from the fact that the only terms in (3.3) corresponding to $A=0$ satisfy
$\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=\delta(I)=\sum_{j=1}^{k}(2n-2d(i_{j}))-2k,$
we deduce that the only contributions to $\psi_{\zeta,H}^{\rm PSS}(q)$ have $k=0$. So $\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=0$, and hence (3.3) simplifies to
$\psi_{\zeta,H}^{\rm PSS}(q)=\sum_{\begin{subarray}{c}p\in\operatorname{Crit}(H),\\ \operatorname{ind}_{H}(q)=\operatorname{ind}_{H}(p)\end{subarray}}n(q,p)\cdot[p,0]$
where $n(q,p)$ is the number of points in $W^{u}(q)\cap W^{s}(p)$. If $p\neq q$, we must have $n(p,q)=0$, otherwise, there is a ${\mathbb{R}}$-action on $W^{u}(q)\cap W^{s}(p)$ which contradicts to $\dim(W^{u}(q)\cap W^{s}(p))=0$. If $p=q$, then we have $n(q,p)=1$ obviously. Therefore, $\psi_{\zeta,H}^{\rm PSS}$ sends each critical point $q$ of $H$ to $[q,0]$. This, together with (3.2), yields the desired equality. ∎
Our next lemma gives a property of the spectral invariant $\rho_{\zeta}$ coming from the non-Archimedean geometry which is very important in the following content.
Lemma 3.8.
Let $H\in{\mathcal{H}}$. If $\alpha=\lambda[M]+\beta$ with $0\neq\lambda\in\Lambda$ and $\beta\in H_{*<2n}(M;\Lambda)$, then
$\rho_{\zeta}(\alpha,H)\geq\rho_{\zeta}([M],H)++\nu(\lambda).$ (3.5)
Proof.
For proving (3.5), we may assume that $H:S^{1}\times M\to{\mathbb{R}}$ is strongly nondegenerate because of the continuity property of $\rho_{\zeta}(a,H)$ with respect to $H$. By Proposition 3.2 (P2), there exists a cycle $c\in CF_{*}(H)$ such that
$\rho_{\zeta}([M],H)=\ell_{H}(c)\quad\hbox{and}\quad\Psi^{\rm PSS}_{\zeta,H}([M])=[c].$
Similarly, if $\beta\neq 0$, then we have a cycle $b\in CF_{*}(H)$ satisfying
$\rho_{\zeta}(\beta,H)=\ell_{H}(b)\quad\hbox{and}\quad\Psi^{\rm PSS}_{\zeta,H}(\beta)=[b].$
As an orthogonalizable $\Lambda$-space $(CF_{*}(H),\ell_{H})$, the image of the boundary map $\partial_{\zeta}^{H,J}$, denoted by $\operatorname{im}(\partial_{\zeta}^{H,J})$, is still an orthogonalizable non-Archimedean normed vector space over $\Lambda$. Take a non-Archimedean orthogonal complement $V$ of $\operatorname{im}(\partial_{\zeta}^{H,J})$ in $CF_{*}(H)$. Consider now the linear map (over $\Lambda$)
$\partial_{\zeta}^{H,J}:V\longrightarrow\operatorname{im}(\partial_{\zeta}^{H,J}).$
Applying Lemma 2.6 to the above map, we see that $CF_{*}(H)$ has a non-Archimedean orthogonal basis
$(x_{0},\ldots,x_{k},y_{1},\ldots,y_{l},z_{1},\ldots,z_{l})$
satisfying the following:
  • $x_{0}=c$;
  • the classes $[x_{0}],\dots,[x_{k}]$ constitute a basis for $HF_{\zeta}(H)$;
  • $\partial_{\zeta}^{H,J}(y_{i})=z_{i}$ for $i=1,\ldots,l$, and $(z_{1},\ldots,z_{l})$ is an orthogonal basis for $\operatorname{im}(\partial_{\zeta}^{H,J})$.
Now we take any cycle $a\in CF_{*}(H)$ such that
$[a]=\Psi^{\rm PSS}_{\zeta,H}(\alpha)=\lambda[c]+[b],$
Then $a-\lambda c-b\in\operatorname{im}(\partial_{\zeta}^{H,J})$. Note that the coefficient of $[x_{0}]$ in the linear expansion of vector $[b]$ with respect to the basis $[x_{0}],\dots,[x_{k}]$ is zero because, otherwise, the preimage $\beta=(\Psi^{\rm PSS}_{\zeta,H})^{-1}[b]$ would contain a multiple of the fundamental class while by our assumption $\beta\in H_{*<2n}(M;\Lambda)$. So, under the above basis of $CF_{*}(H)$, one may expand $a$ linearly as
$a=\lambda c+\sum_{i=1}^{k}f_{i}x_{i}+\sum_{j=1}^{l}g_{j}z_{j}$
where $f_{i},g_{j}\in\Lambda$ for $i\in\{1,\ldots,k\},j\in\{1,\ldots,l\}$. By the non-Archimedean orthogonality of $(x_{0},\ldots,x_{k},y_{1},\ldots,y_{l},z_{1},\ldots,z_{l})$, we obtain
$\ell_{H}(a)=\max\limits_{i,j}\big{\{}\ell_{H}(\lambda c),\ell_{H}(f_{i}x_{i}),\ell_{H}(g_{j}z_{j})\big{\}}\geq\ell_{H}(\lambda c)=\rho_{\zeta}([M],H)+\nu(\lambda)$
from which we conclude that $\rho_{\zeta}(\alpha,H)\geq\rho_{\zeta}(\lambda[M],H)$.∎
Proposition 3.9.
$\rho_{\zeta}([pt],\overline{H})=-\rho_{\zeta}([M],H)$ for any $H\in{\mathcal{H}}$.
Proof.
We may assume that $H\in{\mathcal{H}}$ is strongly nondegenerate due to the continuity property of $\rho_{\zeta}(a,H)$ with respect to $H$. First, Proposition 3.2 (P8) yields
$\rho_{\zeta}([pt],\overline{H})=-\inf\big{\{}\rho_{\zeta}(\alpha,H)\;\big{|}\;{\alpha\in H_{*}(M;\Lambda)\setminus\{0\},\;\textstyle\prod}(\alpha,[pt])\neq 0\big{\}}.$
Next, our goal is to prove that for $\alpha\in H_{*}(M;\Lambda)$, if ${\textstyle\prod}(\alpha,[pt])\neq 0$, then
$\rho_{\zeta}(\alpha,H)\geq\rho_{\zeta}([M],H).$ (3.6)
[…]
$\alpha=(w_{0}+w_{1}\lambda)[M]+\beta$
where $w_{0},w_{1}\in{\mathbb{C}}$ with $w_{0}\neq 0$, $\lambda\in\Lambda$ with $\nu(\lambda)\neq 0$, and $\beta\in H_{*<2n}(M;\Lambda)$. Note that $\rho_{\zeta}(\lambda[M],H)=\rho_{\zeta}([M],H)+\nu(\lambda)\neq\rho_{\zeta}([M],H)$. Hence, by Proposition 3.2 (P5) and (P7), arguing in two cases: (1) $w_{1}=0$; (2) $w_{1}\neq 0$, we obtain
$\rho_{\zeta}((w_{0}+w_{1}\lambda)[M],H)=\max\big{\{}\rho_{\zeta}(w_{0}[M],H),\rho_{\zeta}(w_{1}\lambda[M],H)\big{\}}\geq\rho_{\zeta}([M],H).$ (3.7)
According to Lemma 3.8, we see that
$\rho_{\zeta}(\alpha,H)\geq\rho_{\zeta}((w_{0}+w_{1}\lambda)[M],H).$
Combining this with (3.7), we arrive at (3.6). Since one can take $\alpha$ to be $[M]$, we conclude that $\rho_{\zeta}([pt],\overline{H})=-\rho_{\zeta}([M],H)$ for strongly nondegenerate $H$, and hence for all $H\in{\mathcal{H}}$. ∎
3.3 Hamiltonian Ljusternik–Schnirelman inequality
As we mentioned before, as a vector space over ${\mathbb{C}}$ the big quantum homology $H_{*}(M;\Lambda)$ is isomorphic to $H_{*}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda$. We set
$\widehat{H}_{*}(M;\Lambda)=H_{*<2n}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda.$
Theorem 3.1.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold. Let $H\in{\mathcal{H}}$, and let $\alpha,\beta\in H_{*}(M;\Lambda)$ with $\alpha\neq 0$. Then we have
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)\leq\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta).$
Furthermore, if $\beta\in\widehat{H}_{*}(M,\Lambda)$ is non-zero, then either
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)<\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta)$
or the set ${\rm Fix}(\varphi_{H}^{1})$ is homologically non-trivial in $M$, hence it is an infinite set.
Next we postpone the proof of Theorem 3.1 to the end of this section, and employ this theorem to give the following corollary which will be used in Section 5.
Corollary 3.10.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold. Suppose that there exist $k$ homology classes $u_{i}\in\widehat{H}_{*}(M;\Lambda)$, $i=1,\ldots,k$ and $\beta\in H_{*}(M;\Lambda)$ such that $\beta*_{\zeta}u_{1}*_{\zeta}\cdots*_{\zeta}u_{k}\neq 0$ and $I_{\nu}(u_{i})\leq 0$. Let $H\in{\mathcal{H}}$ such that the ${\rm Fix}(\varphi_{H}^{1})$ is a finite set. Then the action functional $\mathcal{A}_{H}$ has $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ such that
$\rho_{\zeta}(\alpha_{k},H)=\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0})$
where $\alpha_{0}=\beta$ and $\alpha_{i}=\alpha_{i-1}*_{\zeta}u_{i}$ for $i\in\{1,\ldots,k\}$.
Proof.
The condition that $\beta*_{\zeta}u_{1}*_{\zeta}\cdots*_{\zeta}u_{k}\neq 0$ implies that each $\alpha_{i}$ is nonzero. Since ${\rm Fix}(\varphi_{H}^{1})$ is finite and $I_{\nu}(u_{i})\leq 0$ for any $i\in\{1,\ldots,k\}$, by Theorem 3.1 we obtain immediately
$\rho_{\zeta}(\alpha_{i},H)=\rho_{\zeta}(\alpha_{i-1}*_{\zeta}u_{i},H)<\rho_{\zeta}(\alpha_{i-1},H)+I_{\nu}(u_{i})\leq\rho_{\zeta}(\alpha_{i-1},H).$
So we get
$\rho_{\zeta}(\alpha_{k},H)<\ldots<\rho_{\zeta}(\alpha_{0},H).$
Combining this with Proposition 3.2 (P2) yields $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ as required. ∎
Proof of Theorem 3.1. The argument we will make to show the theorem is an adaption of Howard [35], see also [60, 26, 7, 28]. The idea of the proof is to use all properties of the spectral invariant $\rho_{\zeta}$ listed as above. First, by Proposition 3.2 (P6),
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*0)\leq\rho_{\zeta}(\alpha,H)+\rho_{\zeta}(\beta,0).$
This, together with by Proposition 3.2  (P1) and (P3), implies
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*0)\leq\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta)$
[…]
Next, we prove the second statement. It suffices to show that if $\alpha\in H_{*}(M;\Lambda)$, $\beta\in\hat{H}(M;\Lambda)$ are two nonzero classes with the property
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)=\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta),$ (3.8)
then ${\rm Fix}(\varphi_{H}^{1})$ is homologically nontrivial in $M$. For this end, we pick any open neighborhood $V$ of ${\rm Fix}(\varphi_{H}^{1})$. We may assume that $M\setminus\overline{V}$ is nonempty since nothing needs to be proved if ${\rm Fix}(\varphi_{H}^{1})=M$. We shall prove that $\overline{V}$ is homologically nontrivial.
Pick a $C^{2}$-small function $f:M\to{\mathbb{R}}$ such that $f|_{\overline{V}}\equiv 0$ and $f|_{M\setminus\overline{V}}<0$. We shall prove the following
Claim 3.11.
For sufficiently small $\varepsilon>0$, we have
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*\varepsilon f)=\rho_{\zeta}(\alpha*_{\zeta}\beta,H).$
Proof of Claim 3.11. Consider now a family of spectral invariants $\rho_{\zeta}(\alpha*_{\zeta}\beta,H*s\varepsilon f)$ for $s\in[0,1]$. Here to guarantee that the concatenated Hamiltonians $H*s\varepsilon f$ make sense, one can assume that for some small $\epsilon>0$, $H=f\equiv 0$ for $|t|<\epsilon$ by the reparameterization as before (cf. Section 2.4.5). Since $\rho_{\zeta}(\alpha*_{\zeta}\beta,H)$ is continuous in $H$ in $C^{0}$-topology and the action spectra $\operatorname{Spec}(H*s\varepsilon f)$ have measure zero for all $s$, we only need to prove that for $\varepsilon>0$ small enough, the sets $\operatorname{Spec}(H*s\varepsilon f)$ are the same for all $s\in[0,1]$. Note that
${\rm Fix}(\varphi^{1}_{H*\varepsilon f})={\rm Fix}(\varphi_{H}^{1})$
provided that $\varepsilon$ is sufficiently small. In fact, we have $\varphi^{1}_{H*\varepsilon f}=\varphi^{1}_{H}\circ\varphi^{1}_{\varepsilon f}$ and $\varphi^{t}_{\varepsilon f}|_{V}\equiv Id$. Clearly, ${\rm Fix}(\varphi^{1}_{H*\varepsilon f})\supseteq{\rm Fix}(\varphi_{H}^{1})$. Suppose now that $p\in M$ is a fixed point of $\varphi^{1}_{H*\varepsilon f}$, and hence $\varphi^{1}_{\varepsilon f}(p)=\varphi^{-1}_{H}(p)$. Since the set $M\setminus V$ is compact and $q\neq\varphi_{H}^{-1}(q)$ for any $q\in M\setminus V$ (remind that $V$ contains ${\rm Fix}(\varphi_{H}^{1})$), one may choose $\varepsilon>0$ small enough such that $\varphi_{\varepsilon f}^{1}(q)=\varphi_{f}^{\varepsilon}(q)\neq\varphi_{H}^{-1}(q)$ for any $q\in M\setminus V$. Therefore, $p\in V$, and hence $p=\varphi_{H}^{-1}(p)$. So ${\rm Fix}(\varphi^{1}_{H*\varepsilon f})\subseteq{\rm Fix}(\varphi_{H}^{1})$.
We now prove that $\operatorname{Spec}(H*\varepsilon f)=\operatorname{Spec}(H)$ for any given small $\varepsilon>0$. For each $x\in{\rm Fix}(\varphi_{H}^{1})$, we let $x(t)=\varphi_{H}^{t}(x),t\in[0,1]$ and $\overline{x}:D^{2}\to M$ be a cotinuous disc such that $\overline{x}(e^{2\pi it})=x(t)$. It is easy to see that
$\varphi_{H*\varepsilon f}^{t}(x)=\begin{cases}x&0\leq t\leq 1/2,\\ \varphi_{H}^{2t-1}(x)&1/2\leq t\leq 1.\\ \end{cases}$
Corresponding to each critical point $(x,\overline{x})$ of $\mathcal{A}_{H}$, we denote by $x^{\prime}(t)=\varphi_{H*\varepsilon f}^{t}(x)$ and let $\overline{x}^{\prime}:D^{2}\to M$ be a disc given by
$\overline{x}^{\prime}(re^{2\pi it})=\begin{cases}\overline{x}(r)&(r,t)\in[0,1]\times[0,1/2],\\ \overline{x}(re^{2\pi it})&(r,t)\in[0,1]\times[1/2,1].\\ \end{cases}$
Clearly, on the boundary of this disc, $\overline{x}^{\prime}$ restricts as the periodic orbit $x^{\prime}$. So we get a bijection
$\mathcal{J}:\operatorname{Crit}(\mathcal{A}_{H})\longrightarrow\operatorname{Crit}(\mathcal{A}_{H*\varepsilon f}),\quad(x,\overline{x})\mapsto(x^{\prime},\overline{x}^{\prime}).$
A straightforward computation yields $\mathcal{A}_{H*\varepsilon f}(x,\overline{x})=\mathcal{A}_{H}(\mathcal{J}(x^{\prime},\overline{x}^{\prime}))$. This will result in $\operatorname{Spec}(H*\varepsilon f)=\operatorname{Spec}(H)$. Similarly, one can show that $\operatorname{Spec}(H*s\varepsilon f)=\operatorname{Spec}(H)$ for all $s$. This completes the proof of the claim.∎
Now we are in a position to finish the proof. By our assumption, for $\beta\in\widehat{H}(M;\Lambda)$, we may write $\beta=\sum_{i}\beta_{i}\otimes k_{i}$ with $0\neq\beta_{i}\in H_{*<2n}(M;{\mathbb{C}})$ and $k_{i}\in\Lambda$. By Proposition 3.2 (P1), (P4) and (P7), for given sufficiently small $\varepsilon>0$, the following inequalities hold:
$\rho_{\zeta}(\beta,\varepsilon f)$$\leq$ $\max_{i}\big{\{}\rho_{\zeta}(\beta_{i}\otimes k_{i},\varepsilon f)\big{\}}$ (3.9)
$=$ $\max_{i}\big{\{}\rho_{\zeta}(\beta_{i},\varepsilon f)+\nu(k_{i})\big{\}}$
$\leq$ $\max_{i}\big{\{}\rho_{\zeta}(\beta_{i},\varepsilon f)\big{\}}+I_{\nu}(\beta).$
Then it follows from Proposition 3.2 (P6) and Claim 3.11 yields
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)$ $=$ $\rho_{\zeta}(\alpha*_{\zeta}\beta,H*\varepsilon f)$ (3.10)
$\leq$ $\rho_{\zeta}(\alpha,H)+\rho_{\zeta}(\beta,\varepsilon f)$
$\leq$ $\rho_{\zeta}(\alpha,H)+\max_{i}\big{\{}\rho_{\zeta}(\beta_{i},\varepsilon f)\big{\}}+I_{\nu}(\beta),$
where in the last inequality we have used (3.9). On the one hand, from (3.10) we deduce that there exists $i_{0}\in{\mathbb{N}}$ such that
$\rho_{\zeta}(\beta_{i_{0}},\varepsilon f)\geq\rho_{\zeta}(\alpha*_{\zeta}\beta,H)-\rho_{\zeta}(\alpha,H)-I_{\nu}(\beta).$
Then, by (3.8) and Lemma 3.7, we get
$c_{LS}(\beta_{i_{0}},\varepsilon f)=\rho_{\zeta}(\beta_{i_{0}},\varepsilon f)\geq 0.$
On the other hand, it follows from Proposition 3.1.3 that
$c_{LS}(\beta_{i_{0}},\varepsilon f)=c_{LS}([M]\cap\beta_{i_{0}},\varepsilon f)\leq c_{LS}([M],\varepsilon f)=\max\varepsilon f=0.$
Consequently, we obtain
$c_{LS}([M]\cap\beta_{i_{0}},\varepsilon f)=c_{LS}([M],\varepsilon f)=0$
for some nonzero class $\beta_{i_{0}}\in H_{*<2n}(M;\Lambda)$. Then by Proposition 3.1.4, the zero level set $\overline{V}$ of $f$ is homologically nontrivial. This completes the proof of the second statement. ∎
4 Nonvanishing Gromov-Witten invariants
We begin with the following lemma which plays a key role in reproving a remarkable result by Lu [44, Corollary 1.19], namely, bounding Hofer-Zehnder capacity from above in terms of Gromov-Witten invariants.
Lemma 4.1 ([68, Lemma 4.2]).
Let $(M^{2n},\omega)$ be a closed symplectic manifold which admits a nonzero Gromov-Witten invariant of the form:
${\rm GW}_{0,k+3,A}\big{(}[pt],a_{0},[pt],a_{1},\ldots,a_{k}\big{)}$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$ and $a_{0},\ldots,a_{k}$ are rational homology classes of even degree. Then, there is an open dense set of possible choices of the deformation parameter $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ and a homology class $a_{0}^{\prime}\in H_{<2n}(M;{\mathbb{Q}})$ such that
$[pt]*_{\zeta}a_{0}^{\prime}=\lambda[M]+\beta\quad\hbox{and}\quad\nu(\lambda)\geq-\langle[\omega],A\rangle$
where $\lambda\in\Lambda$ and $\beta\in H_{*<2n}(M;\Lambda_{0})$.
Corollary 4.2.
Under the same conditions in Lemma 4.1, for any $H\in{\mathcal{H}}$, we have
$\rho_{\zeta}([M],H)-\rho_{\zeta}([pt],H)\leq\langle[\omega],A\rangle.$ (4.1)
Moreover, the strict inequality holds if the number of elements in #$\#{\rm Fix}(\varphi_{H}^{1})$ is finite.
Remark 4.3.
Combining Theorems 3.9 and Corollary 4.2, one finds that under the same conditions in Lemma 4.1, it holds that
$\gamma_{\zeta}(\varphi)<\langle[\omega],A\rangle$
for any $\varphi\in\operatorname{Ham}(M,\omega)$ with finite #$\#{\rm Fix}(\varphi)$.
Proof of Corollary 4.2. According to Lemma 4.1, one has $[pt]*_{\zeta}a_{0}^{\prime}=\lambda[M]+\beta$ and $\nu(\lambda)\geq-\langle[\omega],A\rangle$ for some $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$, $a_{0}^{\prime}\in H_{<2n}(M;{\mathbb{Q}})$, $\lambda\in\Lambda$ and $\beta\in H_{*<2n}(M;\Lambda)$. We put $\alpha=[pt]*_{\zeta}a_{0}^{\prime}$. It follows from Theorem 3.1 that if the number of ${\rm Fix}(\varphi_{H}^{1})$ is finite, then
$\rho_{\zeta}(\alpha,H)<\rho_{\zeta}([pt],H)+I_{\nu}(a_{0}^{\prime})=\rho_{\zeta}([pt],H).$ (4.2)
Then, Lemma 3.8, together with (4.2), would give
$\rho_{\zeta}([M],H)-\langle[\omega],A\rangle\leq\rho_{\zeta}([M],H)+\nu(\lambda)\leq\rho_{\zeta}(\alpha,H)<\rho_{\zeta}([pt],H),$
from which we obtain the strict inequality as required. Hence, for any $H\in{\mathcal{H}}$, (4.1) follows from this inequality for strongly nondegenerate $H$ and the continuity property of $\rho_{\zeta}$. ∎
Lemma 4.4.
Let $(M^{2n},\omega)$ be a closed symplectic manifold which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a,\ldots,a,b_{0},\ldots,b_{k})$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, the class $a\in H_{*<2n}(M;{\mathbb{Q}})$ occurs $l$ times with $l>2n$, and $b_{0},\ldots,b_{k}$ ($0\leq k\in{\mathbb{Z}}$) are rational homology classes of even degree and $b_{0}\in H_{*<2n}(M;{\mathbb{Q}})$. Then there is an open set of possible choice of deformation parameter $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that $(\underbrace{a*_{\zeta},\ldots,*_{\zeta}a}_{l})_{\langle[\omega],A\rangle}\neq 0$, and in particular the $l$-fold product of $a$ w.r.t. $*_{\zeta}$ is nonzero.
Proof.
The argument what we will do is essentially due to [68, Lemma 4.2] (see also [44, Lemma 7.1]), and we include it for completeness. Note first that the class $A$ appearing in the mixed Gromov-Witten invariant cannot be zero. Indeed, the fundamental class axiom of Gromov-Witten invariants (cf. [48, Proposition 7.5.7]) implies that none of the classes $b_{1},\ldots,b_{k}$ can be $\lambda[M]$ for some $\lambda\in{\mathbb{Q}}\setminus\{0\}$. Assume now that $A=0$ in $H_{2}(M;\mathbb{Z})/{\rm torsion}$. Using the dimension formula (2.2), one finds that
$l(2n-\deg(a))+2n-\deg(b_{0})+\sum_{j=1}^{k}(2n-2-\deg(b_{j}))=2n$
contradicting to our assumption that $l>2n$ and $2n-\deg(a)\geq 1$. So $A\neq 0$ in $H_{2}(M;\mathbb{Z})/{\rm torsion}$. By the divisor axiom of the mixed Gromov-Witten invariants, one may assume that $b_{1},\ldots,b_{k}\in\oplus_{k=0}^{n-2}H_{2i}(M;{\mathbb{Q}})$. We now take a homogeneous basis $\xi_{1},\ldots,\xi_{N}$ for the ${\mathbb{Z}}$-module $\oplus_{i=0}^{n-1}H_{2i}(M;\mathbb{Z})/{\rm torsion}$ such that $\xi_{1},\ldots,\xi_{s}$ is a basis for $H_{2n-2}(M;\mathbb{Z})/{\rm torsion}$ for some $s<N$. Then by the multilinearity and symmetry properties of the Gromov-Witten invariants, we may further assume that our nonzero Gromov-Witten invariant is of the shape
${\rm GW}_{0,l+1+\sum_{j}\alpha_{j},A}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b_{0}^{\prime},\underbrace{\xi_{s+1},\ldots\xi_{s+1}}_{\alpha_{s+1}},\ldots,\underbrace{\xi_{N},\ldots,\xi_{N}}_{\alpha_{N}})$ (4.3)
for some homogeneous class $b_{0}^{\prime}\in\oplus_{i=1}^{n-1}H_{2i}(M;{\mathbb{Q}})$ and $\alpha=(\alpha_{s+1},\ldots,\alpha_{N})\in{\mathbb{N}}^{N-s}$ with $\sum_{j=s+1}^{N}\alpha_{j}\leq k$.
For $\vec{x}=(x_{1},\ldots,x_{s})\in{\mathbb{C}}^{s}$ and $\vec{y}=(y_{s+1},\ldots,y_{N})\in{\mathbb{C}}^{N-s}$, we put
$\zeta(x,y)=\sum_{i=1}^{s}x_{i}\xi_{i}+\sum_{j=s+1}^{N}y_{j}\xi_{j}.$
Let us consider the class $\underbrace{a*_{\zeta},\ldots,*_{\zeta}a}_{l}$ as a function of $(\vec{x},\vec{y})$. Take a homogeneous basis $\{c_{j}\}_{j=1}^{K}$ for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}_{j=1}^{K}$ with respect to the Poincáre homology intersection pairing $\circ$ such that $c_{1}=b_{0}^{\prime}$. According to the symmetry properties of the Gromov-Witten invariants, after expanding out the formula for $a^{*_{\zeta}l}$, we see that the coefficient on the class $c^{1}$ is an expression of the form
$\sum_{g\in\Gamma_{\omega}}\sum_{\beta=(\beta_{s+1},\ldots,\beta_{N})}f_{g,\beta}(\vec{x})y_{s+1}^{\beta_{s+1}}\cdots y_{N}^{\beta_{N}}T^{g}$
where $f_{g,\beta}(\vec{x})$ is a function of $\vec{x}$ which is identically zero for all but finitely many $\beta$ (by the dimension formula (2.2)). In particular, for the period $g=\langle[\omega],A\rangle$ and the value $\beta=\alpha$, we obtain
[…]
$=$ [$\sum_{\stackrel{{\scriptstyle B\in H_{2}(M;\mathbb{Z})/{\rm torsion},}}{{\langle[\omega],B\rangle=\langle[\omega],A\rangle}}}\bigg{(}{\rm GW}_{0,l+1+\sum\alpha_{j},B}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b_{0}^{\prime},\underbrace{\xi_{s+1},\ldots\xi_{s+1}}_{\alpha_{s+1}},\ldots,\underbrace{\xi_{N},\ldots,\xi_{N}}_{\alpha_{N}})$
×$\hskip 284.52756pt\times\bigg{(}\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap B}\bigg{)}\bigg{)}.$
We notice that the coefficient on $\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap A}$ in the above sum is nonzero because of our assumed nonzero Gromov-Witten (4.3). Now $H_{2n-2}(M;\mathbb{Z})/{\rm torsion}$ has $\{\xi_{i}\}_{i=1}^{s}$ as a basis, the map $B\mapsto(\xi_{1}\circ B,\ldots,\xi_{s}\circ B)$ is injective. Consequently, the coefficient on $T^{\langle[\omega],A\rangle}c^{1}$ in the expansion of $a^{*_{\zeta}l}$ is a polynomial in $e^{x_{i}}$ and $y_{j}$ which carries a nonzero coefficient multiplying
$\bigg{(}\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap A}\bigg{)}\prod_{j=s+1}^{N}y_{j}^{\alpha_{j}}.$
Then it follows from the fact that this polynomial does not vanish identically that there is an open dense set of choices of $(\vec{x},\vec{y})$ at which the coefficient on $T^{\langle[\omega],A\rangle}c^{1}$ is not zero. This proves that $(\underbrace{a*_{\zeta(\vec{x},\vec{y})},\ldots,*_{\zeta(\vec{x},\vec{y})}a}_{l})_{\langle[\omega],A\rangle}\neq 0$ for such choices of $(\vec{x},\vec{y})$. ∎
Before finishing this section, we notice that the argument what we have already done in the proof of Lemma 4.4 can be used to obtain the following
Lemma 4.5.
Let $(M^{2n},\omega)$ be a closed symplectic manifold which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],b_{1},\ldots,b_{k})$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{Q}})$ with $2\leq l\in{\mathbb{N}}$, and the classes $b_{1},\ldots,b_{k}$ are rational homology classes of even degree. Then there is an open set of possible choice of deformation parameter $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that the $l$-fold product $a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}=\lambda[M]+\beta$ with $\beta\in H_{*<2n}(M;\Lambda_{0})$ satisfies $\nu(\lambda)\geq-\langle[\omega],A\rangle$.
Sketch of the proof of Lemma 4.5. We first need to show that the class $A$ appearing in our nonzero Gromov-Witten invariant is nonzero. According to the fundamental class axiom of mixed Gromov-Witten invariants (see [48, Proposition 7.5.7]), none of the classes $b_{1},\ldots,b_{k}$ can be a multiple of the fundamental class. Then the dimension formula (2.2) yields
$\sum_{i=1}^{l}(2n-\deg(a_{i}))+2n+\sum_{j=1}^{k}(2n-2-\deg(b_{j}))=2n$
contradicting to our assumption that $2n-\deg(a_{i})\geq 1$ for all $i$. The remaining argument of the proof is the same as that in Lemma 4.4, we omit it. ∎
5 Proof of Theorems 1.1-1.5
Proof of Theorem 1.1. Given $\varphi\in\operatorname{Ham}(M,\omega)$, take $H\in{\mathcal{H}}$ such that $\varphi=\varphi^{1}_{H}$. Without loss of generality, we may assume that the number of elements in ${\rm Fix}(\varphi_{H}^{1})$ is finite, otherwise, nothing needs to be proved. Assume that there are $A\in H_{2}(M;\mathbb{Z})/\hbox{torsion}$ and rational homology classes $a_{0},\ldots,a_{k}$ of even degree such that
${\rm GW}_{0,k+3,A}\big{(}[pt],a_{0},[pt],a_{1}\ldots,a_{k}\big{)}\neq 0.$ (5.1)
Our proof is divided to two steps:
Step 1. By the definition of cuplength (cf. Section 1) and the Poincaré duality, there exist $u_{i}\in H_{<2n}(M,{\mathbb{C}})$, $i=1\ldots,k$ such that $u_{1}\cap\cdots\cap u_{k}=[pt]$ and $\operatorname{cuplength}(M)=k+1$. Take classes $\alpha_{0},\dots,\alpha_{k}$ in $H_{*}(M;\Lambda)$ given by
$\alpha_{0}=[M]\quad\hbox{and}\quad\alpha_{i}=\alpha_{i-1}*_{\zeta}u_{i},\;i=1\ldots,k.$
We shall prove the strict inequality
$\rho_{\zeta}(\alpha_{0},H)-\rho_{\zeta}(\alpha_{k},H)<\langle[\omega],A\rangle.$ (5.2)
Note that the big quantum product is a quantum deformation of the homological intersection product in $H_{*}(M;{\mathbb{C}})$. In particular, we have
[…]
which implies that each $\alpha_{i}$ is nonzero, and that ${\textstyle\prod}(\alpha_{k},[M])\neq 0$. Thus, it follows from Propositions 3.2 (P8) and 3.9 that
$\rho_{\zeta}(\alpha_{k},H)\geq-\rho_{\zeta}([M],\overline{H})=\rho_{\zeta}([pt],H).$ (5.3)
Since $\operatorname{GW}_{0,k+3,A}\big{(}[pt],a_{0},[pt],a_{1}\ldots,a_{k}\big{)}\neq 0$ and $\sharp{\rm Fix}(\varphi_{H}^{1})$ is finite, by Corollary 4.2, we obtain
$\rho_{\zeta}([M],H)-\rho_{\zeta}([pt],H)<\langle[\omega],A\rangle.$
Combining this with (5.3), we get (5.2) as required.
Step 2. The next goal is to prove that
#$\#{\rm Fix}(\varphi)\geq\bigg{\lceil}\frac{p\cdot{\rm cuplength}(M;{\mathbb{C}})}{\langle[\omega],A\rangle}\bigg{\rceil}.$ (5.4)
According to Corollary 3.10, there exist $k+1$ elements $\widetilde{x}_{i}=(x_{i},\overline{x}_{i})\in\operatorname{Crit}(\mathcal{A}_{H})$ such that
$\rho_{\zeta}(\alpha_{k},H)=\mathcal{A}_{H}(\widetilde{x}_{k})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(\widetilde{x}_{0}).$ (5.5)
Set
=$\mathcal{E}=\big{\{}\mathcal{A}_{H}(\widetilde{x}_{i})|i=0,\ldots,k\big{\}},$
and define an equivalent relation on $\mathcal{E}$ as follows: for $a,b\in\mathcal{E}$, $a\sim b$ if and only if $a-b$ is a symplectic period, i.e., $a-b\in\Gamma_{\omega}$. Denote by $\widehat{\mathcal{E}}$ the set of such equivalence classes. Then we see that the number $\#\widehat{\mathcal{E}}$ bounds from below the number of Hamiltonian periodic orbits $x_{i}$ (corresponding to the elements in ${\rm Fix}(\varphi_{H}^{1})$).
We now estimate $\#\widehat{\mathcal{E}}$. Let $l=\langle[\omega],A\rangle/p$ (recall that $p\in\Gamma_{\omega}$ is the positive generator of $\Gamma_{\omega}$). According to (5.2) and (5.5), each equivalence class contains at most $l$ elements in $\mathcal{E}$. So we obtain the inequality (5.4) as required.
Finally, since (5.4) holds for every nonzero Gromov-Witten invariant of the form (5.1), we obtain the desired inequality. This completes the proof.∎
Proof of Theorem 1.2. Take $H\in{\mathcal{H}}$ such that $\varphi=\varphi^{1}_{H}$. Since $\gamma_{\zeta}(\varphi_{H}^{1})<p$, the theorem follows from (5.3) and (5.5). The theorem is proved. ∎
Proof of Theorem 1.3. Assume that ${{\rm qcl}_{\zeta}}(g)>0$ for some $g\in\Gamma_{\omega}$ and $\zeta\in H_{*}(M,\Lambda_{0})$. Set ${{\rm qcl}_{\zeta}}(g)=k+1$ and $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$. Then there exist homology classes $a_{1},\ldots,a_{k}\in H_{<2n}(M;{\mathbb{C}})$ such that $(\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k})_{g}\neq 0$ with $g=\langle[\omega],A\rangle$ for some $A\in H_{*}(M;{\mathbb{Z}})/{\rm torsion}$. Hence,
$\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k}=\sum_{h\in\Gamma_{\omega}}(\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k})_{h}T^{h}\neq 0.$
Set $\beta=\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k}$. Then $I_{\nu}(\beta)\geq-g$. By Proposition 3.2 (P1), we obtain that
$I_{\nu}(\beta)+\int^{1}_{0}\min H_{t}dt\leq\rho_{\zeta}(\beta,H)\quad\hbox{and}\quad\rho_{\zeta}([M],H)\leq\int^{1}_{0}\max H_{t}dt.$ (5.6)
According to Corollary 3.10, there exist $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ such that
$\rho_{\zeta}(\beta,H)=\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\rho_{\zeta}([M],H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$
This, together with (5.6), gives
$-g+\int^{1}_{0}\min H_{t}dt\leq\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\mathcal{A}_{H}(x_{0},\overline{x}_{0})\leq\int^{1}_{0}\max H_{t}dt.$ (5.7)
Similar to the proof of Theorem 5 in Step 2, we let =$\mathcal{E}=\{\mathcal{A}_{H}(x_{i},\overline{x}_{i})\;|\;i=0,\ldots,k\}$ and consider the set $\widehat{\mathcal{E}}$ of equivalence classes as before. It follows from (5.7) that each equivalence class in $\widehat{\mathcal{E}}$ contains at most
$\bigg{\lfloor}\frac{\|H\|+g}{p}\bigg{\rfloor}+1$
elements in $\mathcal{E}$. So we get
#$\#{\rm Fix}(\varphi_{H}^{1})\geq\#\widehat{\mathcal{E}}\geq\bigg{\lceil}\frac{k+1}{\big{[}(\|H\|+g)/p\big{]}+1}\bigg{\rceil}\geq\bigg{\lceil}\frac{p(k+1)}{\|H\|+g+p}\bigg{\rceil}.$
Therefore, the desired estimate for symplectic fixed points in the theorem follows from the above inequality immediately. ∎
Proof of Theorem 1.4. According to Lemma 4.4, for each integer $l>2n$, there exists a class $\zeta_{l}\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that $(\underbrace{a*_{\zeta_{l}},\ldots,*_{\zeta_{l}}a}_{l})_{g_{l}}\neq 0$ for $g_{l}=\langle[\omega],A_{l}\rangle$, and hence ${\rm qcl}_{\zeta_{l}}(g_{l})\geq l+1$. Moreover, for each such $l$, by (4.3), we have the nonzero Gromov-Witten invariant
${\rm GW}_{0,l+1+K,A_{l}}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b^{l},\eta_{1}^{l},\ldots,\eta_{K}^{l})$ (5.8)
for some homogeneous class $b^{l}\in H_{*}(M;{\mathbb{Q}})$ of even degree, and homogeneous classes $\eta_{1}^{l}\ldots,\eta_{K}^{l}\in\oplus_{i=0}^{n-2}H_{2i}(M;{\mathbb{Q}})$ with $0\leq K=K(l)\leq k(l)$ (corresponding to $K=0$ there are no inserted classes $\eta_{i}^{l}$). Then it follows from Theorem 1.3 that for any $\varphi\in\operatorname{Ham}(M,\omega)$,
#$\#{\rm Fix}(\varphi)\geq\limsup_{l}\bigg{\lceil}\frac{p(l+1)}{p+g_{l}+\|\varphi\|}\bigg{\rceil}$ (5.9)
We now estimate the right hand side of the above inequality. Since the invariant (5.8) does not vanish, according to the dimension formula (2.2), we obtain
$l(2n-\deg(a))+2n-\deg(b^{l})+\sum_{j=1}^{K}(2n-\deg(\eta^{l}_{j}))=2n+2\langle c_{1}(TM),A_{l}\rangle+2K.$ (5.10)
Note that $2n-\deg(\eta^{l}_{j})\geq 4$ for all $2n<l\in{\mathbb{N}}$ and all $j\in\{1,\ldots,K(l)\}$. Then (5.10) yields
  • (a)
    $\langle c_{1}(TM),A_{l}\rangle\to\infty$ as $l\to\infty$,
  • (b)
    $\frac{2}{l}\langle c_{1}(TM),A_{l}\rangle\leq 2n-\deg(a)+\frac{1}{\ell}\sum_{j=1}^{K}(2n-2-\deg(\eta^{l}_{j}))-\frac{1}{l}\deg(b^{l})$ for all $l$.
If the function $l\mapsto k(l)$ is bounded and so for $K=K(l)$, then the (b) as above implies
$\limsup_{l}\frac{l}{\langle c_{1}(TM),A_{l}\rangle}\geq\frac{2}{2n-\deg(a)}.$
Since $\omega|_{\pi_{2}(M)}=\lambda c_{1}|_{\pi_{2}(M)}\neq 0$ for some $\lambda>0$, we get that $\omega(A_{l})\to\infty$ as $l\to\infty$, and $p=\lambda N$. Hence,
$\small\limsup_{l}\frac{p(l+1)}{p+g_{l}+\|\varphi\|}\geq\limsup_{l}\frac{l}{\langle c_{1}(TM),A_{l}\rangle}\cdot\frac{\lambda N}{(p+\|\varphi\|)/\langle c_{1}(TM),A_{l}\rangle+\lambda}\geq\frac{2N}{2n-\deg(a)}$
from which, together with (5.9), we obtain the estimate as required. ∎
Proof of Theorem 1.5. Let $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$. The proof of the theorem is divided into two parts.
(1) Since our symplectic manifold $(M,\omega)$ carries the nonzero mixed Gromov-Witten invariant, by Lemma 4.5, there is a class $\zeta\in\oplus_{i=0}^{n-1}H(M;{\mathbb{C}})$ such that
$a_{1}*_{\zeta}\cdots*_{\zeta}a_{l}=\lambda[M]+\beta$ (5.11)
with $\beta\in H_{*<2n}(M;\Lambda_{0})$ and $\nu(\lambda)\geq-\langle[\omega],A\rangle$. Set $\alpha_{0}=[M]$ and
$\alpha_{j}=a_{1}*_{\zeta}\cdots*_{\zeta}a_{j},\quad j=1,\ldots,l.$
It follows from Corollary 3.10 that there exist elements $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{l})\in\operatorname{Crit}(\mathcal{A}_{H})$ such that
$\rho_{\zeta}(\alpha_{l},H)=\mathcal{A}_{H}(x_{l},\overline{x}_{l})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$ (5.12)
According to (5.11) and Lemma 3.8, we have
$\rho_{\zeta}(\alpha_{l},H)\geq\rho_{\zeta}([M],H)+\nu(\lambda)\geq\rho_{\zeta}([M],H)-\langle[\omega],A\rangle.$
Together with (5.12), this proves that
#$\#{\rm Fix}(\varphi^{1}_{H})\geq 1+\bigg{\lceil}\frac{l+1-\langle[\omega],A\rangle/p-1}{\langle[\omega],A\rangle/p}\bigg{\rceil}=\bigg{\lceil}\frac{pl}{\langle[\omega],A\rangle}\bigg{\rceil}.$
This completes the proof of the estimate (1.3).
(2) According to part (1), for each $l\geq 2$, we have
#$\#{\rm Fix}(\varphi^{1}_{H})\geq\bigg{\lceil}\frac{pl}{\langle[\omega],A_{l}\rangle}\bigg{\rceil}.$ (5.13)
Next, we estimate the right hand side of the above inequality. The nonvanishing of our sequence of Gromov-Witten invariants implies
$\sum_{i=1}^{l}(2n-a_{i})+2n=2n+2\langle c_{1}(TM),A\rangle$
from which we deduce that
$l\cdot\min_{i}\{2n-a_{i}\}\leq 2\langle c_{1}(TM),A\rangle\leq l\cdot\max_{i}\{2n-a_{i}\}.$ (5.14)
Hence, $\langle c_{1}(TM),A\rangle>0$ because $2n-2a_{i}\geq 1$ for all $i$ (so the right hand of the inequality (1.4) is positive). Combining (5.13) and (5.14) yields the estimate (1.4). The theorem is proven.∎
Appendix A Coadjoint orbits of compact simple Lie groups
We shall employ Theorem 1.1 to obtain symplecitc fixed point estimates for monotone coadjoint orbits of compact simple Lie groups. Recall that a compact Lie group $G$ acts on its dual Lie algebra ${\mathfrak{g}}^{*}$ by the coadjoint representation. For $\lambda\in{\mathfrak{g}}^{*}$, the coadjoint orbit ${\mathcal{O}}_{\lambda}$ passing through $\lambda$ is endowed with a canonical symplectic form $\omega_{\lambda}$, known as Kostant-Kirillov-Souriau form (cf. [38]). Let (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denote an adjoint invariant inner product on ${\mathfrak{g}}$. There is a canonical isomorphism
${\mathfrak{g}}\longrightarrow{\mathfrak{g}}^{*},\quad X\longmapsto(X,\cdot)$
which induces a scalar product on ${\mathfrak{t}}^{*}$, denoted again by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Let $T\subset G$ be a maximal torus with Lie algebra ${\mathfrak{t}}\subset{\mathfrak{g}}$. Let $\mathfrak{R}\subset{\mathfrak{t}}^{*}$ be the root system of $T$. Hence, the complexifications of ${\mathfrak{g}}$ and ${\mathfrak{t}}$ satisfy the following decomposition
${\mathfrak{g}}_{\mathbb{C}}={\mathfrak{t}}_{\mathbb{C}}\oplus\bigoplus\limits_{\alpha\in\mathfrak{R}}{\mathfrak{g}}_{\alpha}$
where ${\mathfrak{g}}_{\alpha}$ is the root space for which the root $\alpha\in\mathfrak{R}$ is given by
${\mathfrak{g}}_{\alpha}:=\big{\{}Y\in{\mathfrak{g}}_{\mathbb{C}}\;|\;[X,Y]=\sqrt{-1}\alpha(X)Y,\quad\hbox{for all}\;X\in{\mathfrak{t}}\big{\}}.$
Corresponding to each root $\alpha\in\mathfrak{R}$, one has a coroot $\check{\alpha}\in{\mathfrak{t}}$ identified with $2\alpha/(\alpha,\alpha)\in{\mathfrak{t}}^{*}$ via the above invariant inner product. Given $\alpha\in\mathfrak{R}$, we define a reflection $s_{\alpha}$ on ${\mathfrak{t}}^{*}$ as
$s_{\alpha}:{\mathfrak{t}}^{*}\longrightarrow{\mathfrak{t}}^{*},\quad\xi\longmapsto\xi-\frac{2(\alpha,\xi)}{(\alpha,\alpha)}\alpha.$
Fix a subset $\mathcal{S}\subset\mathfrak{R}$ consisting of simple roots $\alpha_{1},\ldots,\alpha_{\ell}$ such that any $\alpha\in\mathfrak{R}$ can be written as
$\alpha=\sum_{k=1}^{\ell}m_{k}\alpha_{k}$
where $m_{k}\in{\mathbb{Z}}$ are all in ${\mathbb{Z}}_{\geq 0}$ or all in ${\mathbb{Z}}_{\leq 0}$. The roots $\alpha$ for which all $m_{k}$ are in ${\mathbb{Z}}_{\geq 0}$ (reps. ${\mathbb{Z}}_{\leq 0}$) are called positive (resp. negative). Write $\mathfrak{R}=\mathfrak{R}^{+}\cup\mathfrak{R}^{-}$ according to the positive and negative roots. Obviously, we have $\mathcal{S}\subset\mathfrak{R}^{+}$.
Let $G_{\mathbb{C}}$ be the complexification of $G$. We choose a parabolic subgroup of $P\subset G_{\mathbb{C}}$ such that ${\mathcal{O}}_{\lambda}\cong G_{\mathbb{C}}/P$. This allows us to endow ${\mathcal{O}}_{\lambda}$ with a complex structure $J$ compatible with $\omega_{\lambda}$ so that the triple $({\mathcal{O}}_{\lambda},\omega_{\lambda},J)$ is a Kähler manifold. Fix a Borel subgroup $B\subset P$ which has Lie algebra
$\mathfrak{b}={\mathfrak{t}}_{\mathbb{C}}\oplus\bigoplus_{\alpha\in\mathfrak{R}^{+}}{\mathfrak{g}}_{\alpha}.$
The group $W$ of linear transformations of ${\mathfrak{t}}^{*}$ generated by all reflections $s_{\alpha}$, $\alpha\in\mathfrak{R}$ is called the Weyl group of $G$ relative to $T$. It is in fact generated by $s_{\alpha_{1}},\ldots,s_{\alpha_{\ell}}$, and is isomorphic to the quotient group $N_{G_{\mathbb{C}}}(T)/T$ where $N_{G_{\mathbb{C}}}(T)$ is the normalizer of $T$ in $G_{\mathbb{C}}$. The length $L(w)$ of an element $w$ of $W$ is the minimum number of simple reflections whose product is $w$. The element of longest length is denoted by $w_{0}$. The parabolic subgroup $P$ of $G_{\mathbb{C}}$ corresponds canonically to a subset $\mathcal{S}_{P}=\{\alpha\in\mathcal{S}|\langle\lambda,\check{\alpha}\rangle=0\}$. Let $\mathfrak{R}^{+}_{P}$ be the set of positive roots can be written as sums of roots in $\mathcal{S}_{P}$. We denote by $W_{P}$ the Weyl group of $P$ generated by the set of simple roots in $\mathcal{S}_{P}$.
For $u\in W/W_{P}$, we let $X(u)=\overline{BuP/P}$ be the corresponding Schubert variety. Here the $u$ should first be interpreted as a representative in $W$ and then a representative in the normalizer $N_{G_{\mathbb{C}}}(T)$ due to the fact that $X(u)$ is independent of these choices. This is a subvariety of $G_{\mathbb{C}}/P$ of real dimension $2L(u)$. Let $B^{-}:=w_{0}Bw_{0}\subset G_{\mathbb{C}}$ be the Borel subgroup opposite to $B$. Similarly, we let $Y(u)=\overline{B^{-}uP/P}$ be the opposite Schubert variety of codimension $2L(u)$. Denote by $\sigma_{u}$ and $\check{\sigma}_{u}$ the fundamental classes in the homology group $H_{*}(G_{\mathbb{C}}/P;{\mathbb{Z}})$ of $Y(u)$ and $X(u)$, respectively. It is well known that the set of Schubert classes $\{\sigma_{u}\}_{u\in W/W_{P}}$ forms a ${\mathbb{Z}}$-basis of $H_{*}(G_{\mathbb{C}}/P;{\mathbb{Z}})$, and the set of Schubert classes $\{\check{\sigma}_{u}\}_{u\in W/W_{P}}$ is the dual basis of $\{\sigma_{u}\}_{u\in W/W_{P}}$ with respect to the Poincaré intersection pairing.
For any $u$ in $W/W_{P}$, we let $x(u)=uP/P$ be the corresponding point in $G_{\mathbb{C}}/P$. Such points are all $T$-fixed points of $G_{\mathbb{C}}/P$. The varieties $X(u)$ and $Y(u)$ meet transversally at the point $x(u)$. Corresponding to a positive root $\alpha\in\mathfrak{R}\setminus\mathfrak{R}_{P}^{+}$, we have a unique irreducible $T$-invariant curve $C_{\alpha}=SL_{2}({\mathbb{C}})_{\alpha}\cdot P/P$ containing $1\cdot P$ and $s_{\alpha}\cdot P$, where $SL_{2}({\mathbb{C}})_{\alpha}\subset G_{\mathbb{C}}$ is the subgroup with Lie algebra ${\mathfrak{g}}_{\alpha}\oplus{\mathfrak{g}}_{-\alpha}\oplus[{\mathfrak{g}}_{\alpha},{\mathfrak{g}}_{-\alpha}]$. Such $C_{\alpha}$ has the symplectic area $\langle\lambda,\check{\alpha}\rangle$ with respect to the symplectic form $\omega_{\lambda}$. If $\alpha\in\mathcal{S}\setminus\mathcal{S}_{P}$, then $C_{\alpha}=X(s_{\alpha})$ is one of the basic Schubert varieties, see [24, 25]. Hence, there is a one-by-one map
${\mathbb{Z}}\check{\mathcal{S}}/{\mathbb{Z}}\check{\mathcal{S}}_{P}\longrightarrow H_{2}(G_{\mathbb{C}}/P;{\mathbb{Z}}),\quad\check{\alpha}+{\mathbb{Z}}\check{\mathcal{S}}_{P}\longmapsto[C_{\alpha}]$
where ${\mathbb{Z}}\check{\mathcal{S}}:=\{k\check{\alpha}|\alpha\in\mathcal{S},\;k\in{\mathbb{Z}}\}$ and ${\mathbb{Z}}\check{\mathcal{S}}_{P}:=\{k\check{\alpha}|\alpha\in\mathcal{S}_{P},\;k\in{\mathbb{Z}}\}$.
With the above preparation we now cite a result of Caviedes Castro [8] on the estimate of the quantity $\Theta({\mathcal{O}}_{\lambda},\omega_{\lambda})$ which is used to bound Hofer-Zehnder capacity of coadjoint orbits of compact Lie groups via Lu’s theorem [44, Corollary 1.19].
Lemma A.1 ([8, Theorem 5.2]).
Let $G$ be a compact simple Lie group, and $({\mathcal{O}}_{\lambda},\omega_{\lambda})$ a monotone coadjoint orbit passing through $\lambda\in{\mathfrak{t}}^{*}$. If there exist orthogonal positive roots $\alpha_{1},\ldots,\alpha_{r}\in\mathfrak{R}$ such that the longest element $w_{0}$ of the Weyl group $W$ has the composition $w_{0}=s_{\alpha_{1}}\cdots s_{\alpha_{r}}$, then
$\Theta({\mathcal{O}}_{\lambda},\omega_{\lambda})\leq\sum_{k=1}^{r}|\langle\lambda,\check{\alpha}_{k}\rangle|.$
Combining the above lemma with Theorem 1.1, we obtain
Theorem A.1.
Let $G$ be a compact simple Lie group, and $({\mathcal{O}}_{\lambda},\omega_{\lambda})$ a monotone coadjoint orbit passing through $\lambda\in{\mathfrak{t}}^{*}$. Then
#$\#{\rm Fix}(\varphi)\geq\sup\limits_{\begin{subarray}{c}\alpha_{1},\ldots,\alpha_{r}\in\mathfrak{R},\\ w_{0}=s_{\alpha_{1}}\cdots s_{\alpha_{r}}\end{subarray}}\bigg{\lceil}\frac{p\cdot{\rm cuplength}(M;{\mathbb{C}})}{\sum_{k=1}^{r}|\langle\lambda,\check{\alpha}_{k}\rangle|}\bigg{\rceil}$ (A.1)
for all $\varphi\in\operatorname{Ham}({\mathcal{O}}_{\lambda},\omega_{\lambda})$, where
$p=\min\bigg{\{}\sum_{i=1}^{\ell}\ell_{i}\langle\lambda,\check{\alpha}_{i}\rangle>0\;\big{|}\;\alpha_{i}\in\mathcal{S}\setminus\mathcal{S}_{P},\;\ell_{i}\in{\mathbb{Z}}\bigg{\}}$
is the minimal sympletic period of $({\mathcal{O}}_{\lambda},\omega_{\lambda})$.
Remark A.2.
In [8] Caviedes Castro used the monotonicity condition to identify the algebraic Gromov–Witten invariant defined on complex homogeneous projective manifolds with the symplectic Gromov–Witten invariants, and then resorted to the works of Fulton-Woodward [25] and Postnikov [56] to obtain the upper bound of $\Theta({\mathcal{O}}_{\lambda},\omega_{\lambda})$. If the equivalence of these two definitions of symplectic Gromov–Witten invariants and their algebraic counterparts holds true in general, then one can drop the monotonicity assumption and still has the inequality (A.1).
Appendix B Symplectic toric manifolds
We recall here a definition and some facts on symplectic toric manifolds, and refer the reader to the books [4, 48] for more details on this subject.
Let ${\mathbb{C}}^{n}$ be the standard $2n$-dimensional real vector space equipped with the standard symplectic form $\omega_{0}$, i.e., the imaginary part of $-\frac{1}{2}\sum dz_{i}\wedge d\bar{z}_{i}$. Let ${\mathbb{T}}^{k}$ be a $k$-dimensional torus with Lie algebra $\mathfrak{t}$. Denote the integer lattice of $\mathfrak{t}$ by
$\Delta:=\{\xi\in{\mathfrak{t}}\;|\;\exp(\xi)=1\}$
and its dual lattice by
$\Delta^{*}:=\{w\in\mathfrak{t}^{*}|\langle w,\xi\rangle\in{\mathbb{Z}}\;\;\hbox{for any}\;\xi\in\Delta\}.$
Suppose that ${\mathbb{T}}^{k}$ acts on ${\mathbb{C}}^{n}$ by Hamiltonian linear transformations
$z_{i}\longmapsto\exp(-2\pi\sqrt{-1}\langle{\mathsf{w}}_{i},\xi\rangle)z_{i},\quad i=1,\ldots,n$
where ${\mathsf{w}}_{i}\in\Delta^{*}$. The moment map of this action is given by
$\mu(z)=\pi\sum_{i=1}^{n}|z_{i}|{\mathsf{w}}_{i}$
for any $z\in{\mathbb{C}}^{n}$. Here we require that $\mu$ is proper and the vectors ${\mathsf{w}}_{i}$ span the vector space ${\mathfrak{t}}^{*}$, and hence the linear space of solutions $(\alpha_{1},\ldots,\alpha_{n})\in{\mathbb{R}}^{n}$ for the linear equation $\sum_{i=1}^{n}\alpha_{i}{\mathsf{w}}_{i}=0$ in ${\mathfrak{t}}^{*}$ has dimension $n-k$.
Definition B.1.
A toric symplectic manifold is a symplectic quotient at a regular value $\tau\in{\mathfrak{t}}^{*}$ of the moment map $\mu$ by ${\mathbb{T}}^{k}$ such that ${\mathbb{T}}^{k}$ acts freely on $\mu^{-1}(\tau)$, namely
$M_{\tau}=\mu^{-1}(\tau)/{\mathbb{T}}^{k}$
which has a real dimension $2m=2n-2k$ and a symplectic form $\omega_{\tau}\in\Omega^{2}(M_{\tau})$ induced by the symplectic reduction of $({\mathbb{C}}^{n},\omega_{0})$.
We denote by $H^{*}(M_{\tau})$, respectively $H_{*}(M_{\tau})$, the quotients of the integral (co)homology groups by their torsion subgroups. For $j\in I:=\{1,\ldots,n\}$, we denote by
${\rm cone}(I_{j})=\big{\{}\sum_{i\neq j}\alpha_{i}{\mathsf{w}}_{i}\;\big{|}\;\alpha_{i}\geq 0\big{\}}$
the positive cone spanned by the vectors ${\mathsf{w}}_{i},\;i\in I\setminus\{j\}$. Consider now the subspace
$E_{j}=\{z=(z_{1},\ldots,z_{n})\in{\mathbb{C}}^{n}\;|\;z_{j}=0\}.$
Obviously, $\tau\in{\rm cone}(I_{j})$ if and only if $E_{j}\cap\mu^{-1}(\tau)\neq 0$. In this case, one can see that the corresponding submanifold
$X_{j}=E_{j}\cap\mu^{-1}(\tau)/{\mathbb{T}}^{k}=\{[z]\in M_{\tau}\;|\;z_{j}=0\}.$
is a smooth divisor of $M_{\tau}$ and determines a codimension-$2$ class $[X_{j}]$ in $H_{2m-2}(M_{\tau})$. It is shown in [30] that $[X_{j}]$ ($1\leq j\leq n$) generate the homology ring of $H_{*}(M_{\tau})$ w.r.t. the usual homology intersection product $\cap$ in the following sense:
  • the homology classes $[X_{1}],\ldots,[X_{n}]$ span $H_{2m-2}(M_{\tau})$,
  • every relation $\sum_{i=1}^{n}\alpha_{i}{\mathsf{w}}_{i}=0$ implies a corresponding relation $\sum_{i=1}^{n}\alpha_{i}[X_{i}]=0$,
  • for any subset $J=\{i_{1},\ldots,i_{j}\}\subset I$, the relation $\tau\notin\{\sum_{i\in I\setminus J}\alpha_{i}{\mathsf{w}}_{i}\;\big{|}\;\alpha_{i}\geq 0\}$ implies a corresponding relation $[X_{i_{1}}]\cap\cdots\cap[X_{i_{j}}]=0$,
  • the above relations are the only ones among the classes $[X_{i}]$ in $H_{*}(M_{\tau})$.
From this fact one can see that $H_{2}(M_{\tau})$ can be identified with the sublattice of $\Delta\subset{\mathfrak{t}}$:
$\Delta(\tau)=\{\xi\in\Delta\;|\;\langle{\mathsf{w}}_{j},\xi\rangle=0\quad\hbox{for any}\;j\;\hbox{with}\;\tau\notin{\rm cone}(I_{j})\},$
and we denote the bijective map by
$\Delta(\tau)\longrightarrow H_{2}(M_{\tau}),\quad\eta\longmapsto A_{\eta}.$ (A.1)
Then, the dual $H^{2}(M_{\tau})$ corresponds to a quotient of the integer lattice in ${\mathfrak{t}}^{*}$. Further, the first Chern class $c_{1}(TM)$ and the cohomology class of $\omega_{\tau}$ are given by
$c_{1}(TM)=\overline{{\mathsf{w}}}_{1}+\ldots+\overline{{\mathsf{w}}}_{n},\quad[\omega_{\tau}]=\overline{\tau}$
where $\overline{\gamma}$ is the image of $\gamma\in{\mathfrak{t}}^{*}$ under the above correspondence
${\mathfrak{t}}^{*}=\Delta^{*}\otimes{\mathbb{R}}\longrightarrow H^{2}(M_{\tau};{\mathbb{R}}),\quad\gamma\longmapsto\overline{\gamma}.$
Let $C(\tau)\subset{\mathfrak{t}}^{*}$ be the component of $\tau$ in the set of regular values of the moment map $\mu$ which is called the Kähler cone of $M_{\tau}$. Its dual cone in $\Delta(\tau)$ is given by
$\Delta^{\rm eff}(\tau)=\{\xi\in\Delta(\tau)\;|\;\langle t,\xi\rangle\geq 0\quad\hbox{for any}\;t\in C(\tau)\}.$
It is well known that each nontrivial class in $H_{2}(M_{\tau})$ corresponding to an element in $\Delta^{\rm eff}(\tau)$ can be represented by a holomorphic curve in $M_{\tau}$. Such classes are called effective classes.
Definition B.2.
We say that a toric manifold $M_{\tau}$ is Fano if the vector $c_{1}=\sum_{i=1}^{n}{\mathsf{w}}_{i}$ satisfies $\langle c_{1},\xi\rangle>0$ for all $\xi\in\Delta^{\rm eff}(\tau)\setminus\{0\}$.
Let $\mathcal{D}^{\rm eff}(\tau)\subset{\mathbb{Z}}^{n}$ denote the cone
$\mathcal{D}^{\rm eff}(\tau)=\{(\langle{\mathsf{w}}_{1},\xi\rangle,\ldots,\langle{\mathsf{w}}_{n},\xi\rangle)\;|\;\xi\in\Delta^{\rm eff}(\tau)\}.$ (A.2)
Then we have a bijection $\Delta^{\rm eff}(\tau)\to\mathcal{D}^{\rm eff}(\tau)$, and denote its inverse map by
$\mathcal{D}^{\rm eff}(\tau)\longrightarrow\Delta^{\rm eff}(\tau),\quad d\longmapsto\xi_{d}.$ (A.3)
For $\eta\in\Delta(\tau)$, we denote by $A_{\eta}\in H_{2}(M_{\tau})$ the corresponding homology class as in (A.1). Let ${\rm GW}_{A}([X]^{*d},\alpha)$ denote the Gromov-Witten invariants as in (1.7). From the Batyrev-Givental formula for the quantum homology of $M_{\tau}$, it is not hard to see the following
Theorem B.1.
[48, Theorem 11.3.3] If $M_{\tau}$ is a Fano toric manifold, then the equality
${\rm GW}_{A_{\eta}}([X]^{*d^{+}},\alpha)={\rm GW}_{A_{\eta}-A_{\xi_{d}}}([X]^{*d^{-}},\alpha)$
holds for each $d\in\mathcal{D}^{\rm eff}(\tau)$, each $\eta\in\Delta(\tau)$ and each $\alpha\in H_{*}(M_{\tau})$ with the $n$-tuples $d^{+}$ and $d^{-}$ given by
$d_{i}^{+}:=\max\{d_{i},0\},\quad d_{i}^{-}:=\max\{-d_{i},0\}.$