### v5

# Degenerate symplectic fixed points and Gromov-Witten invariants

Abstract
We establish a connection between Gromov-Witten invariants and the number of fixed points of Hamiltonian diffeomorphisms on a closed rational symplectic manifold via deformed Hamiltonian spectral invariants. We generalize Givental’s symplectic fixed point theorem for Fano toric manifolds to closed rational symplectic manifolds which admit nonzero Gromov-Witten invariants with fixed marked points and one point insertion. We prove a new cuplength estimate of symplectic fixed points involved in deformed spectral invariants. We extend Schwarz’s quantum cuplength to the notion of deformed quantum cuplength for symplectic periods and employ it to estimate the number of fixed points of Hamiltonian diffeomorphisms on monotone symplectic manifolds with nonzero mixed Gromov-Witten invariants.
1 Introduction
The Arnol’d conjecture [2, 3] asserts that every Hamiltonian diffeomorphism $\varphi$ of a closed symplectic manifold $(M,\omega)$ possesses at least as many fixed points as a smooth function $f:M\to{\mathbb{R}}$ on $M$ possesses critical points. This longstanding conjecture has a rich history, detailed in the work of Hofer and Zehnder [32, Chapter 6]. The most general answer obtained by Rudyak-Oprea [59] who showed that this conjecture holds for closed symplectically aspherical manifolds. The main motivation of this paper is the following weaker version, i.e., the homological Arnol’d conjecture:
$\{\hbox{fixed points of}\;\varphi\}\geq{\rm cuplength}(M;R).$
Here ${\rm cuplength}(M;R)$ denotes the cuplength of $M$ over a coefficient ring $R$ which is defined as the maximal number $k+1\in{\mathbb{N}}$ such that there exist $k$ cohomology classes $a_{i}\in H^{*}(M;R)$ with $\deg(a_{i})>0$ satisfying $a_{1}\cup\cdots\cup a_{k}\neq 0$. If all the fixed points of $\varphi$ are a priori nondegenerate, then the nondegenerate homological Arnol’d conjecture asserts that the number of the fixed points of $\varphi$ is not less than the Betti sum of $M$ over any coefficient ring. Since the advent of Floer homology [15, 16, 17], there has been a huge progress in the nondegenerate version, for instance, in the 1990s, Fukaya-Ono [20], Liu-Tian [40] and Ruan [58] independently settled Arnol’d conjecture over rational numbers; more recently, Abouzaid-Blumberg [1], and Bai-Xu [5] and Rezchikov [56] extended these results to the coefficients over any finite field and intergers, respectively. For earlier results on the nondegenerate Arnol’d conjecture, we refer to [11, 17, 30, 51], etc. It is worth to mention that Pardon [53] and Filippenko-Wehrheim [14] employed different versions of the virtual technique to reprove Arnol’d conjecture over rational numbers.
However, compared to the nondegenerate case, progress on the degenerate homological Arnold conjecture has not kept pace and, in fact, remains far from being solved except for some special cases, see for instance a standard torus $(\mathbb{T}^{2n},\omega_{0})$ by Conley and Zehnder [11], symplectic manifolds $(M,\omega)$ with $\pi_{2}(M)=0$ by Floer [18] and Hofer [29], the complex projective space ${\mathbb{P}}^{n}$ with the Fubini-Study form $\omega_{\rm FS}$ by Fortune [19], and negatively monotone symplectic manifolds $(M^{2n},\omega)$ with minimal Chern number at least $n$ by Le and Ono [38], etc. As far as the author knows, $({\mathbb{P}}^{n},\omega_{\rm FS})$ is the only monotone symplectic manifold, i.e., $\omega|_{\pi_{2}(M)}=\lambda c_{1}|_{\pi_{2}(M)}\neq 0$ with $\lambda>0$, for which the degenerate homological Arnol’d conjecture has been proven until now.
As the title of [17] indicated, Floer was aware of the significance of $J$-homolophic spheres toward the degenerate Arnol’d conjecture more than 35 years ago. It is well known that $J$-holomorphic curves serve as the fundamental building blocks in the construction of Gromov-Witten invariants. These invariants have some important applications in contact/symplectic dynamics. For example, one was given by Liu-Tian [41] and Taubes [63] to prove Weinstein conjecture for certain contact manifolds in high dimensions and for arbitrary closed $3$-dimensional contact manifolds respectively by using the nonvanishing property of these invariants (prior to these two results, Hofer and Viterbo [31] and Chen [10] had been aware of the important role of the appearance of $J$-holomorphic spheres toward Weinstein conjecture); another one was first discovered by Lu [43] and latter reproved by Usher [68] for bounding Hofer-Zehnder capacity from above for a large class of symplectic manifolds, etc.
It was commented by McDuff and Salamon [47, Remark 9.4.12]: “The existence of a nontrivial Gromov–Witten invariant is more significant, though as yet the geometric and dynamical consequences of this condition are far from being fully understood. So far, most applications have come via properties of the small quantum cohomology ring …” In this paper, we will use nonzero mixed Gromov-Witten invariants via spectral invariants of homology classes in the big quantum homology $H_{*}(M;\Lambda)$ with quantum products $_{\zeta}$111if ${\rm dim}M=2n$, the deformation parameter $\zeta$ is an element of $\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$ for a certain subring $\Lambda_{0}$ of $\Lambda$ containing $\mathbb{C}$}, see Section 2.2. to estimate the number of fixed points of Hamiltonian diffeomorphisms where $\Lambda$ is the Novikov ring over $\mathbb{C}$}.
The mixed Gromov-Witten invariant ${\rm GW}_{0,l+k,A}^{\{1,\ldots,l\}}(a_{1},\ldots,a_{l},b_{1},\ldots,b_{k})$ from Ruan and Tian [57] (see also [47, Definition 7.3.7]) formally counts (with appropriate signs) the tuples $(u,\vec{w})=(u,\{w_{j}\}_{j=1}^{k})$ consisting of a $\{J_{z}\}$-holomorphic sphere $u:S^{2}\to M$ representing the class $A$ with $u(z_{i})\in\alpha_{i}$ for $1\leq i\leq l$ and $u(w_{j})\in\beta_{j}$ for $1\leq j\leq k$, where $\{z_{i}\}_{i=1}^{l}$ is a tuple of fixed pairwise distinct points on $S^{2}$ and the pairwise distinct marked points $w_{j}\in S^{2}\setminus\{z_{i}\}_{i=1}^{l}$ are allowed to vary freely for $1\leq j\leq k$, and $\alpha_{i},\beta_{j}$ are generic representatives of the classes $a_{i}$ and $b_{j}$ respectively. Note that when fixing $3$ pairwise distinct points on $S^{2}$, the mixed Gromov-Witten invariant ${\rm GW}_{0,k+3,A}^{\{1,2,3\}}(a_{1},a_{2},a_{3},b_{1},\ldots,b_{k})$ is precisely the usual Gromov-Witten invariant ${\rm GW}_{0,k+3,A}(a_{1},a_{2},a_{3},b_{1},\ldots,b_{k})$ as defined in [47, Section 7.1] for the semipositive case and [20] for the general case.
Our first theorem gives a concrete connection, for a rational symplectic manifold $(M,\omega)$, between the number of fixed points of Hamiltonian diffeomorphisms and nonzero mixed Gromov-Witten invariants with one point insertion. Here “rational” means that $\omega(\pi_{2}(M))=p\cdot{\mathbb{Z}}$ for some positive number $p$, called the minimal symplectic period of $(M,\omega)$. Throughout this paper, we set $H_{2}(M)=H_{2}(M;\mathbb{Z})/{\rm torsion}$.
Theorem 1.1.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],b_{1},\ldots,b_{k})$
where $A\in H_{2}(M)$, $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{Q}})$ with $2\leq l\in{\mathbb{N}}$, and the classes $b_{1},\ldots,b_{k}$ are rational homology classes of even degree (when $k=0$ there is no class $b_{i}$ insertion). Assume that
$\langle[\omega],A\rangle=\inf\big\{\langle[\omega],B\rangle\big|\;{\rm GW}_{0,l+i+1,B}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},c,b_{1},\ldots,b_{i})\neq 0\big\}$ (1.1)
where the infimum is taken over any $c\in H_{*}(M;{\mathbb{Q}})$, any $i\in{\mathbb{N}}\cup\{0\}$ and any $b_{j}\in H_{\rm even}(M;{\mathbb{Q}})$ with $j\in\{1,\ldots,i\}$. Then, for each $\varphi\in\operatorname{Ham}(M,\omega)$,
${\rm Fix}(\varphi)\geq\bigg\lceil\frac{pl}{\langle[\omega],A\rangle}\bigg\rceil$ (1.2)
where $\lceil\cdot\rceil$ denotes the smallest integer that is greater or equal to the given number.
A direct consequence of Theorem 1.1 is the following two corollaries.
Corollary 1.1.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant of the form:
${\rm GW}_{0,k+3,A}(a_{0},a_{1},[pt],b_{1},\ldots,b_{k})$
where $A\in H_{2}(M)$, $a_{0},a_{1}\in H_{*<2n}(M;{\mathbb{Q}})$, and the classes $b_{1},\ldots,b_{k}$ are rational homology classes of even degree. If $\langle[\omega],A\rangle$=p, then each Hamiltonian diffeomorphism $\varphi$ of $(M,\omega)$ has at least two fixed points.
A symplectic manifold is called uniruled if there is a nonzero Gromov-Witten invariant involving a point insertion and with a nontrivial class curve. Uniruledness is a fundamental concept in birational geometry and is invariant under symplectic birational cobordism (see [34]). For more results on the notion of uniruledness and related applications, we refer to [36, 58, 45, 39, 43, 44], etc.
The symplectic effective cone of a closed symplectic manifold $(M,\omega)$ is defined by
$K^{\rm eff}(M,\omega)=\big\{\beta\in H_{2}(M)\;|\;\exists\;\beta_{1},\dots,\beta_{N}\in H_{2}(M)\;\hbox{such that}\;\beta=\sum_{i}^{N}\beta_{i},\;{\rm GW}_{0,3,\beta_{i}}\neq 0\big\}.$
A symplectic Fano manifold $(M,\omega)$ is defined to be one in which the class $c_{1}(TM)$ is positive on the nonzero spherical elements in $K^{\rm eff}(M,\omega)$ (cf. [47, p.434]).
Corollary 1.2.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant with fixed marked points:
${\rm GW}_{0,l+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt])$
where $A\in H_{2}(M)$, $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{Q}})$ with $2\leq l\in{\mathbb{N}}$. Let
$\hbar^{\rm eff}(M,\omega)=\inf\{\langle[\omega],\beta\rangle\;|\;\beta\in K^{\rm eff}(M,\omega)\}.$
If $\langle[\omega],A\rangle=\hbar^{\rm eff}(M,\omega)$, then
${\rm Fix}(\varphi)\geq\bigg\lceil\frac{2p}{\max_{i}\{2n-a_{i}\}}\cdot\frac{\langle c_{1}(TM),A\rangle}{\langle[\omega],A\rangle}\bigg\rceil$ (1.3)
for all $\varphi\in\operatorname{Ham}(M,\omega)$.
For $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, we denote by $\gamma_{\zeta}(\varphi)$ the $\zeta$-defomed spectral norm of $\varphi\in\operatorname{Ham}(M;\omega)$ (see Definition 3.5). We have the following fixed point estimate of $\varphi$ in terms of $\gamma_{\zeta}(\varphi)$.
Theorem 1.2.
Let $(M,\omega)$ be a closed rational symplectic manifold with period $p>0$. Then, for any $\zeta\in H(M;\Lambda_{0})$ and any $\varphi\in\operatorname{Ham}(M;\omega)$, we have
$\operatorname{Fix}(\varphi)\geq\frac{\operatorname{cuplength}(M;{\mathbb{C}})}{\lfloor\gamma_{\zeta}(\varphi)/p\rfloor+1}$
where $\lfloor a\rfloor$ denotes the largest integer that is less than or equal to $a$.
Since the Hofer norm $\|\varphi\|$ of $\varphi\in\operatorname{Ham}(M;\omega)$ is not less than the $\zeta$-deformed spectral norm $\gamma_{\zeta}(\varphi)$ (cf. (3.1)), we have the following corollary which is a straightforward generalization of [26, Theorem 8].
Corollary 1.3.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$. If the Hofer norm of $\varphi$ satisfies $\|\varphi\|<p$, then the degenerate Arnold conjecture over complex numbers holds for $\varphi$.
Remark 1.4.
Prior to Corollary 1.3, Schwarz has proved that Arnol’d conjecture over ${\mathbb{Z}}_{2}$ holds for $\|\varphi\|<p$; see [61, Theorem 1.1].
Through introducing the $\zeta$-deformed quantum cuplength ${\rm qcl}_{\zeta}$} for symplectic periods (cf. Definition 2.4), our approach developed in this paper gives the following extension of Schwarz’s estimate [61, Theorem 1.2] for symplectic fixed points to all closed rational symplectic manifolds.
Theorem 1.3.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with period $p>0$. Then, for each $\varphi\in\operatorname{Ham}(M,\omega)$, we have the estimate
${\rm Fix}(\varphi)\geq\sup\limits_{\zeta\in H_{*}(M;\Lambda_{0})}\sup\limits_{g\in\Gamma_{\omega}}\bigg\lceil\frac{p\cdot{{\rm qcl}_{\zeta}}(g)}{p+g+\|\varphi\|}\bigg\rceil.$
To the best of the author’s knowledge, for closed monotone symplectic manifolds, Schwarz [61] has extended all previously known estimates for symplectic fixed points via his approach of quantum cuplength estimate. Our next theorem reveals that his result [61, Corollary 1.3] can be reinterpreted in terms of nonzero mixed Gromov-Witten invariants.
Theorem 1.4.
Let $(M^{2n},\omega)$ be a closed monotone symplectic manifold with minimal Chern number $N\geq 1$, and let $a\in H_{*<2n}(M;{\mathbb{Q}})$. Suppose that for each $2n<l\in{\mathbb{N}}$, there exists a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A_{l}}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b_{0}^{l},\ldots,b_{k}^{l})$ (1.4)
for some integer $k=k(l)\in{\mathbb{N}}\cup\{0\}$, where $A_{l}\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, and $b_{0}^{l},\ldots,b_{k}^{l}$ are rational homology classes of even degree. Then, for any $\varphi\in\operatorname{Ham}(M,\omega)$,
${\rm Fix}(\varphi)\geq\frac{2N}{2n-\deg(a)}$
provided that the function $l\mapsto k(l)$ is bounded.
For $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M,\Lambda_{0})$, we call a class $a\in H_{*<2n}(M;{\mathbb{C}})$ a $\zeta$-nonnilpotent element in the $\zeta$-deformed quatum homology ring $QH_{*}(M,\Lambda)_{\zeta}$ if $\underbrace{a*_{\zeta}\cdots*_{\zeta}a}_{l}\neq 0$ for all $l\in{\mathbb{N}}$. Note that the existence of a $\zeta$-nonnilpotent element implies that for each $l\geq 2n$, one has a nonzero mixed Gromov-Witten invariant as in (1.4). Furthermore, if $\zeta=0$, the function $k(l)$ as in Theorem 1.4 can be taken to be identically one. Hence, we obtain
Corollary 1.5 ([61, Corollary 1.3]).
Let $(M,\omega)$ be a monotone symplectic manifold with minimal Chern number $N$ determined by $c_{1}(\pi_{2}(M))=N\cdot{\mathbb{Z}}$. Suppose that $a\in H_{k}(M;{\mathbb{C}})$ ($0\leq k<2n$) is a 0-nonnilpotent element in $QH_{*}(M,\Lambda)_{0}$. Then, for all $\varphi\in\operatorname{Ham}(M,\omega)$,
${\rm Fix}(\varphi)\geq\frac{2N}{2n-k}.$
Using Corollary 1.5, Schwarz [61, Sect.2.5] proved the following:
  • (i)
    Let $M={\mathbb{P}}^{n}\times W$ be the product symplectic manifold where ${\mathbb{P}}^{n}$ is equipped with the standard symplectic form, and $W$ is symplectically aspherical. Then, each Hamiltonian diffeomorphism of $M$ has at least $n+1$ fixed points.
  • (ii)
    Let ${\rm Gr}(k,n)$ be the Grassmannian of $k$-planes in ${\mathbb{C}}^{n}$ equipped with the symplectic form $\omega$ satisfying $c_{1}({\rm Gr}(k,n))=n[\omega]$. Then, each $\varphi\in\operatorname{Ham}({\rm Gr}(k,n),\omega)$ has at least $n$ fixed points. In particular, this recovers Fortune’s theorem [19].
  • (iii)
    Let ${\rm F}_{n+1}$ be the complete flag manifold of sequences $V^{1}\subset\ldots\subset V^{n}$ of subspaces of ${\mathbb{C}}^{n+1}$ with ${\rm dim}_{\mathbb{C}}V^{i}=i$ equipped with the symplectic form $\omega$ satisfying $c_{1}({\rm F}_{n+1})=2[\omega]$. Then, for any $\varphi\in\operatorname{Ham}({\rm F}_{n+1},\omega)$, ${\rm Fix}(\varphi)\geq 2$.
Inspired by Schwarz’s quantum cuplength estimate, a different way for estimating symplectic fixed points was obtained by the author via introducing the so-called fundamental quantum factorization; see [26, Definition 23]. Using the big quantum homolgy ring (cf. Section 2.2), one can extend this notion as follows.
Definition 1.6.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$. We say that $M$ has a principal fundamental quantum factorization associated to $\zeta\in\oplus_{i=1}^{n-1}H_{2i}(M,\Lambda_{0})$ (denoted by PFQF for short) of length $l\in{\mathbb{N}}$ with order $g\in{\mathbb{N}}$ if there exist $u_{1},\ldots,u_{l}\in H_{*<2n}(M;{\mathbb{C}})$ such that
$u_{1}*_{\zeta}\cdots*_{\zeta}u_{l}=kT^{gp}[M]+\beta$
with $I_{\nu}(\beta)\leq-gp$, where $0\neq k\in{\mathbb{C}}$, $\beta\in\oplus_{i=0}^{2n-1}H_{i}(M;\Lambda)$ and $I_{\nu}:H_{*}(M;\Lambda)\to\Gamma_{\omega}$ is the valuation map as given in (2.4).
Theorem 1.5 ([26, Theorem 9]).
Let $(M,\omega)$ be a closed rational symplectic manifold admitting a PFQF of length $l$ with order $g$. Then each $\varphi\in\operatorname{Ham}(M,\omega)$ has at least $\lceil\frac{l}{g}\rceil$ fixed points.
In the small quantum homology $H_{*}(M;\Lambda)$ with product $_{0}$ (i.e., $\zeta=0$), if there is a nonzero class $\beta\in H_{<2n}(M;\Lambda_{0})$ such that
$pt]*\beta=T^{kp}[M]$
for some $k\in{\mathbb{N}}$, we say that $M$ is point invertible of order $k$. The class of point invertible manifolds includes, for example, ${\mathbb{P}}^{n}$ and the quadric $Q^{2n}\subset{\mathbb{P}}^{n+1}$. Moreover, using the product formula (see [47, Exercise 11.1.19]) for Gromov–Witten invariants, this class is closed with respect to products.
Theorem 1.6.
Let $(M^{2n},\omega)$ be a closed symplectic manifold with minimal symplectic period $p>0$ and minimal Chern number $N\geq n$. Assume that $M$ is point invertible of order $g$. If $N\geq n+1$, then, for any $\varphi\in\operatorname{Ham}(M,\omega)$,
$\operatorname{Fix}(\varphi)\geq\bigg\lceil\frac{\operatorname{cuplength}(M;{\mathbb{C}})}{g}\bigg\rceil.$
If $N=n$ and $(M^{2n},\omega)$ is monotone, then
$\operatorname{Fix}(\varphi)\geq\min\bigg\{\bigg\lceil\frac{\operatorname{cuplength}(M;{\mathbb{C}})}{g}\bigg\rceil,\operatorname{cuplength}(M;{\mathbb{C}})-1\bigg\}.$
Clearly, a point invertible manifold has a nonzero Gromov-Witten invariant with two point insertions. We remark that the condition (1.1) in Theorem 1.1 is just a technical one. The author expects that this condition can be removed. Moreover, we propose the following
Conjecture 1.7.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold with minimal symplectic period $p>0$, which admits a nonzero Gromov-Witten invariant of the form:
${\rm GW}_{0,k+3,A}\big([pt],a_{0},[pt],a_{1}\ldots,a_{k}\big)$ (1.5)
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, and $a_{0},\ldots,a_{k}$ are rational homology classes of even degree. Then, for any $\varphi\in\operatorname{Ham}(M,\omega)$,
$\operatorname{Fix}(\varphi)\geq\bigg\lceil\frac{p\cdot\operatorname{cuplength}(M;{\mathbb{C}})}{\Theta(M,\omega)}\bigg\rceil$
where $\Theta(M,\omega):=\inf\big\{\langle[\omega],A\rangle\big|{\rm GW}_{0,k+3,A}\big([pt],a_{0},[pt],a_{1}\ldots,a_{k}\big)\neq 0\big\}$. In particular, if $\Theta(M,\omega)=p$, then each $\varphi\in\operatorname{Ham}(M,\omega)$ has at least $\operatorname{cuplength}(M;{\mathbb{C}})$ fixed points.
Clearly, we have $\Theta\geq p$ in the above conjecture because $\operatorname{GW}_{0,k+3,A}\neq 0$ with two point insertions implies that the homology class $A$ can be represented by a $J$-holomorphic stable map of genus zero which has a positive symplectic area. The question of whether there are Gromov-invariants with two point insertions is related to the notion of rationally connectedness in algebraic geometry. Voisin [71] and Tian [64, 65] studied some important examples including rationally connected $3$-folds and $4$-folds on this kind of question. Actually, a folklore conjecture asserts that every rationally connected projective manifold carries a nonzero genus zero Gromov-Witten invariant with two point insertions; see for instance [35, Conjecture 3.4]. If Conjecture 1.7 holds true, then one can apply it to estimate symplectic fixed points in many cases.
1.1 Applications and examples
1.1.1 The quadric in ${\mathbb{P}}^{n+1}$
Let $Q\subset{\mathbb{P}}^{n+1}$ ($n\geq 2$) be the smooth complex $n$-dimensional quadric given by
$Q=\{[z_{0},\ldots,z_{n+1}]\in{\mathbb{P}}^{n+1}|z_{0}^{2}+\cdots+z_{n}^{2}=z_{n+1}^{2}\}.$
We endow $Q$ with the symplectic structure $\omega$ induced from $({\mathbb{P}}^{n+1},\omega_{\rm FS})$. We use the normalization that the symplectic structure $\omega_{\rm FS}$ of ${\mathbb{P}}^{n+1}$ satisfies $\int_{{\mathbb{P}}^{1}}\omega_{\rm FS}=1$. For $n=2$, $Q\subset{\mathbb{P}}^{3}$ is symplectomorphic to $({\mathbb{P}}^{1}\times{\mathbb{P}}^{1},\omega_{\rm FS}\oplus\omega_{\rm FS})$. Let $h\in H_{2n-2}(Q;{\mathbb{Z}})$ be the class of a hyperplane section induced from the embedding $Q\subset{\mathbb{P}}^{n+1}$. Note that $(Q,\omega)$ is monotone and that $c_{1}(Q)=n{\rm PD}_{Q}(h)$. Since for a simply connected algebraic manifold $M$, $\operatorname{cuplength}(M;{\mathbb{C}})={\rm dim}_{{\mathbb{C}}}(M)+1$ (see [7]), we have $\operatorname{cuplength}(Q;{\mathbb{C}})=n+1$. From Beauville’s computation of quantum cohomology (see [6]), the quantum product * on $H_{*}(Q;\Lambda)$ satisfies the following identities:
  • $h^{*k}=h^{\cap j}$ for any $0\leq k\leq n-1$;
  • $h^{*n}=2[pt]+2T^{1}[Q]$ (remind that $\int_{{\mathbb{P}}^{1}}\omega_{\rm FS}=1$);
  • $h^{*(n+1)}=4T^{1}h$.
By the second identity, one can see that
${\rm GW}_{0,n+1,A}^{\{1,\ldots,n+1\}}(\underbrace{h,\ldots,h}_{n},[pt])\neq 0$
for the generator $A$ of $H_{2}(Q;{\mathbb{Z}})$ with $\langle[\omega],A\rangle$=1. According to Theorem 1.1, we obtain
${\rm Fix}(\varphi)\geq n$
for any $\varphi\in\operatorname{Ham}(Q,\omega)$.
1.1.2 Fano toric manifolds
Let $(M^{2m}_{\tau},\omega_{\tau})$ be a toric manifold, namely a symplectic quotient ${\mathbb{C}}^{n}/\mkern-5.0mu/{\mathbb{T}}^{k}$ of ${\mathbb{C}}^{n}$ by a subtorus ${\mathbb{T}}^{k}\subset{\mathbb{T}}^{n}$; see Definition A.1. The homology $H_{2m-2}(M_{\tau};{\mathbb{Z}})/{\rm torsion}$ is spanned by the classes $X_{1}],\ldots,[X_{n}]$ corresponding to $n$ smooth divisors $X_{1},\ldots,X_{n}$ of $M_{\tau}$. For an $n$-tuple $d=(d_{1},\dots,d_{n})$ of nonnegative integers, we denote
$X]^{*d}=[X_{1}]*\cdots*[X_{1}]*\cdots*[X_{n}]*\cdots*[X_{n}].$ (1.6)
where each class $X_{i}]$ occurs $d_{i}$ times, and the right hand side stands for the (undefomed) quantum product * on $H_{*}(M_{\tau};\Lambda)$.
Notice that for a toric Fano manifold $(M_{\tau},\omega_{\tau})$, the effective cone $\Delta^{\rm eff}(\tau)$ (cf. (A.2)), when identified with the set of effective classes in $H_{2}(M_{\tau};{\mathbb{Z}})/{\rm torsion}$, is precisely the symplectic effective cone $K^{\rm eff}(M_{\tau},\omega_{\tau})$ generated by the spheres represented by the edges of Delzant polytope associated to $(M_{\tau},\omega_{\tau})$; see [47, Rmk.11.3.2]. Let $\mathcal{D}^{\rm eff}(\tau)$ be the cone in ${\mathbb{Z}}^{n}$ which corresponds to the set $\Delta^{\rm eff}(\tau)$ (see (A.4)). For each $d\in\mathcal{D}^{\rm eff}(\tau)$ with $d_{k}\geq 0$, according to Theorem A.1, if $(M_{\tau},\omega_{\tau})$ is Fano then we have
$X]^{*d}=T^{\langle[\omega],A_{\xi_{d}}\rangle}[M]$ (1.7)
where the homology class $A_{\xi_{d}}\in H_{2}(M_{\tau};{\mathbb{Z}})/{\rm torsion}$ is an effective one given by (A.1) and (A.4). Consequently,
${\rm GW}^{\{1,\ldots,\sum_{i}d_{i}+1\}}_{0,1+\sum_{i}d_{i},A_{\xi_{d}}}([X_{1}],\ldots,[X_{1}],\ldots,[X_{n}],\ldots,[X_{n}],[pt])\neq 0.$ (1.8)
Using Theorem 1.5, we now recover the following result of Givental.
Corollary 1.8 ([25, Theorem 1.1]).
Let $(M,\omega)$ be a Fano toric manifold with minimal symplectic period $p>0$. Let $\mathcal{E}$ be the set of nonzero effective homology classes of $M$ which correspond to points $d=(d_{1},\ldots,d_{n})$ in $\mathcal{D}^{\rm eff}(\tau)$ with $d_{i}\geq 0$. Then
${\rm Fix}(\varphi)\geq\sup_{A\in\mathcal{E}}\bigg\lceil\frac{p\cdot\langle c_{1}(TM),A\rangle}{\langle[\omega],A\rangle}\bigg\rceil$
for any $\varphi\in\operatorname{Ham}(M,\omega)$.
Remark 1.9.
In [25, Theorem 1.1], Givental assumed that the symplectic form is primitive, i.e., $\omega(\pi_{2}(M))={\mathbb{Z}}$. That’s why the minimal symplectic period $p$ does not occur in his lower bound. As first observed by Givental, Corollary 1.8 gives examples of non-trivial symplectic fixed point results for non-monotone symplectic manifolds; see page 453 in [25].
Proof of Corollary 1.8. For any $d\in\mathcal{D}^{\rm eff}(\tau)$ with $d_{k}\geq 0$, it follows from (1.7) that $(M,\omega)$ has a PFQF of length $\ell$ with order $\langle[\omega],A_{\xi_{d}}\rangle/p$, where $\ell$ is the number of positive components of $d$. According to Theorem 1.5, for any $\varphi\in\operatorname{Ham}(M,\omega)$ we have
${\rm Fix}(\varphi)\geq\bigg\lceil\frac{p\ell}{\langle[\omega],A_{\xi_{d}}\rangle}\bigg\rceil.$
Since we have the nonzero Gromov-Witten invariant with fixed marked points as in (1.8), the dimension formula (2.2) implies that $\langle c_{1}(TM),A_{\xi_{d}}\rangle=\ell$. Hence, we obtain
${\rm Fix}(\varphi)\geq\bigg\lceil\frac{p\cdot\langle c_{1}(TM),A_{\xi_{d}}\rangle}{\langle[\omega],A_{\xi_{d}}\rangle}\bigg\rceil.$
for any non-negative $d\in\mathcal{D}^{\rm eff}(\tau)$. The proof is complete. ∎
For a monotone toric manifold $(M,\omega)$, we can also use Theorem 1.1 to obtain the following
Corollary 1.10.
Let $(M,\omega)$ be a monotone toric manifold with minimal Chern number $N\geq 1$. Let $d=(d_{1},\ldots d_{n})$ be a point in $\mathcal{D}^{\rm eff}(\tau)$ such that
$\langle\omega,A_{\xi_{d}}\rangle=\inf\{\langle\omega,A_{\xi_{d^{\prime}}}\rangle|d^{\prime}\in\mathcal{D}^{\rm eff}(\tau)\}=\hbar(M,\omega).$
If $d$ belongs to the first orthant of ${\mathbb{Z}}^{n}$, then each Hamiltonian diffeomorphism $\varphi$ has at least $N$ fixed points.
Proof.
By (1.8), there is a nonzero Gromov-Witten invariant with fixed marked points and one point insertion in the class $A_{\xi_{d}}$. Since each divisor $X_{i}$ has codimension two, according to Corollary 1.2, we see that for any $\varphi\in\operatorname{Ham}(M,\omega)$,
${\rm Fix}(\varphi)\geq\bigg\lceil\frac{p\cdot\langle c_{1}(TM),A_{\xi_{d}}\rangle}{\langle[\omega],A_{\xi_{d}}\rangle}\bigg\rceil.$
The monotonicity condition implies
$\frac{\langle[\omega],A_{\xi_{d}}\rangle}{p}=\frac{\langle c_{1}(TM),A_{\xi_{d}}\rangle}{N}.$
Therefore, ${\rm Fix}(\varphi)\geq N$. ∎
1.1.3 $\mathbb{P}^{n}$ blown up at one point
The process of blowing up in complex geometry is well-known. Blowing up replaces a point $x$ in a complex manifold $X$ by the set $\Sigma$ of all lines through this point, which is called an exceptional divisor biholomorphic to the complex projective space $\mathbb{P}^{n-1}$. In the symplectic setting, the role of a point is played by a symplectically embedded standard ball, and blowing up amounts to removing the interior of a symplectic ball and collapsing the bounding sphere via the Hopf map to the exceptional divisor.
Let $(M,\omega)$ be a closed symplectic manifold with minimal symplectic period $p>0$ which admits a nonzero Gromov-Witten invariant of the form (1.5). Recall that the Gromov width $c_{G}$ of a symplectic manifold $(M^{2n},\omega)$ is defined as
$c_{G}:=\sup\{\pi r^{2}|\exists\;\hbox{a symplectic embedding}\;B^{2n}(r)\hookrightarrow M\}$
where $B^{2n}(r)$ stands for the open ball of radius $r>0$ with center at the origin in ${\mathbb{R}}^{2n}$. It is well known that $c_{G}\leq\Theta(M,\omega)$; see [27, 43, 44].
[…]
$\widetilde{a}=\pi_{M}^{*}a,\quad\widetilde{b}=\pi_{M}^{*}b.$
[…]
$\widetilde{\omega}]=\widetilde{[\omega]}-\pi\lambda^{2}e,\quad c_{1}(T\widetilde{M})=\widetilde{c_{1}(TM)}-(n-1)e$ (1.9)
[…]
${\rm Fix}(\varphi)\geq{\rm gcd}(n+1,2).$
1.1.4 Projective bundles over ${\mathbb{P}}^{n}$
Let $V$ be a rank $r$ bundle over ${\mathbb{P}}^{n}$, and let ${\mathbb{P}}(V)$ be the corresponding projective bundle. Let $H$ and $\xi$ be the cohomology classes of a hyperplane in ${\mathbb{P}}^{n}$ and the tautological line bundle in ${\mathbb{P}}(V)$. Denote by $\pi:{\mathbb{P}}(V)\to{\mathbb{P}}^{n}$ the natural projection. Let $h=\pi^{*}H$. If ${\mathbb{P}}(V)$ is a Fano variety, one can choose the symplectic form on ${\mathbb{P}}(V)$ to be the Kähler form $\omega$ such that $\omega]=c_{1}({\mathbb{P}}(V))$ (see [55]). Assume now that $V=\oplus^{r}_{i=1}\mathcal{O}_{{\mathbb{P}}^{n}}(m_{i})$, where $m_{i}\geq 1$ for each $i$, $m_{i}=1$ for at least one $i$, and $\sum_{i=1}^{r}m_{i}<(2n+2+r)/2$. It follows from [55, Proposition 5.19] that
$\prod_{i=1}^{r}(\bar{\xi}-\bar{h}/m_{i})=T^{r}[{\mathbb{P}}(V)]$ (1.10)
where the left-hand side stands for the product in the small quantum homology ring $H_{*}({\mathbb{P}}(V),\Lambda)$, and $\bar{\xi}$ and $\bar{h}$ are Poincaré dual to $\xi$ and $h$ respectively.
To apply Theorem 1.5 to estimate fixed points of Hamiltonian diffeomorphisms on $({\mathbb{P}}(V),\omega)$, we calculate the minimal Chern number of ${\mathbb{P}}(V)$. Since
$T{\mathbb{P}}(V)\cong\pi^{*}T{\mathbb{P}}^{n}\oplus T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}}$
where $T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}}$ donotes the vertical tangent bundle along the fibres, the first Chern class is
$c_{1}({\mathbb{P}}(V))=\pi^{*}c_{1}({\mathbb{P}}^{n})+c_{1}(T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}}).$
For a projective bundle, the relative Euler sequence (cf. [52]) reads
$0\longrightarrow\mathcal{O}_{{\mathbb{P}}(V)}\longrightarrow\pi^{*}V^{*}\otimes\mathcal{O}_{{\mathbb{P}}(V)}(1)\longrightarrow T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}}\longrightarrow 0$
where $\mathcal{O}_{{\mathbb{P}}(V)}(1)$ is the tautological line bundle on ${\mathbb{P}}(V)$. Hence,
$c_{1}(T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}})=c_{1}(\pi^{*}V\otimes\mathcal{O}_{{\mathbb{P}}(V)}(1)).$
Write $H=c_{1}(\mathcal{O}_{{\mathbb{P}}}(1))$. Then we see that
$c_{1}(V)=mH,\quad m=\sum_{i=1}^{r}m_{i}.$
Let $\xi=c_{1}(\mathcal{O}_{{\mathbb{P}}(V)}(1))$. Using the formula for the first Chern class of a tensor product of a vector bundle with line bundle,
$c_{1}(\pi^{*}V^{*}\otimes\mathcal{O}_{{\mathbb{P}}(V)}(1))={\rm rank}(V)\cdot\xi+c_{1}(\pi^{*}V^{*})=r\xi-m\pi^{*}H.$
So
$c_{1}(T_{{\mathbb{P}}(V)/{\mathbb{P}}^{n}})=r\xi-m\pi^{*}H.$
Since $c_{1}({\mathbb{P}}^{n})=(n+1)H$, we see that
$c_{1}({\mathbb{P}}(V))=(n+1)\pi^{*}H+(r\xi-m\pi^{*}H)=r\xi+(n+1-m)\pi^{*}H.$ (1.11)
Next, we determine the values of $c_{1}({\mathbb{P}}(V))$ on $\pi_{2}({\mathbb{P}}(V))$. Consider the fibre bundle
${\mathbb{P}}^{r-1}\hookrightarrow{\mathbb{P}}(V)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}{\mathbb{P}}^{n}.$
The homotopy exact sequence yields
$0\longrightarrow\pi_{2}({\mathbb{P}}^{r-1})\longrightarrow\pi_{2}({\mathbb{P}}(V))\longrightarrow\pi_{2}({\mathbb{P}}^{n})\longrightarrow 0.$
So for $r\geq 2$ we obtain
$\pi_{2}({\mathbb{P}}(V))\cong\pi_{2}({\mathbb{P}}^{r-1})\oplus\pi_{2}({\mathbb{P}}^{n})={\mathbb{Z}}\oplus{\mathbb{Z}}.$
Let $A$ and $B$ be lines in the base ${\mathbb{P}}^{n}$ and a fibre ${\mathbb{P}}^{r-1}$ respectively which satisfy
$\int_{A}\pi^{*}H=1,\quad\int_{A}\xi=0,\quad\int_{B}\pi^{*}H=0,\quad\int_{B}\xi=1.$
Then $\pi_{2}({\mathbb{P}}(V))$ is freely generated by $A$ and $B$. For any class $aA+bB\in\pi_{2}({\mathbb{P}}(V))$, by (1.11) we see that
$\int_{aA+bB}c_{1}({\mathbb{P}}(V))=a(n+1-m)+br.$
Therefore, the minimal Chern number of ${\mathbb{P}}(V)$ is
$N_{{\mathbb{P}}(V)}={\rm gcd}(|n+1-m|,r).$
By our choice of symplectic form $\omega$ on ${\mathbb{P}}(V)$, the minimal symplectic period of $({\mathbb{P}}(V),\omega)$ is also the minimal Chern number $N_{{\mathbb{P}}(V)}$. According to (1.10), $({\mathbb{P}}(V),\omega)$ has the PFQF (with respect to the undeformed product $_{0}$) of length $r$ and order $r/N_{{\mathbb{P}}(V)}$. Hence, using Theorem 1.5 we have the following
Theorem 1.7.
Let $V=\oplus^{r}_{i=1}\mathcal{O}_{{\mathbb{P}}^{n}}(m_{i})$, where $r\geq 2$, $m_{i}\geq 1$ for each $i$, $m_{i}=1$ for at least one $i$, and $\sum_{i=1}^{r}m_{i}<(2n+2+r)/2$. Let $\omega$ be a Kähler form such that $\omega]=c_{1}({\mathbb{P}}(V))$. Then, for any $\varphi\in\operatorname{Ham}({\mathbb{P}}(V),\omega)$,
${\rm Fix}(\varphi)\geq{\rm gcd}\big(|n+1-\sum_{i=1}^{r}m_{i}|,r\big).$
1.2 Organization of the paper
The paper is structured as follows: Section 2 covers preliminaries, including the Novikov ring, big quantum homology, deformed Floer homology and PSS isomorphisms. Section 3 develops deformed Hamiltonian spectral invariants and proves the key Ljusternik-Schnirelman inequality. Section 4 establishes critical links between nonzero Gromov-Witten invariants and nonvanishing big quantum products. Section 5 provides the proofs of Theorems 1.11.6 on degenetate symplectic fixed point estimates. Appendice A gives some background on symplectic toric manifolds following the books [4, 47].
Acknowledgements
I thank Jinxin Xue for his help and encouragement and for his grant supports NSFC 12271285 and the New Cornerstone investigator program. I am indebted to Egor Shelukhin for pointing out an error in the proof of Conjecture 1.7 in the first version of this manuscript. I thank Guanheng Chen, Vincent Humilière, Tian-Jun Li, Guangcun Lu, Zhiyu Tian, Jun Zhang and Shuo Zhang for valuable and inspiring remarks. I am grateful to Kaoru Ono and Michael Usher for helpful discussions about spectral invariants with bulk studied in K. Fukaya, Y.-G. Oh, H. Ohta and K. Ono [22] and M. Usher [68], respectively. Last but not the least, I thank Huagui Duan, Hui Liu, Jian Wang, Guowei Yu, Duanzhi Zhang, Qinglong Zhou and Chaofeng Zhu, the organizers of the International Conference on Advanced Topics in Dynamical Systems in Tianjin (October 2025), Kaoru Ono and Yong-Geun Oh, the organizers of RIMS & IBS-CGP joint workshop in Kyoto (November 2025), for giving me an opportunity to present a preliminary version of this work and for the superb job they did in organizing these conferences.
2 Preliminaries
2.1 Notations and conventions
The set of spherical symplectic periods of $(M,\omega)$ is denoted by
$\Gamma_{\omega}=\{\langle[\omega],A\rangle\;|\;A\in\pi_{2}(M)\}.$
We say that a symplectic manifold $(M,\omega)$ is rational if there is a number $p>0$ such that $\Gamma_{\omega}=p\cdot{\mathbb{Z}}$.
Definition 2.1.
  1. (1)
    The Novikov ring associated to $\Gamma_{\omega}$ is the ring of formal Laurent series given by
    $\Lambda=\bigg\{\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\;\big|\;a_{i}\in{\mathbb{C}},\;\lambda_{i}\in\Gamma_{\omega},\;\forall C>0,\;\#\{i\in{\mathbb{N}}|a_{i}\neq 0,\lambda_{i}<C\}<\infty\bigg\}$
    where $T$ is a formal variable.
  2. (2)
    The valuation map on $\Lambda$ is the map
    $\nu:\Lambda\to\Gamma_{\omega},\quad\nu\big(\sum_{i=1}^{\infty}a_{i}T^{\lambda_{i}}\big)=\max\{-\lambda_{i}\;|\;a_{i}\neq 0\}.$
Denote
$\Lambda_{0}:=\bigg\{\sum_{i}a_{i}T^{g_{i}}\in\Lambda\;\big|\;g_{i}\geq 0\;\forall i\bigg\}.$
Note that $\Lambda_{0}=\nu^{-1}((-\infty,0])$ and $\Lambda$ is the field of fractions of the ring $\Lambda_{0}$. We define the map $\exp:\Lambda_{0}\to\Lambda_{0}$ by the usual Taylor series
$\exp(a)=\sum_{k=1}^{\infty}\frac{a^{k}}{k!}.$
It is not difficult to see that this is a well-defined map.
Here we remark that in this paper we do not grade the Novikov ring $\Lambda$ since we are going to use a simplified set-up where bulk-deformed Floer homology and big quantum homology will be viewed as ungraded theories.
Definition 2.2.
Let $V$ be a vector space over $\Lambda$, and $\ell:V\to{\mathbb{R}}\cup\{-\infty\}$ a function. The pair $(V,\ell)$ is called a non-Archimedean normed vector space if
  • (i)
    $\ell(0)=$-$\infty$, and $\ell(c)\in{\mathbb{R}}$ for all nonzero $c\in V$;
  • (ii)
    $\ell(\alpha\cdot c)=\ell(c)+\nu(\alpha)$ for all $\alpha\in\Lambda$ and $c\in V$;
  • (iii)
    $\ell(c_{1}+c_{2})\leq\max\{\ell(c_{1}),\ell(c_{2})\}$ for all $c_{1},c_{2}\in V$.
Remark 2.3.
There is a symbolic difference between our property (ii) in the above definition and the property (F2) in [70, Definition 2.2] because of the symbolic difference for the valuation map $\nu$. It is shown that the equality in the (iii) holds if $\ell(c_{1})\neq\ell(c_{2})$; see [13, Proposition 2.1] or [70].
2.2 Big quantum homology ring
In this subsection we will review the big quantum homology which is closely related to the deformed spectral invarints.
Given $A\in H_{2}(M;{\mathbb{Z}})$, for homogeneous classes $a_{i}\in H_{*}(M;{\mathbb{Z}})/{\rm torsion},i=1\ldots,k$ with $k\geq 3$ which satisfy
$\sum_{i=1}^{k}(2n-{\rm deg}(a_{i}))=2n+2\langle c_{1}(TM),A\rangle+2k-6,$ (2.1)
we denote ${\rm GW}_{0,k,A}(a_{1},\ldots,a_{k})$ the Gromov-Witten invariant by counting (in the appropriate virtual sense) the number of $J$-holomorphic spheres in the class $A$ passing through the submanifolds $N_{i}$, where $J$ is a generic almost complex structure compactible with $\omega$ and the $N_{i}$ are generic cycles representing the classes $a_{i}$. See Ruan and Tian [57], Fukaya and Ono [20] and Ruan [58]. We put $\operatorname{GW}_{0,k,A}(a_{1},\ldots,a_{k})=0$ unless (2.1) holds, and define
${\rm GW}_{k}(a_{1},\ldots,a_{k})=\sum_{A\in H_{2}(M,{\mathbb{Z}})}{\rm GW}_{0,k,A}(a_{1},\ldots,a_{k})T^{\langle[\omega],A\rangle}.$
Then, by the Gromov compactness theorem, one can extend the above definition linearly over $\Lambda$ to a $\Lambda$-module homomorphism
${\rm GW}_{k}:H_{*}(M;\Lambda)^{\otimes k}\longrightarrow\Lambda.$
For each $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, we define the $\zeta$-deformed quantum multiplication on $H_{*}(M;\Lambda)$ as follows: for any $a,b\in H_{*}(M;\Lambda)$, the product
$_{\zeta}:H_{*}(M;\Lambda)\otimes H_{*}(M;\Lambda)\to H_{*}(M;\Lambda)$
is given by the formula
$\langle a*_{\zeta}b,c\rangle_{\rm PD_{M}}=\sum_{k=0}^{\infty}\frac{1}{k!}{\rm GW}_{k+3}(a,b,c,\underbrace{\zeta,\ldots,\zeta}_{k})$
for any $c\in H_{*}(M;\Lambda)$ where $\langle\cdot,\cdot\rangle_{\rm PD_{M}}$ denotes the Poincare duality. Alternatively, one may express this product in terms of a homogeneous basis as follows. Let $\{c_{j}\}_{j=1}^{K}$ be a homogeneous basis for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}$, namely, $c_{i}\circ c^{j}=\delta_{i}^{j}$ where $\circ$ denotes the Poincaré intersection pairing. Then the product $_{\zeta}$ can be obtained by extending linearly from the following formula
$a*_{\zeta}b=\sum_{A\in H_{2}(M;{\mathbb{Z}})}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}{\rm GW}_{0,k+3,A}(a,b,c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\langle[\omega],A\rangle}c^{j}.$
It is well-known that the product $_{\zeta}$ is associative and supercommutative with a parity decomposition $H_{*}(M;\Lambda)=H_{\rm even}(M;\Lambda)\oplus H_{\rm odd}(M;\Lambda)$; see [37] or [47, Chapter 11.5]. So we obtain a ring
$QH_{*}(M,\omega)_{\zeta}=\big(H_{*}(M;\Lambda),*_{\zeta}\big)$
which obviously has the fundamental class $[M]$ as its unity, and which is called the $\zeta$-deformed homology ring of $(M,\omega)$.
Following Ruan and Tian [57], we denote by ${\rm GW}_{0,l+k,A}^{\{1,\ldots,l\}}(a_{1},\ldots,a_{l},b_{1},\ldots,b_{k})$ the mixed Gromov-Witten invariant. This invariant formally counts (with appropriate signs) the tuples $(u,\vec{w})=(u,\{w_{j}\}_{j=1}^{k})$ consisting of a $J$-holomorphic sphere $u:S^{2}\to M$ representing the class $A$ with $u(z_{i})\in\alpha_{i}$ and $u(w_{j})\in\beta_{j}$ for all $i,j$, where $\{z_{i}\}_{i=1}^{l}$ is a tuple of fixed pairwise distinct points on $S^{2}$ and $w_{j}\in S^{2}\setminus\{z_{i}\}_{i=1}^{l}$ are pairwise distinct marked points for all $j$. Here $\alpha_{i}$ and $\beta_{j}$ are generic representatives of the classes $a_{i}$ and $b_{j}$, respectively. This quantity is zero unless
$\sum_{i=1}^{l}(2n-\deg(a_{i}))+\sum_{j=1}^{k}(2n-\deg(b_{j}))=2n+2\langle c_{1}(TM),A\rangle+2k.$ (2.2)
Using the splitting, multilinearity and symmetry properties of the mixed Gromov-invariants (cf. [57] or [47, Chapter 11]), one can see that an $l$-fold product is then obtained by extending linearly from the formula
$a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}=\sum_{A\in H_{2}(M;{\mathbb{Z}})}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}{\rm GW}_{0,l+k+1,A}^{I}(a_{1}\ldots,a_{l},c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\langle[\omega],A\rangle}c^{j}$ (2.3)
where $I=\{1,\ldots,l+1\}$, and $\{c_{j}\}_{j=1}^{K}$ is a homogeneous basis for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}$.
As a vector space over $\mathbb{C}$}, the big quantum homology $H_{*}(M;\Lambda)$ is isomorphic to $H_{*}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda$. Each element $a\in H_{*}(M;\Lambda)$ can be written as $a=\sum_{g\in\Gamma_{\omega}}a_{g}T^{g}$ with $a_{g}\in H_{*}(M;{\mathbb{C}})$ and the property that for any $C>0$,
$\{g\in\Gamma_{\omega}\;|\;a_{g}\neq 0,g<C\}<\infty.$
This latter finiteness condition gives rise to a valuation function $I_{\nu}:H_{*}(M;\Lambda)\to\Gamma_{\omega}$ defined by
$I_{\nu}(a)=\max\{-g|a_{g}\neq 0\}.$ (2.4)
There exists a pairing $\Delta:H_{*}(M;\Lambda)\times H_{*}(M;\Lambda)\to\Lambda$ defined by
$\Delta\bigg(\sum a_{g}T^{g},\sum b_{h}T^{h}\bigg)=\sum a_{g}\circ b_{h}\cdot T^{g+h}$
where $\circ$ stands for the Poincaré intersection pairing. It is easy to verify that the pairing is well-defined, and nondegenerate in the sense that $\Delta(a,b)=0$ for all $b$ then $a=0$. Then, by taking coefficients of the zero-order term in $\Delta(a,b)$ with respect to the formal variable $T$, one can define a $\mathbb{C}$}-valued nondegenerate paring ${\textstyle\prod}:H_{*}(M;\Lambda)\times H_{*}(M;\Lambda)\to{\mathbb{C}}$ by
$\prod\bigg(\sum a_{g}T^{g},\sum b_{h}T^{h}\bigg)=\sum a_{g}\circ b_{-g}$ (2.5)
(see [68, p. 1359] or [22, Chapter 15]).
2.2.1 Deformed quantum cuplength
Schwarz [61, Section 2.2] generalized the notion of classical cuplength to small quantum cohomology, called quantum cuplength, and employed it to obtain some estimates for the number of 1-periodic solutions of Hamiltonian equations. Now we further generalize this invariant to our deformed quantum homology as follows.
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, for homology classes $\alpha_{1},\ldots,\alpha_{k}\in H_{*}(M;\Lambda)$, using the splitting property of the Gromov invariants, the $k$-fold big quantum product can be represented as a formal series in $H_{*}(M;\Lambda)$
$\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k}=\sum(\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k})_{g}T^{g}.$
Definition 2.4.
The $\zeta$-deformed quantum cuplength for $g\in\Gamma_{\omega}$ is defined as
${{\rm qcl}_{\zeta}}(g)=\max\big\{k+1\;|\;\exists\;a_{i}\in H_{*<2n}(M;{\mathbb{C}})\;\hbox{such that}\;(a_{1}*_{\zeta}\cdots*_{\zeta}a_{k})_{g}\neq 0\big\}$
if there exist homology classes $a_{i}\in H_{*<2n}(M;{\mathbb{C}})$ such that $(a_{1}*_{\zeta}\cdots*_{\zeta}a_{k})_{g}\neq 0$, and we let ${{\rm qcl}_{\zeta}}(g)=0$ if $(a_{1}*_{\zeta}\cdots*_{\zeta}a_{k})_{g}=0$ for arbitrary nonzero classes $a_{1},\cdots,a_{k}\in H_{*}(M;{\mathbb{C}})$.
Remark 2.5.
Note that $g>0$ if ${{\rm qcl}_{\zeta}}(g)>0$. From the properties of Gromov-Witten invariants–in particular, considering terms with no $\zeta$ insertions and no $J$-holomorphic spheres (i.e., $A=0$), one finds that
$a_{1}*_{\zeta}\cdots*_{\zeta}a_{k}=a_{1}\cap\cdots\cap a_{k}+\sum_{0<g\in\Gamma_{\omega}}b_{g}T^{g}$
for any $a_{1},\ldots,a_{k}\in H_{*}(M;{\mathbb{C}})$ where $\cap$ stands for the usual homology intersection product. Therefore, the classical cuplength coincides with ${{\rm qcl}_{\zeta}}(0)$, and ${{\rm qcl}_{\zeta}}(g)=0$ for all $g\in\{\gamma\in\Gamma_{\omega}|\gamma<0\}$.
2.2.2 Morse theoretical big quantum homology
The $\zeta$-deformed quantum homology ring $QH_{*}(M,\omega)_{\zeta}$ has a Morse theoretical interpretation which we describe as follows. Let $\{\beta_{i}\}_{i=1}^{m}$ be a basis of $\oplus_{i=0}^{n-2}H_{2i}(M;{\mathbb{Q}})$. Thom’s theorem [66, Théorème II.29] implies that for each $\beta_{i}$, one may choose a representative given by an embedding
$f_{i}:\Delta_{i}\longrightarrow M$ (2.6)
(i.e., $f_{i*}[\Delta_{i}]=\beta_{i}$). Here, each $\Delta_{i}$, $i=1,\ldots,m$, is a smooth closed oriented manifold of dimension $2d(i)$. For simplicity, hereafter we denote
$\delta(I):=\sum^{k}_{j=1}(2n-2d(i_{j})-2).$
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M,\Lambda_{0})$, we let $\theta\in\Omega^{2}(M;\Lambda_{0})$ be a closed Novikov-ring-valued two-form which is Poincaré dual to the degree-(2n2)(2n-2) component of $\zeta$. Then we have
$\zeta-{\rm PD_{M}}[\theta]=\sum_{i=1}^{m}a_{i}\beta_{i}\in\bigoplus_{i=0}^{n-2}H_{2i}(M;\Lambda_{0})$ (2.7)
where $a_{i}\in\Lambda_{0}$ for all $i\in\{1,\ldots,m\}$. For any $I=(i_{1},\ldots,i_{k})\in\{1,\ldots,m\}^{k}$, we set
$a_{I}=a_{i_{1}}\ldots a_{i_{k}}.$
Let $f\in C^{\infty}(M,{\mathbb{R}})$ be a Morse function. Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, we choose a generic Riemannian metric $\rho$ on $M$ such that the stable and unstable manifolds $W^{s}(f),W^{u}(f)$ for $x,y\in{\rm Crit}(f)$ are transverse both to each other and to the maps $f_{i}:\Delta_{i}\to M$. A grading
${\rm ind}_{f}:{\rm Crit}(f)\longrightarrow\{0,\ldots,2n\}$
is given by the Morse index. Denote
$\mathcal{M}_{x,y}(f,\rho)=\big\{\gamma\in C^{\infty}({\mathbb{R}},M)\;\big|\;\dot{\gamma}+{\rm grad}_{\rho}f(\gamma)=0,\;\gamma(-\infty)=x,\;\gamma(+\infty)=y\big\}.$
Since $\dim\widehat{\mathcal{M}}_{x,y}(f,\rho)={\rm ind}_{f}(x)-{\rm ind}_{f}(y)$ and $\widehat{\mathcal{M}}_{x,y}(f,\rho)$ carries a free $\mathbb{R}$}-action, if ${\rm ind}_{f}(x)={\rm ind}_{f}(y)+1$ then the moduli space
$\mathcal{M}_{x,y}(f,\rho)=\widehat{\mathcal{M}}_{x,y}(f,\rho)/{\mathbb{R}}$
is a compact 0-dimensional manifold. Equipping unstable manifolds with arbitrary orientations, together with the orientation of $M$, gives rise to orientations on $\widehat{\mathcal{M}}_{x,y}(f,\rho)$, which, compared with the canonical orientation by the flow, assign a sign to every element $\gamma]\in\mathcal{M}_{x,y}(f,\rho)$ provided that ${\rm ind}_{f}(x)={\rm ind}_{f}(y)+1$. Then we obtain a Morse complex
$CM_{k}(f;{\mathbb{Z}})=\sum_{\begin{subarray}{c}p\in{\rm Crit}(f),\\ {\rm ind}_{f}(p)=k\end{subarray}}{\mathbb{Z}}\langle p\rangle$
with differential
$\partial:CM_{k+1}(f;{\mathbb{Z}})\to CM_{k}(f;{\mathbb{Z}}),\quad\partial x=\sum_{{\rm ind}_{f}(x)={\rm ind}_{f}(y)+1}\sharp_{\rm alg}\mathcal{M}_{x,y}(f,\rho)\cdot y$
where $\sharp_{\rm alg}\mathcal{M}_{x,y}(f,\rho)\in{\mathbb{Z}}$ is the number of the negative gradient flowlines counted with sign. The homology of the complex $(CM(f;{\mathbb{Z}}),\partial)$ is called Morse homology of $(f,\rho)$ with integer coefficients, naturally isomorphic to the singular homology $H_{*}(M;{\mathbb{Z}})$.
We now consider a complex with $\Lambda$-coefficients
$CM_{*}(f;\Lambda)=CM_{*}(f;{\mathbb{Z}})\otimes_{{\mathbb{Z}}}\Lambda$
with differential $\partial^{f}:CM(f;\Lambda)\to CM(f;\Lambda)$ given by $\partial\otimes_{{\mathbb{Z}}}{\rm id}$. The homology of this complex is independent of the choice of $\rho$, and naturally isomorphic to $H_{*}(M;\Lambda)$. We denote it by $HM_{*}(f;\Lambda)$.
The $\zeta$-deformed quantum multiplication on $H_{*}(M;\Lambda)$ has the following Morse theoretical description. Let $(f,\rho^{0}),(g,\rho^{1}),(h,\rho^{2})$ be three Morse-Smale pairs as above. Let $J$ be an almost complex structure on $M$ which smoothly depends on $z\in{\mathbb{C}}P^{1}$. Given $A\in\pi_{2}(M)$, we consider the space of $J$-holomorphic spheres
$\mathcal{M}_{A}(J)=\{u:{\mathbb{C}}P^{1}\to M\;|\;du(z)\circ j=J(z,u(z))du(z),\;[u]=A\}.$
Now we define
$\mathcal{M}_{k}(A,J)=\big\{(u,\vec{z})\in\mathcal{M}_{A}(J)\times({\mathbb{C}}P^{1})^{k}\big\},$
and denote by $\overline{\mathcal{M}}_{k}(A,J)$ the compactification of this space with evaluation maps $ev_{i}\;(1\leq i\leq k)$ at the $i$-th marked point. For the marked space $({\mathbb{C}}P^{1},0,-1,\infty)$, each element $(u,\vec{z})$ in $\mathcal{M}_{k}(A,J)$ carries the natural evaluation maps $ev_{0},ev_{-1},ev_{\infty}$ given by $ev_{0}(u,\vec{z})=u(0)$, $ev_{-1}(u,\vec{z})=u(-1)$ and $ev_{\infty}(u,\vec{z})=u(\infty)$. For $p\in\operatorname{Crit}(f),q\in\operatorname{Crit}(g),r\in\operatorname{Crit}(h)$, let $\iota^{0}:W^{u}(p)\to M$, $\iota^{1}:W^{u}(q)\to M$, $\iota^{2}:W^{s}(r)\to M$ denote the inclusion maps. The fiber product
$\overline{\mathcal{M}}_{k}(A,J)_{(ev_{0},ev_{-1},ev_{\infty},ev_{1},\ldots,ev_{k})}\times_{(\iota^{0},\iota^{1},\iota^{2},f_{i_{1}},\ldots,f_{i_{k}})}W^{u}(p)\times W^{u}(q)\times W^{s}(r)\times\Delta_{i_{1}}\times\ldots\times\Delta_{i_{k}}$
has a Kuranishi structure (cf. [20, 21, 23]) with corners of dimension
$\operatorname{ind}_{f}(p)+\operatorname{ind}_{g}(q)-\operatorname{ind}_{h}(r)+2\langle c_{1}(TM),A\rangle-2n-\delta(I).$
Then where the associated multisections are denoted by $\mathsf{s}_{P,I}$, the $\zeta$-deformed quantum product
$_{\zeta}^{\rm Morse}:CM_{*}(f;\Lambda)\times CM_{*}(g;\Lambda)\longrightarrow CM_{*}(h;\Lambda)$
is defined by extending linearly from $p*_{\zeta}^{\rm Morse}q$ for $p\in\operatorname{Crit}(f)$ and $q\in\operatorname{Crit}(g)$ given by
$p*_{\zeta}^{\rm Morse}q=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}A\in\pi_{2}(M),\\ r\in\operatorname{Crit}(h)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \operatorname{ind}_{f}(p)+\operatorname{ind}_{g}(q)-\operatorname{ind}_{h}(r)=2n+\delta(I)-2c_{1}(A)\end{subarray}}|\mathsf{s}_{P,I}^{-1}(0)|\exp\big(\int_{A}\theta\big)a_{I}T^{\int_{A}\omega}r$
where $|\mathsf{s}_{P,I}^{-1}(0)|$ denotes the sum of the rational multiplicities of the points of the zero-dimensional vanishing locus of $\mathsf{s}_{P,I}$. On homology level (after identifying the Morse homology with the $\zeta$-deformed homology), the map $_{\zeta}^{\rm Morse}$ induces a map from $H_{*}(M;\Lambda)\otimes H_{*}(M;\Lambda)$ to $H_{*}(M;\Lambda)$ by sending $a\otimes b$ to
$\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}\sum_{A\in\pi_{2}(M)}\sum_{I\in\{1,\ldots,m\}^{k}}{\rm GW}_{0,k+3,A}(a,b,c_{j},a_{i_{1}}\beta_{i_{1}},\ldots,a_{i_{k}}\beta_{i_{k}})\exp\big(\int_{A}\theta\big)T^{\int_{A}\omega}c^{j}$
[…]
$\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{j=1}^{K}\sum_{A\in\pi_{2}(M)}{\rm GW}_{0,k+3,A}(a,b,c_{j},\underbrace{\zeta,\ldots,\zeta}_{k})T^{\int_{A}\omega}c^{j}$
which coincides with the $\zeta$-deformed quantum multiplication $_{\zeta}$ on $H_{*}(M;\Lambda)$.
Remark 2.6.
In the construction of big quantum homology ring, one can avoid using Kuranishi structures by imposing the assumption that $(M,\omega)$ is strongly semipositive (i.e., for any $A\in\pi_{2}(M)$ with $2-n\leq\langle c_{1}(TM),A\rangle<0$ we have $\int_{A}\omega\leq 0$) and instead follow the constructions of Hofer and Salamon [30]. See [68, Section 3.2.2]. This will not affect our applications in Section 1.1 since those examples listed there meet the latter condition. The generalization to all closed symplectic manifolds requires more subtle technical arguments. However, these arguments are less involved than the technical hurdles overcome in the construction of $A_{\infty}$–algebras associated to appropriate Lagrangian submanifolds [21]; therefore one can adapt the methods from that work or [23].
2.3 Deformed Hamiltonian Floer homology
Following Usher [68] (see also [22]), we review the construction of deformed Hamiltonian Floer homology with additional algebraic structures.
2.3.1 Deformed Floer complexes
Let ${\mathcal{H}}=C^{\infty}(S^{1}\times M,{\mathbb{R}})$. For $H\in{\mathcal{H}}$, we denote by $\{\varphi_{H}^{t}\}_{t\in[0,1]}$ the Hamiltonian isotopy of $H$ which is given by integrating the time-dependent vector field $X_{H_{t}}$, where $H_{t}=H(t,\cdot)$ and $X_{H_{t}}$ is determined uniquely by $dH_{t}=\omega(X_{H_{t}},\cdot)$. Denote by ${\rm Per}(H)$ the set of contractible $1$-periodic orbits of the Hamiltonian flow of $H$. We say that $H$ is normalized if $\int^{1}_{0}H_{t}\omega^{n}=0$ for all $t\in S^{1}$.
We consider the set $\mathcal{L}_{pt}(M)$ of all pairs $(x,\bar{x})$ where $x:S^{1}\to M$ is a loop and $\bar{x}:D^{2}\to M$ a disc with $\bar{x}|_{\partial D^{2}}=x$, called a capping of $x$. We denote by $\pi_{2}(x)$ the set of homotopy classes of such $\bar{x}^{\prime}s$. We identify $(x,\bar{x})$ and $(x^{\prime},\bar{x}^{\prime})$ if $x=x^{\prime}$ and $\bar{x}$ is homotopic to $\bar{x}^{\prime}$ relative to the boundary $x$.
Given $H\in{\mathcal{H}}$, we consider the action functional $\mathcal{A}_{H}:\mathcal{L}_{pt}(M)\longrightarrow{\mathbb{R}}$ given by
$\mathcal{A}_{H}(x,\bar{x})=-\int_{D}\bar{x}^{*}\omega+\int^{1}_{0}H(t,x(t))dt.$
Let ${\rm Crit}(\mathcal{A}_{H})$ denote the set of critical points of $\mathcal{A}_{H}$. It is easy to see that
${\rm Crit}(\mathcal{A}_{H})=\big\{(x,\bar{x})\in\mathcal{L}_{pt}(M)\;\big|\;\dot{x}(t)=X_{H_{t}}(x(t))\big\}.$
For each $H\in{\mathcal{H}}$, we define the action spectrum of $H$ as
${\rm Spec}(H)=\big\{\mathcal{A}_{H}(x,\bar{x})\;\big|\;(x,\bar{x})\in{\rm Crit}(\mathcal{A}_{H})\big\}.$
It is shown that for any $(M,\omega)$ and for any $H\in{\mathcal{H}}$, the set ${\rm Spec}(H)$ has measure zero in $\mathbb{R}$} (see Oh [48] for the general case and Schwarz [60] for aspherical $(M,\omega)$). Moreover, for any one-parameter smooth family of normalized periodic Hamiltonians $H^{s}:S^{1}\times M\to{\mathbb{R}},s\in[0,1]$ with
$\varphi_{H^{s}}^{0}\equiv id,\quad\varphi_{H^{s}}^{1}\equiv\varphi_{H^{0}}^{1}\quad\hbox{for all}\;s\in[0,1],$
we have ${\rm Spec}(H^{s})={\rm Spec}(H^{0})$; see [22, Corollary 10.3].
We say that $H$ is nondegenerate if for each $p\in{\rm Fix}(\varphi_{H}^{1})$, $1$ is not the eigenvalue of the linearization $d_{p}\varphi_{H}^{1}:T_{p}M\to T_{p}M$. Clearly, for a nondegenerate Hamiltonian $H$, the cardinality of the set ${\rm Per}(H)$ of one-periodic solutions of $\dot{x}(t)=X_{H_{t}}(x(t))$ is finite. In this case, we have an integer grading
$\mu_{H}:{\rm Crit}(\mathcal{A}_{H})\longrightarrow{\mathbb{Z}}$
induced by the Conley-Zehnder index. More specifically, in this paper, the Conley-Zehnder index is taken to be the negative of the one defined in [11]. Consequently, for a $C^{2}$-small Morse function $f$, each constant Hamiltonian one-periodic orbit $q$ (together with its constant capping $\bar{q}$) satisfies $\mu_{H}(q,\bar{q})=n-{\rm ind}_{q}(f)$, where ${\rm ind}_{q}(f)$ denotes the Morse index of $f$ at $q$.
We introduce an equivalence relation $\sim$ on ${\rm Crit}(\mathcal{A}_{H})$ such that $(x_{1},\bar{x}_{1})\sim(x_{2},\bar{x}_{2})$ if and only if $x_{1}=x_{2}$ and
$\int_{D^{2}}\bar{x}_{1}^{*}\omega=\int_{D^{2}}\bar{x}_{2}^{*}\omega.$
Denote by $\mathcal{O}(H)$ the set of such equivalence classes. The equivalence class of $(x,\bar{x})\in{\rm Crit}(\mathcal{A}_{H})$ is denoted by $x,\bar{x}]$. Then we have the maps
${\rm Crit}(\mathcal{A}_{H})\longrightarrow\mathcal{O}(H)\longrightarrow{\rm Per}(H)$
by $(x,\bar{x})\mapsto[x,\bar{x}]$ and $x,\bar{x}]\mapsto x$.
We define the Floer complex $CF_{*}(H)$ of a nondegenerate Hamiltonian $H\in{\mathcal{H}}$ as the set of elements having the formal sum
$\sum_{[x,\bar{x}]\in\mathcal{O}(H)}\alpha_{[x,\bar{x}]}[x,\bar{x}],\;\alpha_{[x,\bar{x}]}\in{\mathbb{C}}$ (2.8)
with the property that for any $C>0$,
$\big\{[x,\bar{x}]|\alpha_{[x,\bar{x}]}\neq 0,\mathcal{A}_{H}([x,\bar{x}])>C\big\}<\infty.$
The Novikov field $\Lambda$ acts on $CF_{*}(H)$ as follows: for each $\lambda\in\Gamma_{\omega}$, we pick an arbitrary class $A_{\lambda}\in\pi_{2}(M)$ satisfying $\langle[\omega],A_{\lambda}\rangle=\lambda$ and define the action
$\big(\sum a_{i}T^{\lambda_{i}}\big)\cdot\big(\sum\beta_{[x,\bar{x}]}[x,\bar{x}]\big)=\sum a_{i}\beta_{[x,\bar{x}]}[x,\bar{x}\sharp A_{\lambda_{i}}].$
It is easy to verify that this action is well-defined. Consequently, $CF_{*}(H)$ is a vector space over $\Lambda$ of finite dimension equal to $\sharp{\rm Per}(H)$. More precisely, for each $x\in{\rm Per}(H)$, if we pick a lifting $x,\bar{x}]\in\mathcal{O}(H)$, then any element $c$ of $CF_{*}(H)$ is uniquely written as a finite sum
$c=\sum_{x\in{\rm Per}(H)}a_{x}[x,\bar{x}],\;a_{x}\in\Lambda.$ (2.9)
Definition 2.7.
The filtration function $\ell_{H}$ on $CF_{*}(H)$ is defined by
$\ell_{H}(c)=\max\big\{\mathcal{A}_{H}([x,\bar{x}])\;\big|\;\alpha_{[x,\bar{x}]}\neq 0\big\}$
for any $c=\sum\alpha_{[x,\bar{x}]}[x,\bar{x}]$ of the form (2.8).
Equivalently, after fixing a lifting $x\to[x,\bar{x}]$ for each $x\in{\rm Per}(H)$, one can define the value of $\ell_{H}$ at $c$ given in the form (2.9) by
$\ell_{H}(c)=\max\{\mathcal{A}_{H}([x,\bar{x}])+\nu(a_{x})|a_{x}\neq 0\}$
where $\nu$ is the valuation map on $\Lambda$ (cf. Definition 2.1(2)). The function $\ell_{H}$ gives rise to a filtration on $CF_{*}(H)$ by
$CF_{*}^{\lambda}(H)=\{c\in CF_{*}(H)\;|\;\ell_{H}(c)\leq\lambda\}\quad\hbox{for any}\;\lambda\in{\mathbb{R}}.$
2.3.2 Gradings
As a vector space over $\Lambda$, $CF_{*}(H)$ carries a natural ${\mathbb{Z}}_{2}$-grading which we define as follows.
$CF_{0}(H)=\bigoplus_{\begin{subarray}{c}x\in{\rm Per}(H)\\ n-\mu_{H}(x,\bar{x})=0\;{\rm mod}\;2\end{subarray}}\Lambda[x,\bar{x}],$
$CF_{1}(H)=\bigoplus_{\begin{subarray}{c}x\in{\rm Per}(H)\\ n-\mu_{H}(x,\bar{x})=1\;{\rm mod}\;2\end{subarray}}\Lambda[x,\bar{x}].$
Note that the parity of $\mu_{H}(x,\bar{x})$ depends only on the orbit $x\in{\rm Per}(H)$ and not on its capping $\bar{x}$. Indeed, for any two cappings $w$ and $u$ of $x$, we have
$\mu_{H}(x,w)-\mu_{H}(x,u)=2\langle c_{1}(TM),[w\#\bar{u}]\rangle,$
where $w\#\bar{u}:S^{2}\to M$ denotes the disc $u$ with opposite orientation so that $w\#\bar{u}$ is a $2$-sphere.
2.3.3 Maslov index of homotopy classes of cylinders
Let $H_{\pm}\in{\mathcal{H}}$ and $x_{\pm}\in{\rm Per}(H_{\pm})$. We denote by $\pi_{2}(x_{-},x_{+})$ the set of path components of the space of continuous maps $u:[-\infty,+\infty]\times S^{1}\to M$ satisfying $u(\{\pm\infty\}\times S^{1})=x_{\pm}$.
Given $H_{i}\in{\mathcal{H}}$ and $x_{i}\in{\rm Per}(H_{i}),i=0,1,2$, there is the natural gluing map
$\pi_{2}(x_{0},x_{1})\times\pi_{2}(x_{1},x_{2})\longrightarrow\pi_{2}(x_{0},x_{2})$
$(C_{0},C_{1})\longmapsto C_{0}\sharp C_{1}$
where $C_{0}\sharp C_{1}$ is the equivalence class of a gluing map $w$ of any two representatives $u,v$ of $C_{0}$ and $C_{1}$. More precisely, let $u,v:[-\infty,+\infty]\times S^{1}\to M$ be two continuous maps which represent $C_{0}$ and $C_{1}$ respectively, and define $w:[-\infty,+\infty]\times S^{1}\to M$ by putting $w(s,t)=u(-\log(-s),t)$ for $s\leq 0$ and $w(s,t)=v(\log(s),t)$ for $s\geq 0$. Here $\log:[0,\infty]\to[-\infty,\infty]$ is an extension of the natural logarithm.
Similarly, one can define a concatenation map
$\pi_{2}(x_{0})\times\pi_{2}(x_{0},x_{1})\longrightarrow\pi_{2}(x_{1}),\quad(\omega,C)\longmapsto\omega\sharp C.$
For $x_{\pm}\in{\rm Per}(H_{\pm})$, we now use the Conley-Zehnder index to define a map $\mu$ on $\pi_{2}(x_{-},x_{+})$ as follows. For $C\in\pi_{2}(x_{-},x_{+})$, we pick an arbitrary representative $u:[-\infty,+\infty]\times S^{1}\to M$ of $C$, and pick any disc $w$ such that $w|_{\partial D^{2}}=x_{-}$ (note that $x_{-}$ is contractible by our assumption). We then glue the negative end of $u$ to the boundary of $w$ along $x_{-}$ to obtain a disc $w\sharp u:D^{2}\to M$ with $w\sharp u|_{\partial D^{2}}=x_{+}$. Then we define
$\mu(C)=\mu_{H_{+}}(x_{+},w\sharp u)-\mu_{H_{-}}(x_{-},w).$
Clearly, $\mu(C)$ is independent of the choice of $u$ and $w$, and has the property
$\mu(C_{0}\sharp C_{1})=\mu(C_{0})+\mu(C_{1})\quad\hbox{for all}\;C_{0}\in\pi_{2}(x_{0},x_{1}),C_{1}\in\pi_{2}(x_{1},x_{2}).$
2.3.4 Deformed Floer boundary maps
Definition 2.8.
We say that $H\in{\mathcal{H}}$ is strongly nondegenerate if for each $p\in{\rm Fix}(\varphi_{H}^{1})$, $1$ is not an eigenvalue of the linearization $d_{p}\varphi_{H}^{1}:T_{p}M\to T_{p}M$, and each submanifold $f_{i}(\Delta_{i})$, $i=1,\ldots,m$, does not intersect the set $\{\varphi_{H}^{t}(p)|t\in[0,1]\}$.
By the Sard-Smale theorem, one can show that the space of strongly nondegenerate Hamiltonians is residual in $\mathcal{H}$} in the $C^{l}$-topology for all $2\leq l\leq\infty$. See [68, Appendix A].
Let $J=\{J_{t}\}_{t\in S^{1}}$ be an $S^{1}$-family of $\omega$-compatible almost complex structures on $M$, and $H\in{\mathcal{H}}$ a strongly nondegenerate Hamiltonian. Given $x_{\pm}\in{\rm Per}(H)$, we consider the solution $u:{\mathbb{R}}\times S^{1}\to M$ of the Floer equation:
$\overline{\partial}_{J,H}u:=\partial_{s}u+J(t,u(s,t))\big(\partial_{t}u-X_{H}(t,u(s,t))\big)=0$
with the asymptotic boundary condition $u(s,t)=x_{\pm}(t)$ for $s\to\pm\infty$. Then $u$ extends continuously to a map $u:[-\infty,\infty]\times S^{1}\to M$ with $u(\pm\infty,t)=x_{\pm}(t)$, which represents some class $C\in\pi_{2}(x_{-},x_{+})$. We denote by $\widehat{\mathcal{M}}(x_{-},x_{+},C;H,J)$ the moduli space of such solutions with
$\int_{{\mathbb{R}}\times S^{1}}\big|\partial_{s}u\big|^{2}_{J_{t}}dsdt<\infty$
where $|\cdot|_{J_{t}}$ is the norm associated to an $S^{1}$-family of metrics induced by $J$ for all $t$.
Consider the product space
$\widehat{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)=\widehat{\mathcal{M}}(x_{-},x_{+},C;H,J)\times({\mathbb{R}}\times S^{1})^{k}.$
Note that $\mathbb{R}$} acts on each factor by $s$-translation, we have a quotient
$\mathcal{M}_{k}(x_{-},x_{+},C;H,J)=\widehat{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)/{\mathbb{R}}.$
Using the constructions of Kuranishi structures from [20, Theorem 19.44], one can show that the compactification of $\mathcal{M}_{k}(x_{-},x_{+},C;H,J)$, denoted by $\overline{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)$, admits an oriented Kuranishi structure with corners; see [68, Section 3.2.2]. Futhermore, one has evaluation maps
$\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}:\overline{\mathcal{M}}_{k}(x_{-},x_{+},C;H,J)\longrightarrow M$
given by evluating at the $i$-th marked point, i.e. $\operatorname{ev}_{i}([u,z_{1},\ldots,z_{k}])=u(z_{i})$ for all $u,z_{1},\ldots,z_{k}]$ in the dense subset $\mathcal{M}_{k}(x_{-},x_{+},C;H,J)$ and $i\in\{1,\ldots,k\}$. Given an ordered set $I=(i_{1},\ldots,i_{k})$ with each $i_{j}\in\{1,\ldots,m\}$, we consider the fiber product
$\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I}):=\overline{\mathcal{M}}(x_{-},x_{+},C;H,J)_{\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}}\times_{f_{i_{1}},\ldots,f_{i_{k}}}(\Delta_{i_{1}}\times\cdots\times\Delta_{i_{k}}).$
It follows from [21, Appendix A1] or [23, Section 3] that there is an oriented Kuranishi structure with corners on $\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})$ which has the (virtual) dimension
$\dim\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})=\mu(C)-\sum^{k}_{j=1}(2n-2d(i_{j})-2)-1.$
When $\mu(C)=\delta(I)+1$, we let $\mathsf{s}_{C,I}$ denote the multisection associated to Kuranishi structure on $\widetilde{\mathcal{M}}(x_{-},x_{+},C;\Delta_{I})$, and denote $|\mathsf{s}_{C,I}^{-1}(0)|$ the sum of the rational multiplicities of the points of the zero-dimensional vanishing locus of $\mathsf{s}_{C,I}$.
Following [68, Proposition 3.5] (see also [22, Chapter 6]), we will deform the standard Floer differential by a general homology class $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$. For any $x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\in\mathcal{O}(H)$, we denote
$\mathsf{n}_{(H,J);k}\big([x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\big)=\sum_{\begin{subarray}{c}C\in\pi_{2}(x_{-},x_{+})\\ [x_{+},\bar{x}_{+}]=[x_{+},\bar{x}_{-}\sharp C]\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k}\\ \mu(C)=\delta(I)+1\end{subarray}}\frac{1}{k!}|\mathsf{s}_{C,I}^{-1}(0)|\exp\big(\int_{C}\theta\big)a_{I}.$
Here $\bar{x}_{-}\sharp C$ stands for the concatenation of $\bar{x}$ and any representative $u$ of the class $C$. Clearly, the equivalence class $x_{+},\bar{x}_{-}\sharp C]$ does not depend on the choice of $u$. Now we define the map $\partial_{\zeta}^{H,J}:CF_{*}(H)\to CF_{*}(H)$ by extending $\Lambda$-linearly from
$\partial_{\zeta}^{H,J}[x_{-}\bar{x}_{-}]=\sum_{[x_{+},\bar{x}_{+}]\in\mathcal{O}(H)}\sum_{k=0}^{\infty}\mathsf{n}_{(H,J);k}\big([x_{-},\bar{x}_{-}],[x_{+},\bar{x}_{+}]\big)[x_{+},\bar{x}_{+}].$
Then Propositions 3.5 and 3.6 in [68] imply the following.
Proposition 2.9.
For any $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, the $\partial_{\zeta}^{H,J}$ as above is a well-defined endomorphism and satisfies
$\big(\partial_{\zeta}^{H,J}\big)^{2}=0.$
With this preparation regarding the complex, now we can define the $\zeta$-deformed Floer homology of $(CF_{*}(H),\partial_{\zeta}^{H,J})$ as follows.
Definition 2.10.
The homology of the complex $(CF_{*}(H),\partial_{\zeta}^{H,J})$ is defined by
$HF_{\zeta}(H)=\frac{{\rm ker}\;\partial_{\zeta}^{H,J}}{{\rm im}\;\partial_{\zeta}^{H,J}}.$
It is shown in [68, Proposition 3.8] that for a strongly nondegenerate $H\in{\mathcal{H}}$, $HF_{\zeta}(H)$ is independent of the choice of the almost complex structure $J$. This is the reason why we suppress the $J$ from the notation of the deformed Floer homology. Note that the differential $\partial_{\zeta}^{H,J}$ preserves the $\mathbb{C}$}-space $CF_{*}^{\lambda}(H)$, the corresponding homology is denoted by $HF_{\zeta}^{\lambda}(H)$, called filtered Floer homology of $H$.
2.3.5 Pair of pants products
For any Hamiltonian $H:S^{1}\times M\to{\mathbb{R}}$, one can modify $H$ in a standard way such that the associated Hamiltonian isotopy is constant near $1]\in S^{1}$, namely, letting $H^{\prime}(t,x)=\chi^{\prime}(t)H(\chi(t),x)$ for some smooth function $\chi:[0,1]\to[0,1]$ with the property that $\chi=0$ near $t=0$ and $\chi=1$ near $t=1$. In what follows, we always assume that for some small $\epsilon>0$, $H(t,0)\equiv 0$ for $|t|<\epsilon$. Given two such Hamiltonians $H,K$, we define a new Hamiltonian $H*K:S^{1}\times M\to{\mathbb{R}}$ by
[…]
We now assume that $H,K$ and HKH*K are strongly nondegenerate Hamiltonians. Then there exists a deformed pair of pants product on chain level, denoted by
$\bar{*}_{\zeta}^{\rm PP}:CF_{*}(H)\otimes CF(K)\to CF(H*K),$
which resembles the construction of the product on Hamiltonian Floer homology in [54] (see also [42]). The precise definition of the product $\bar{*}_{\zeta}^{\rm PP}$ is not used in this paper, and hence we will not go into details about it. Here we only mention that $\bar{*}_{\zeta}^{\rm PP}$ is a chain map that restricts to a map
$\bar{*}_{\zeta}^{\rm PP}:CF^{\lambda}(H)\otimes CF^{\mu}(K)\to CF^{\lambda+\mu}(H*K)$
for all $\lambda,\mu\in{\mathbb{R}}$. On the homology level, it induces the map
$_{\zeta}^{\rm PP}:HF^{\lambda}(H)\otimes HF^{\mu}(K)\to HF^{\lambda+\mu}(H*K).$
See [68, Proposition 3.11].
2.4 PSS isomorphisms
Following [68, Proposition 3.3.2], we will define a deformed Piunikhin-Salamon-Schwarz isomorphism from the Morse complex $(CM(f;\Lambda),\partial^{f})$ to the $\zeta$-deformed Floer complexes $(CF_{*}(H),\partial_{\zeta}^{H,J})$. Isomorphisms of this type have been studied in [54, 42].
For a generic Morse-Smale pair $(f,\rho)$ and a strongly nondegenerate Hamiltonian $H$ on $M$, the deformed PSS map
$\psi^{\operatorname{PSS}}_{\zeta}:(CM(f;\Lambda),\partial^{f})\longrightarrow(CF_{*}(H),\partial_{\zeta}^{H,J})$
is defined by emumerating “spiked discs” with incidence conditions corresponding to $I=(i_{1},\ldots,i_{k})$ as follows.
Let $\chi:{\mathbb{R}}\to[0,1]$ be a smooth cutoff function satisfying $\chi(s)=0$ for $s\leq 0$, $\chi(s)=1$ for $s\geq 1$, and $\chi^{\prime}(s)\geq 0$ for all $s$. Given an $S^{1}$-dependent family of almost complex structures $J$, let $\{J_{s}\}_{s\in[0,1]}$ be a two-parameter family of almost complex structures such that $J_{0,t}=J_{0}$, $J_{1,t}=J_{t}$ and $J_{s,0}=J_{0}$ for all $s,t\in[0,1]$. Here $J_{0}$ is a time-independent almost complex structure compatible with $\omega$. We define the ${\mathbb{R}}\times S^{1}$-dependent pair $(H_{\chi},J_{\chi})$ by
$H_{\chi}(s,t)=\chi(s)H_{t},\quad J_{\chi}(s,t)=J_{\chi(s),t}.$
Given $\gamma\in{\rm Per}(H)$ and $C\in\pi_{2}(\gamma)$, we choose any disc $u$ representing $C$ and denote $\mu(C)=\mu_{H}(\gamma,u)$. Let $\mathcal{M}^{\operatorname{PSS}}(\gamma)$ denote the space of the solutions $u:{\mathbb{R}}\times S^{1}\to M$ of the Floer equation
$\partial_{s}u+J_{\chi}\big(\partial_{t}u-X_{H_{\chi}}(u)\big)=0$ (2.10)
satisfying asymptotic boundary condition $u(+\infty,t)=\gamma(t)$ for all $t$ and having finite energy
$E(u)=\int_{{\mathbb{R}}\times S^{1}}\big|\partial_{s}u\big|^{2}_{J_{\chi}}dsdt.$
Note that since $H_{\chi}\equiv 0$ and $J_{\chi}\equiv J_{0}$ for $s<0$, the equation (2.10) restricts to $J_{0}$-holomorphic curve on $(-\infty,0)\times S^{1}$ with finite energy. By the removable singularity theorem, the limit $\lim_{s\to-\infty}u(s,t)$ exists and is independent of $t$. So one can define the map $ev_{-\infty}:\mathcal{M}^{\operatorname{PSS}}(\gamma)\to M$ by
$\operatorname{ev}_{-\infty}(u)=\lim_{s\to-\infty}u(s,t).$
We now define
$\mathcal{M}^{\operatorname{PSS}}_{k}(\gamma,C)=\big\{(u,\vec{z})\in\mathcal{M}^{\operatorname{PSS}}(\gamma)\times({\mathbb{R}}\times S^{1})^{k}\;\big|\;[u]=C\big\}.$
Denote its standard compactification by $\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(\gamma,C)$. The evaluation maps $ev_{i},i\in\{1,\ldots,k\}$ at the marked points, as well as the map $ev_{-\infty}$, extend naturally to this compactification.
Given $p\in\operatorname{Crit}(f)$ and $I=(i_{1},\ldots,i_{k})\in\{1,\ldots,m\}^{k}$, let $\iota_{p}:W^{u}(p)\hookrightarrow M$ denote the inclusion of the unstable manifold of $p$. We consider the fiber product:
$\mathcal{M}^{\operatorname{PSS}}_{k}(\gamma,C,p;\Delta_{I})=\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(\gamma,C)_{\operatorname{ev}_{-\infty},\operatorname{ev}_{1},\ldots,\operatorname{ev}_{k}}\times_{\iota_{p},f_{i_{1}},\dots,f_{i_{k}}}\big(W^{u}(p)\times\Delta_{i_{i}}\times\cdots\times\Delta_{i_{k}}\big).$
This space carries a Kuranishi structure with corners, and its virtual dimension is
$\mu(C)+\operatorname{ind}_{f}(p)-n-\delta(I).$
Let $\mathsf{s}_{\operatorname{PSS}}$ denote the corresponding multisection. Then the PSS map
$\psi^{\operatorname{PSS}}_{\zeta,H}:(CM(f;\Lambda),\partial^{f})\longrightarrow(CF_{*}(H),\partial_{\zeta}^{H,J})$
is defined by extending $\Lambda$-linearly from
$\psi^{\operatorname{PSS}}_{\zeta,H}p=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}\gamma\in{\rm Per}(H),\\ C=[u]\in\pi_{2}(\gamma)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \mu(C)+\operatorname{ind}_{f}(p)=n+\delta(I)\end{subarray}}|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|\exp\big(\int u^{*}\theta\big)a_{I}[\gamma,u].$
It is shown in [68, Proposition 3.9] that the deformed PSS map $\psi^{\operatorname{PSS}}_{\zeta,H}$ is a chain map which induces a ring isomorphism (over $\Lambda$)
$\Psi^{\rm PSS}_{\zeta,H}:HM(f;\Lambda)\longrightarrow HF_{\zeta}(H).$
Moreover, this map respects the ${\mathbb{Z}}_{2}$-gradings of the respective complexes. Moreover, the map $\Psi^{\operatorname{PSS}}_{\zeta,H}$ is independent of the choice of the function $\chi$, the metric $\rho$ and the almost complex structures $\{J_{s}\}_{s\in[0,1]}$.
3 Spectral invariants
3.1 Minmax critical value selector
Fix a ground field $\mathbb{F}$}. Let $X$ be a closed smooth manifold of positive dimension. We will denote by $H_{*}(X)$ the singular homology of $X$ with coefficient field $\mathbb{F}$}. Let $f\in C^{\infty}(X,{\mathbb{R}})$ be a smooth function. For any $\lambda\in{\mathbb{R}}$, we define
$X^{\lambda}:=\{x\in X\;|\;f(x)<\lambda\}.$
To a non-zero singular homology class $a\in H_{*}(X)\setminus\{0\}$, we associate a numerical invariant defined by
$c_{LS}(a,f)=\inf\{\lambda\in{\mathbb{R}}\;|\;a\in{\rm im}(i^{\lambda}_{*})\},$
where $i^{\lambda}_{*}:H_{*}(X^{\lambda})\to H_{*}(X)$ is the map induced by the inclusion $i^{\lambda}:X^{\lambda}\to X$. The function
$c_{LS}:H_{*}(X)\setminus\{0\}\times C^{\infty}(X,{\mathbb{R}})\to{\mathbb{R}}$
is called the minmax critical value selector of $f$. The following properties are well-known from classical Ljusternik-Schnirelman theory; see, e.g., [32, 9, 12, 24].
Proposition 3.1.
The minmax critical value selector $c_{LS}$ satisfies the following properties.
  1. 1.
    $c_{LS}(a,f)$ is a critical value of $f$, and $c_{LS}(ka,f)=c_{LS}(a,f)$ for any nonzero $k\in{\mathbb{F}}$.
  2. 2.
    $c_{LS}(a,f)$ is Lipschitz continuous in $f$ with respect to the $C^{0}$-distance.
  3. 3.
    Let $[pt]$ and $[X]$ denote the point class and the fundamental class, respectively. Then
    $c_{LS}([pt],f)=\min f\leq c_{LS}(a,f)\leq\max f=c_{LS}([X],f).$
  4. 4.
    If $b\neq k[X]$ and $c_{LS}(a\cap b,f)=c_{LS}(a,f)$, then the set
    $\Sigma=\{x\in\operatorname{Crit}(f)\;|\;f(x)=c_{LS}(a,f)\}$
    is homologically non-trivial; that is, for every open neighborhood $V$ of $\Sigma$, the induced map $i_{*}:H_{k}(V)\to H_{k}(X)$ by the inclusion $i:V\hookrightarrow X$ is nonzero for some $k>0$. In particular, such a set $\Sigma$ must be infinite.
3.2 Deformed Hamiltonian spectral invariants
In this section, we follow the works of Schwarz [60] and Oh [50] to define the spectral invariants and review their properties which will be useful in this paper.
For $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$ and a strongly nondegenerate Hamiltonian $H\in{\mathcal{H}}$, as in the previous section we have the PSS isomorphism
$\Psi^{\rm PSS}_{\zeta,H}:H_{*}(M;\Lambda)\longrightarrow HF_{\zeta}(H).$
Given $a\in H_{*}(M;\Lambda)\setminus\{0\}$, we put
$\rho_{\zeta}(a,H)=\inf\big\{\lambda\;|\;\Psi^{\rm PSS}_{\zeta,H}(a)\in\operatorname{im}(i^{\lambda})\big\}$
where $i^{\lambda}:HF_{\zeta}^{\lambda}(H)\to HF_{\zeta}(H)$ is the natural map induced by the inclusion of $CF_{*}^{\lambda}(H)$ to $CF_{*}(H)$. Then $\infty<\rho_{\zeta}(a,H)<+\infty$, and for any two strongly nondegenerate Hamiltonians $H,K\in{\mathcal{H}}$, $\rho_{\zeta}(a,\cdot)$ satisfies
$\int^{1}_{0}\min\limits_{M}(K_{t}-H_{t})dt\leq\rho_{\zeta}(a,H)-\rho_{\zeta}(a,K)\leq\int^{1}_{0}\max\limits_{M}(K_{t}-H_{t})dt.$
This has been proved in [68, 22]; see also Oh [50] for the undeformed case. Recall that the Hofer norm on the space of continuous functions $H:S^{1}\times M\to{\mathbb{R}}$ is given by
$\|H\|=\int^{1}_{0}\big(\max_{M}H(t,x)-\min_{M}H(t,x)\big)dt.$
Hence, if $H,K$ are both normalized, then
$|\rho_{\zeta}(a,H)-\rho_{\zeta}(a,K)|\leq\|H-K\|.$
Consequently, one can extend $\rho_{\zeta}(a,\cdot)$ by continuity to the set of all normalized continuous functions $H:S^{1}\times M\to{\mathbb{R}}$. Note that adding a time-dependent smooth function $c:S^{1}\to{\mathbb{R}}$ to $H$ does not change the Hamiltonian flow of $H$ but only shifts the value of $\mathcal{A}_{H}$ at each critical point by the constant $\int^{1}_{0}c(t)dt$, so we get
$\rho_{\zeta}(a,H+c)=\rho_{\zeta}(a,H)+\int^{1}_{0}c(t)dt.$
for all nonzero classes $a\in H(M,\Lambda)$. Then one can further extend $\rho_{\zeta}(a,\cdot)$ to a map on $C^{0}(S^{1}\times M,{\mathbb{R}})$. The numbers $\rho_{\zeta}(a,H)$ are called deformed spectral invariants of an element $H\in C^{0}(S^{1}\times M,{\mathbb{R}})$ (if $a=0$, we put $\rho_{\zeta}(0,H)=$-$\infty$).
Proposition 3.2 ([68, Proposition 3.13] or [22, Theorem 7.8]).
The deformed spectral invariant
$\rho_{\zeta}:H_{*}(M;\Lambda)\times C^{\infty}(S^{1}\times M,{\mathbb{R}})\to[-\infty,\infty)$
has the following properties:
  • (P1)
    Normalization: $I_{\nu}(a)+\int^{1}_{0}\min H(t,\cdot)dt\leq\rho_{\zeta}(a,H)\leq I_{\nu}(a)+\int^{1}_{0}\max H(t,\cdot)dt$ for all $H\in{\mathcal{H}}$. In particular, $\rho_{\zeta}(a,0)=I_{\nu}(a)$.
  • (P2)
    Spectrality: If $H$ is strongly nondegenerate, then there is $c\in CF_{*}(H)$ such that $c]=\Psi^{\rm PSS}_{\zeta,H}(a)$ and $\rho_{\zeta}(a,H)=\ell_{H}(c)$, and hence $\rho_{\zeta}(a,H)\in\operatorname{Spec}(H)$. If $M$ is rational, for any $H\in{\mathcal{H}}$ one also has $\rho_{\zeta}(a,H)\in\operatorname{Spec}(H)$.
  • (P3)
    Homotopy invariance: $\rho_{\zeta}(a,H)=\rho_{\zeta}(a,K)$, when $\varphi_{H}=\varphi_{K}$ in the universal covering of the group of Hamiltonian diffeomorphisms $\widetilde{{\rm Ham}}(M,\omega)$, and $H,K$ are normalized.
  • (P4)
    Quantum shift: $\rho_{\zeta}(\lambda a,H)=\rho_{\zeta}(a,H)+\nu(\lambda)$ for all $\lambda\in\Lambda$.
  • (P5)
    Projective invariance: $\rho_{\zeta}(wa,H)=\rho_{\zeta}(a,H)$ for any $0\neq w\in{\mathbb{C}}$.
  • (P6)
    Triangle inequality: $\rho_{\zeta}(a*_{\zeta}b,H*K)\leq\rho_{\zeta}(a,H)+\rho_{\zeta}(a,K)$ for any $H,K\in{\mathcal{H}}$.
  • (P7)
    Valuation inequality: $\rho_{\zeta}(a+b,H)\leq\max\{\rho_{\zeta}(a,H),\rho_{\zeta}(b,H)\}$. And if $\rho_{\zeta}(a,H)\neq\rho_{\zeta}(b,H)$, then this equality holds.
  • (P8)
    Duality: For the Hamiltonian $\overline{H}$ given by $\overline{H}(t,x)=-H(t,\varphi_{H}^{t}(x))$, one has
    $\rho_{\zeta}(a,\overline{H})=-\inf\big\{\rho_{\zeta}(b,H)\;|\;b\in H_{*}(M;\Lambda),\;{\textstyle\prod}(b,a)\neq 0\big\}.$
Remark 3.3.
The properties in Proposition 3.2 follow by adapting arguments in the undeformed case, see for instance [50]. Here we only mention that the spectrality property for a strongly nondegenerate $H$ follows from [67, Corollary 1.3] straightforwardly, and then for arbitrary $H$ on a rational symplectic manifold follows from [49, Lemma 7.2], and that the second statement in the (P7) follows from Remark 2.3.
Remark 3.4.
If $(M,\omega)$ is strongly semipositive, we do not need any Kuranishi structure in the constructions of the deformed spectral invariants $\rho_{\zeta}$ and just appeal to [68, Section 3.2] without missing any property listed as above. In this case Proposition 3.2 is enough in the following content and for our applications in Section 1.1. However, if one would like to obtain Theorems 1.1-1.6 without the strongly semipositive condition constraint, then one of course will need Kuranishi structures or something similar to deal with multiple covers of spheres of negative Chern number, but at least in principle one could use constructions just at the level of Fukaya and Ono [20] rather than [21]; see [68, Section 3].
Definition 3.5.
Given $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;\Lambda_{0})$, for any $H\in{\mathcal{H}}$, we define
$\gamma_{\zeta}(H):=\rho_{\zeta}([M],H)+\rho_{\zeta}([M],\overline{H})$
and
$\gamma_{\zeta}(\varphi):=\inf\{\gamma_{\zeta}(H)\;|\;\varphi=\varphi_{H}^{1},\;H\in{\mathcal{H}}\}.$
Remark 3.6.
According to the (P1), (P3) and (P6) as above, we see that $\gamma_{\zeta}$ satisfies
$0\leq\gamma_{\zeta}(\varphi)\leq\|\varphi\|:=\inf\{\|H\|\;|\;\varphi=\varphi_{H}^{1},\;H\in{\mathcal{H}}\}.$ (3.1)
In fact, $\gamma_{\zeta}$ gives rise to a norm on $\operatorname{Ham}(M,\omega)$, called a $\zeta$-spectral norm. This can be proven by straightforwardly adapting arguments that are well-known in the $\zeta=0$ case. See [50] or [69, Remark 2.2]. Since this fact is not explicitly used in the present paper, we leave its proof to the reader.
Spectral invariants are notoriously difficult to compute in general. However, for a $C^{2}$-small autonomous Hamiltonian $H:M\to{\mathbb{R}}$, the deformed Hamiltonian spectral invariants $\rho_{\zeta}$ of $H$ can be computed in terms of the minmax critical sector $c_{LS}$ for $H$.
Lemma 3.7.
Let $H:M\to{\mathbb{R}}$ be a $C^{2}$-small smooth function (i.e., the $C^{2}$-norm $\|H\|_{C^{2}}$ is sufficiently small). Then $\rho_{\zeta}(a,H)=c_{LS}(a,H)$ for all $a\in H_{*}(M,{\mathbb{C}})\setminus\{0\}$.
Proof.
First we notice that for a closed manifold $M$ with positive dimension, the set of smooth functions on $M$ is dense in the space of Morse functions in $C^{2}$-topology. One may perturb $H$ to a $C^{2}$-small smooth Morse function in $C^{2}$-norm, and then further to an autonomous Hamiltonian whose critical points miss the images of the maps $f_{i}:\Delta_{i}\to M$ as in (2.6), i.e., a strongly nondegenerate autonomous Hamiltonian $H$. Since $c_{LS}(a,H)$ and $\rho_{\zeta}(a,H)$ are both Lipschitz continuous in $H$, it suffices to prove the lemma for a $C^{2}$-small Morse and strongly nondegenerate $H$.
Since $H$ is a $C^{2}$-small Morse function, the set ${\rm Per}(H)$ indeed consists of critical points of $H$ (cf. [62, Section 7]). Note that for each $q\in\operatorname{Crit}(H)$, the set $\pi_{2}(q)$ consisting of the homotopy classes of discs with boundary mapping to $q$ is isomorphic to $\pi_{2}(M)$. Given $A\in\pi_{2}(M)$, let $A_{q}\in\pi_{2}(q)$ be the element corresponding to $A$. According to our convention for Conley-Zehnder index of Hamiltonian periodic orbits, we have $\mu(A_{q})=n-\operatorname{ind}_{H}(q)+2\langle c_{1}(TM),A\rangle$. For $q\in\operatorname{Crit}(H)$, we denote by $[q,0]$ the class of the constant periodic orbit $q$ with constant capping. Clearly,
$\mathcal{A}_{H}([q,0])=H(q).$ (3.2)
Given $q\in\operatorname{Crit}(H)$, recall that the PSS map $\psi_{\zeta,H}^{\rm PSS}:CM(H;\Lambda)\to CF_{*}(H)$ defined as in Section 2.4 is
$\psi^{\operatorname{PSS}}_{\zeta,H}q=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{\begin{subarray}{c}p\in{\rm Crit}(H),\\ A\in\pi_{2}(M)\end{subarray}}\sum_{\begin{subarray}{c}I\in\{1,\ldots,m\}^{k},\\ \operatorname{ind}_{f}(q)-\operatorname{ind}_{f}(p)+2c_{1}(A)=\delta(I)\end{subarray}}|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|\exp\big(\int_{A}\theta\big)a_{I}[p,A_{p}].$ (3.3)
Here $\mathsf{s}_{\operatorname{PSS}}$ is the multisection of the Kuranishi structure with corners on the fiber product
$\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})=\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)_{ev_{-\infty},ev_{1},\ldots,ev_{k}}\times_{\iota_{q},f_{i_{1}},\dots,f_{i_{k}}}\big(W^{u}(q)\times\Delta_{i_{i}}\times\cdots\times\Delta_{i_{k}}\big)$
where $\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)$ is the compactification of the space of perturbed holomorphic planes with $k$ marked points representing $A$ and asymptotic to $p$.
Now we may choose a time-independent almost complex structure $J$ compatible with $\omega$ such that the pair $(H,\rho_{J})$ is a Morse-Smale pair where $\rho_{J}(\cdot,\cdot)=\omega(\cdot,J\cdot)$, see Theorem 8.1 in [62]. Then we choose the two-parameter family $\{J_{s}\}_{s\in[0,1]}$ of almost complex structures as in Section 2.4 such that $J_{s}\equiv J$. In this case, we will show that the corresponding PSS map
$\psi_{\zeta,H}^{\rm PSS}:CM(H;\Lambda)\to CF_{*}(H)$
sends $q$ to $[q,0]$.
A typical element of $\overline{\mathcal{M}}^{\operatorname{PSS}}_{k}(p,A)$ consists of a disc component and possibly some other components, each of which is cylindrical solution to the Floer equation for $H$ or a $J$-holomorphic sphere.
The space $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})$ has an extra $S^{1}$-symmetry by the $S^{1}$ action of disc, cylinder components of the domain ${\mathbb{R}}\times S^{1}$, and by rotating the spherical components. Following Fukaya and Ono [20, Mainlemma 22.4], we see that the fixed locus $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ of this action is isolated, and away from the fixed locus the action is locally free. So, after taking a quotient by this $S^{1}$ action, one can construct a Kuranishi structure of dimension 1-1 on the quotient of the complement of the fixed locus and choose a multisection on it which is transversal to 0. Then we lift this multisection to the complement $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})\setminus\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ such that the quantity $|\mathsf{s}_{\operatorname{PSS}}^{-1}(0)|$ is given by contributions only coming from $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$. Note that the $S^{1}$-action on the spherical components is locally free except those corresponding to a topological trivial sphere, and that each element in $\mathcal{M}^{\operatorname{PSS}}_{k}(p,A,q;\Delta_{I})^{S^{1}}$ is independent of the $S^{1}$-coordinate, and hence, a solution $\gamma:[0,\infty)\to M$ of
$\dot{\gamma}(s)=-\chi(s)\;{\rm grad}_{\rho_{J}}H(\gamma(s))\quad\hbox{with}\;\gamma(0)\in W^{u}(q;H)$ (3.4)
(corresponding to the disc component), and possibly some number of negative gradient flowlines of $H$ (corresponding to the cylindrical components), or constant maps $u:S^{2}\to M$ with image $\gamma(0)$, where all subject to suitable incidence conditions. This implies that $A=0$ since the nonspherical components are $S^{1}$-independent, and that there will be no cylindrical components since their images in $M$ are $1$-dimensional objects which do not satisfy nontrivial incidence conditions. So the only contributions to $\psi_{\zeta,H}^{\rm PSS}(q)$ come from the solutions $\gamma$ of (3.4) with $\gamma(\infty)=p\in\operatorname{Crit}(H)$ with incidence conditions, corresponding to the points $\gamma(0)\in W^{u}(q)\cap W^{s}(p)$ which meet the cycles $\beta_{1}^{I},\dots,\beta_{k}^{I}$. Furthermore, we may ask that these later intersections are mutually transverse to each other and to the $W^{u}(q)$ and $W^{s}(p)$, then a dimension counting yields
$\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=\dim\big(W^{u}(q)\cap W^{s}(p)\big)=\sum_{j=1}^{k}(2n-2d(i_{j})).$
Consequently, from the fact that the only terms in (3.3) corresponding to $A=0$ satisfy
$\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=\delta(I)=\sum_{j=1}^{k}(2n-2d(i_{j}))-2k,$
we deduce that the only contributions to $\psi_{\zeta,H}^{\rm PSS}(q)$ have $k=0$. So $\operatorname{ind}_{H}(q)-\operatorname{ind}_{H}(p)=0$, and hence (3.3) simplifies to
$\psi_{\zeta,H}^{\rm PSS}(q)=\sum_{\begin{subarray}{c}p\in\operatorname{Crit}(H),\\ \operatorname{ind}_{H}(q)=\operatorname{ind}_{H}(p)\end{subarray}}n(q,p)\cdot[p,0]$
where $n(q,p)$ is the number of points in $W^{u}(q)\cap W^{s}(p)$. If $p\neq q$, we must have $n(p,q)=0$, otherwise, there is a $\mathbb{R}$}-action on $W^{u}(q)\cap W^{s}(p)$ which contradicts to $\dim(W^{u}(q)\cap W^{s}(p))=0$. If $p=q$, then we have $n(q,p)=1$ obviously. Therefore, $\psi_{\zeta,H}^{\rm PSS}$ sends each critical point $q$ of $H$ to $[q,0]$. This, together with (3.2), yields the desired equality. ∎
3.3 Hamiltonian Ljusternik–Schnirelman inequality
As we mentioned before, as a vector space over $\mathbb{C}$} the big quantum homology $H_{*}(M;\Lambda)$ is isomorphic to $H_{*}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda$. We set
$\widehat{H}_{*}(M;\Lambda)=H_{*<2n}(M;{\mathbb{C}})\otimes_{\mathbb{C}}\Lambda.$
Theorem 3.1.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold. Let $H\in{\mathcal{H}}$, and let $\alpha,\beta\in H_{*}(M;\Lambda)$ with $\alpha\neq 0$. Then we have
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)\leq\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta).$
Furthermore, if $\beta\in\widehat{H}_{*}(M,\Lambda)$ is non-zero, then either
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)<\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta)$
or the set ${\rm Fix}(\varphi_{H}^{1})$ is homologically non-trivial in $M$ (and hence infinite).
Next we postpone the proof of Theorem 3.1 to the end of this section, and employ this theorem to give the following corollary which will be used in Section 5.
Corollary 3.8.
Let $(M^{2n},\omega)$ be a closed rational symplectic manifold. Suppose that there exist $k$ homology classes $u_{i}\in\widehat{H}_{*}(M;\Lambda)$, $i=1,\ldots,k$ and $\beta\in H_{*}(M;\Lambda)$ such that $\beta*_{\zeta}u_{1}*_{\zeta}\cdots*_{\zeta}u_{k}\neq 0$ and $I_{\nu}(u_{i})\leq 0$. Let $H\in{\mathcal{H}}$ such that the ${\rm Fix}(\varphi_{H}^{1})$ is a finite set. Then the action functional $\mathcal{A}_{H}$ has $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ such that
$\rho_{\zeta}(\alpha_{k},H)=\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0})$
where $\alpha_{0}=\beta$ and $\alpha_{i}=\alpha_{i-1}*_{\zeta}u_{i}$ for $i\in\{1,\ldots,k\}$.
Proof.
The condition that $\beta*_{\zeta}u_{1}*_{\zeta}\cdots*_{\zeta}u_{k}\neq 0$ implies that each $\alpha_{i}$ is nonzero. Since ${\rm Fix}(\varphi_{H}^{1})$ is finite and $I_{\nu}(u_{i})\leq 0$ for any $i\in\{1,\ldots,k\}$, by Theorem 3.1 we obtain immediately
$\rho_{\zeta}(\alpha_{i},H)=\rho_{\zeta}(\alpha_{i-1}*_{\zeta}u_{i},H)<\rho_{\zeta}(\alpha_{i-1},H)+I_{\nu}(u_{i})\leq\rho_{\zeta}(\alpha_{i-1},H).$
So we get
$\rho_{\zeta}(\alpha_{k},H)<\ldots<\rho_{\zeta}(\alpha_{0},H).$
Combining this with Proposition 3.2 (P2) yields $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ as required. ∎
Proof of Theorem 3.1. The argument we will make to show the theorem is an adaption of Howard [33]; see also [60, 24, 8, 26]. The key idea of the proof is to use Lemma 3.7 to transform the desired inequality on the spectral invariant $\rho_{\zeta}$ to the 4th property of $c_{LS}$ in Proposition 3.1.
First, by Proposition 3.2 (P6), we see that
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*0)\leq\rho_{\zeta}(\alpha,H)+\rho_{\zeta}(\beta,0).$
This, together with by Proposition 3.2 (P1) and (P3), implies
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*0)\leq\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta)$
where we have used the fact that H0H*0 and $H$ have the same mean value, that is, $\int^{1}_{0}\int(H*0)_{t}\omega^{n}dt=\int^{1}_{0}\int H_{t}\omega^{n}dt$ and $\varphi_{H*0}^{1}=\varphi_{H}^{1}$ in $\widetilde{{\rm Ham}}(M,\omega)$.
Next, we prove the second statement. It suffices to show that if $\alpha\in H_{*}(M;\Lambda)$, $\beta\in\hat{H}(M;\Lambda)$ are two nonzero classes with the property
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)=\rho_{\zeta}(\alpha,H)+I_{\nu}(\beta),$ (3.5)
then ${\rm Fix}(\varphi_{H}^{1})$ is homologically nontrivial in $M$. For this end, we pick any open neighborhood $V$ of ${\rm Fix}(\varphi_{H}^{1})$. We may assume that $M\setminus\overline{V}$ is nonempty since nothing needs to be proved if ${\rm Fix}(\varphi_{H}^{1})=M$. We shall prove that $\overline{V}$ is homologically nontrivial.
Pick a $C^{2}$-small function $f:M\to{\mathbb{R}}$ such that $f|_{\overline{V}}\equiv 0$ and $f|_{M\setminus\overline{V}}<0$. We shall prove the following
Claim 3.9.
For sufficiently small $\varepsilon>0$, we have
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H*\varepsilon f)=\rho_{\zeta}(\alpha*_{\zeta}\beta,H).$
Proof of Claim 3.9. Consider now a family of spectral invariants $\rho_{\zeta}(\alpha*_{\zeta}\beta,H*s\varepsilon f)$ for $s\in[0,1]$. Here to guarantee that the concatenated Hamiltonians $H*s\varepsilon f$ make sense, one can assume that for some small $\epsilon>0$, $H=f\equiv 0$ for $|t|<\epsilon$ by the reparameterization as before (cf. Section 2.3.5). Since $\rho_{\zeta}(\alpha*_{\zeta}\beta,H)$ is continuous in $H$ in $C^{0}$-topology and the action spectra $\operatorname{Spec}(H*s\varepsilon f)$ have measure zero for all $s$, we only need to prove that for $\varepsilon>0$ small enough, the sets $\operatorname{Spec}(H*s\varepsilon f)$ are the same for all $s\in[0,1]$. Note that
${\rm Fix}(\varphi^{1}_{H*\varepsilon f})={\rm Fix}(\varphi_{H}^{1})$
provided that $\varepsilon$ is sufficiently small. In fact, we have $\varphi^{1}_{H*\varepsilon f}=\varphi^{1}_{H}\circ\varphi^{1}_{\varepsilon f}$ and $\varphi^{t}_{\varepsilon f}|_{V}\equiv Id$. Clearly, ${\rm Fix}(\varphi^{1}_{H*\varepsilon f})\supseteq{\rm Fix}(\varphi_{H}^{1})$. Suppose now that $p\in M$ is a fixed point of $\varphi^{1}_{H*\varepsilon f}$, and hence $\varphi^{1}_{\varepsilon f}(p)=\varphi^{-1}_{H}(p)$. Since the set $M\setminus V$ is compact and $q\neq\varphi_{H}^{-1}(q)$ for any $q\in M\setminus V$ (remind that $V$ contains ${\rm Fix}(\varphi_{H}^{1})$), one may choose $\varepsilon>0$ small enough such that $\varphi_{\varepsilon f}^{1}(q)=\varphi_{f}^{\varepsilon}(q)\neq\varphi_{H}^{-1}(q)$ for any $q\in M\setminus V$. Therefore, $p\in V$, and hence $p=\varphi_{H}^{-1}(p)$. So ${\rm Fix}(\varphi^{1}_{H*\varepsilon f})\subseteq{\rm Fix}(\varphi_{H}^{1})$.
We now prove that $\operatorname{Spec}(H*\varepsilon f)=\operatorname{Spec}(H)$ for any given small $\varepsilon>0$. For each $x\in{\rm Fix}(\varphi_{H}^{1})$, we let $x(t)=\varphi_{H}^{t}(x),t\in[0,1]$ and $\overline{x}:D^{2}\to M$ be a continuous disc such that $\overline{x}(e^{2\pi it})=x(t)$. It is easy to see that
[…]
Corresponding to each critical point $(x,\overline{x})$ of $\mathcal{A}_{H}$, we denote by $x^{\prime}(t)=\varphi_{H*\varepsilon f}^{t}(x)$ and let $\overline{x}^{\prime}:D^{2}\to M$ be a topological disc given by
[…]
Clearly, on the boundary of this disc, $\overline{x}^{\prime}$ restricts as the periodic orbit $x^{\prime}$. So we get a bijection
$\mathcal{J}:\operatorname{Crit}(\mathcal{A}_{H})\longrightarrow\operatorname{Crit}(\mathcal{A}_{H*\varepsilon f}),\quad(x,\overline{x})\mapsto(x^{\prime},\overline{x}^{\prime}).$
A straightforward computation yields $\mathcal{A}_{H*\varepsilon f}(x,\overline{x})=\mathcal{A}_{H}(\mathcal{J}(x^{\prime},\overline{x}^{\prime}))$. This will result in $\operatorname{Spec}(H*\varepsilon f)=\operatorname{Spec}(H)$. Similarly, one can show that $\operatorname{Spec}(H*s\varepsilon f)=\operatorname{Spec}(H)$ for all $s$. This completes the proof of the claim.
Now we are in a position to finish the proof. By our assumption, for $\beta\in\widehat{H}(M;\Lambda)$, we may write $\beta=\sum_{i}\beta_{i}\otimes k_{i}$ with $0\neq\beta_{i}\in H_{*<2n}(M;{\mathbb{C}})$ and $k_{i}\in\Lambda$. By Proposition 3.2 (P1), (P4) and (P7), for given sufficiently small $\varepsilon>0$, the following inequalities hold:
$\rho_{\zeta}(\beta,\varepsilon f)$ $\leq$ $\max_{i}\big\{\rho_{\zeta}(\beta_{i}\otimes k_{i},\varepsilon f)\big\}$ (3.6)
$=$ $\max_{i}\big\{\rho_{\zeta}(\beta_{i},\varepsilon f)+\nu(k_{i})\big\}$
$\leq$ $\max_{i}\big\{\rho_{\zeta}(\beta_{i},\varepsilon f)\big\}+I_{\nu}(\beta).$
Then it follows from Proposition 3.2 (P6) and Claim 3.9 yields
$\rho_{\zeta}(\alpha*_{\zeta}\beta,H)$ $=$ $\rho_{\zeta}(\alpha*_{\zeta}\beta,H*\varepsilon f)$ (3.7)
$\leq$ $\rho_{\zeta}(\alpha,H)+\rho_{\zeta}(\beta,\varepsilon f)$
$\leq$ $\rho_{\zeta}(\alpha,H)+\max_{i}\big\{\rho_{\zeta}(\beta_{i},\varepsilon f)\big\}+I_{\nu}(\beta),$
where in the last inequality we have used (3.6). On the one hand, from (3.7) we deduce that there exists $i_{0}\in{\mathbb{N}}$ such that
$\rho_{\zeta}(\beta_{i_{0}},\varepsilon f)\geq\rho_{\zeta}(\alpha*_{\zeta}\beta,H)-\rho_{\zeta}(\alpha,H)-I_{\nu}(\beta).$
Then, by (3.5) and Lemma 3.7, we get
$c_{LS}(\beta_{i_{0}},\varepsilon f)=\rho_{\zeta}(\beta_{i_{0}},\varepsilon f)\geq 0.$
On the other hand, it follows from Proposition 3.1.3 that
$c_{LS}(\beta_{i_{0}},\varepsilon f)=c_{LS}([M]\cap\beta_{i_{0}},\varepsilon f)\leq c_{LS}([M],\varepsilon f)=\max\varepsilon f=0.$
Consequently, we obtain
$c_{LS}([M]\cap\beta_{i_{0}},\varepsilon f)=c_{LS}([M],\varepsilon f)=0$
for some nonzero class $\beta_{i_{0}}\in H_{*<2n}(M;{\mathbb{C}})$. Then by Proposition 3.1.4, the zero level set $\overline{V}$ of $f$ is homologically nontrivial. This completes the proof of the second statement. ∎
4 Nonvanishing Gromov-Witten invariants
We begin with the following lemma whose proof is essentially due to Usher [68, Lemma 4.2] (see also [45, Lemma A.3] or [43, Lemma 7.1]).
Lemma 4.1.
Let $(M^{2n},\omega)$ be a closed symplectic manifold which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],b_{1},\ldots,b_{k})$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, $a_{1},\ldots,a_{l}\in H_{*<2n}(M;{\mathbb{Q}})$ with $2\leq l\in{\mathbb{N}}$, and the classes $b_{1},\ldots,b_{k}$ are rational homology classes of even degree. Then there is an open set of possible choice of deformation parameter $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that the $l$-fold product $a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}=\lambda[M]+\beta$ with $\beta\in H_{*<2n}(M;\Lambda_{0})$ satisfies $\nu(\lambda)\geq-\langle[\omega],A\rangle$.
Proof.
First we show that the class $A$ appearing in our nonzero Gromov-Witten invariant is nonzero. According to the fundamental class axiom of mixed Gromov-Witten invariants (see [47, Proposition 7.5.7]), none of the classes $b_{1},\ldots,b_{k}$ can be a multiple of the fundamental class. If $A=0$, then the dimension formula (2.2) yields
$\sum_{i=1}^{l}(2n-\deg(a_{i}))+2n+\sum_{j=1}^{k}(2n-2-\deg(b_{j}))=2n$
contradicting our assumption that $2n-\deg(a_{i})\geq 1$ for all $i$.
By the divisor axiom of the mixed Gromov-Witten invariants, one may assume that $b_{1},\ldots,b_{k}\in\oplus_{i=0}^{n-2}H_{2i}(M;{\mathbb{Q}})$. Let $\xi_{1},\ldots,\xi_{N}$ be a homogeneous basis for the $\mathbb{Z}$}-module $\oplus_{i=0}^{n-1}H_{2i}(M;\mathbb{Z})/{\rm torsion}$. We may order them so that $\xi_{1},\ldots,\xi_{s}$ form a basis for $H_{2n-2}(M;\mathbb{Z})/{\rm torsion}$ for some $s<N$. Then by the multilinearity and symmetry properties of the Gromov-Witten invariants, we may further assume that our nonzero Gromov-Witten invariant is of the form
${\rm GW}_{0,l+1+\sum_{j}\alpha_{j},A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],\underbrace{\xi_{s+1},\ldots,\xi_{s+1}}_{\alpha_{s+1}},\ldots,\underbrace{\xi_{N},\ldots,\xi_{N}}_{\alpha_{N}})$ (4.1)
for some $\alpha=(\alpha_{s+1},\ldots,\alpha_{N})\in{\mathbb{N}}^{N-s}$ with $\sum_{j=s+1}^{N}\alpha_{j}\leq k$.
For $\vec{x}=(x_{1},\ldots,x_{s})\in{\mathbb{C}}^{s}$ and $\vec{y}=(y_{s+1},\ldots,y_{N})\in{\mathbb{C}}^{N-s}$, we put
$\zeta(x,y)=\sum_{i=1}^{s}x_{i}\xi_{i}+\sum_{j=s+1}^{N}y_{j}\xi_{j}.$
Let us consider the class $a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}$ as a function of $(\vec{x},\vec{y})$. Take a homogeneous basis $\{c_{j}\}_{j=1}^{K}$ for $H_{*}(M;{\mathbb{Q}})$ with dual basis $\{c^{j}\}_{j=1}^{K}$ with respect to the Poincaré homology intersection pairing $\circ$. According to the symmetry properties of the Gromov-Witten invariants, after expanding out the formula for $a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}$, we see that the coefficient on the class $[M]$ is an expression of the form
$\sum_{g\in\Gamma_{\omega}}\sum_{\beta=(\beta_{s+1},\ldots,\beta_{N})}f_{g,\beta}(\vec{x})y_{s+1}^{\beta_{s+1}}\cdots y_{N}^{\beta_{N}}T^{g}$
where $f_{g,\beta}(\vec{x})$ is a function of $\vec{x}$ which is identically zero for all but finitely many $\beta$ (by the dimension formula (2.2)). In particular, for the symplectic period $g=\langle[\omega],A\rangle$ and the value $\beta=\alpha$, we obtain
$\bigg($ $\prod\limits_{i=s+1}^{N}$ $\big(\alpha_{i}!\big)\bigg)f_{\langle[\omega],A\rangle,\alpha}(\vec{x})$
$=$ $\sum_{\stackrel{{\scriptstyle B\in H_{2}(M;\mathbb{Z})/{\rm torsion},}}{{\langle[\omega],B\rangle=\langle[\omega],A\rangle}}}\bigg({\rm GW}_{0,l+1+\sum\alpha_{j},B}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt],\underbrace{\xi_{s+1},\ldots\xi_{s+1}}_{\alpha_{s+1}},\ldots,\underbrace{\xi_{N},\ldots,\xi_{N}}_{\alpha_{N}})$
$\hskip 284.52756pt\times\bigg(\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap B}\bigg)\bigg).$
We notice that the coefficient on $\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap A}$ in the above sum is nonzero because of our assumed nonzero Gromov-Witten (4.1). Now $H_{2n-2}(M;\mathbb{Z})/{\rm torsion}$ has $\{\xi_{i}\}_{i=1}^{s}$ as a basis, the map $B\mapsto(\xi_{1}\circ B,\ldots,\xi_{s}\circ B)$ is injective. Consequently, in the expansion of $a_{1}*_{\zeta},\ldots,*_{\zeta}a_{l}$, the coefficient of $T^{\langle[\omega],A\rangle}[M]$ is a polynomial in $e^{x_{i}}$ and $y_{j}$ having a nonzero coefficient multiplying
$\bigg(\prod_{i=1}^{s}(e^{x_{i}})^{\xi_{i}\cap A}\bigg)\prod_{j=s+1}^{N}y_{j}^{\alpha_{j}}.$
Then it follows from the fact that this polynomial does not vanish identically that there is an open dense set of choices of $(\vec{x},\vec{y})$ at which the coefficient on $T^{\langle[\omega],A\rangle}[M]$ is not zero. This proves that for such choice of $(\vec{x},\vec{y})$, writing $a_{1}*_{\zeta(\vec{x},\vec{y})},\ldots,*_{\zeta(\vec{x},\vec{y})}a_{l}=\lambda[M]+\beta$ with $\beta\in H_{*<2n}(M;\Lambda_{0})$, we have $\nu(\lambda)\geq-\langle[\omega],A\rangle$. ∎
Before finishing this section, we notice that the argument used in the proof of Lemma 4.1 also yields the following
Lemma 4.2.
Let $(M^{2n},\omega)$ be a closed symplectic manifold which admits a nonzero mixed Gromov-Witten invariant of the form:
${\rm GW}_{0,l+k+1,A}^{\{1,\ldots,l+1\}}(a,\ldots,a,b_{0},\ldots,b_{k})$
where $A\in H_{2}(M;\mathbb{Z})/{\rm torsion}$, the class $a\in H_{*<2n}(M;{\mathbb{Q}})$ occurs $l$ times for some $l>2n$, and $b_{0},\ldots,b_{k}$ (with $k\geq 0$) are rational homology classes of even degree with $b_{0}\in H_{*<2n}(M;{\mathbb{Q}})$. Then there is an open dense set of possible choices of the deformation parameter $\zeta\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that $(\underbrace{a*_{\zeta},\ldots,*_{\zeta}a}_{l})_{\langle[\omega],A\rangle}\neq 0$, and in particular the $l$-fold product of $a$ w.r.t. $_{\zeta}$ is nonzero.
Sketch of the proof of Lemma 4.2. First note that the class $A$ appearing in the mixed Gromov-Witten invariant cannot be zero. Indeed, the fundamental class axiom of Gromov-Witten invariants (cf. [47, Proposition 7.5.7]) implies that none of the classes $b_{1},\ldots,b_{k}$ can be $\lambda[M]$ for some $\lambda\in{\mathbb{Q}}\setminus\{0\}$. Assume now that $A=0$ in $H_{2}(M;\mathbb{Z})/{\rm torsion}$. Using the dimension formula (2.2), one finds that
$l(2n-\deg(a))+2n-\deg(b_{0})+\sum_{j=1}^{k}(2n-2-\deg(b_{j}))=2n,$
contradicting our assumption that $l>2n$ and $2n-\deg(a)\geq 1$. So $A\neq 0$ in $H_{2}(M;\mathbb{Z})/{\rm torsion}$. The remaining argument of the proof parallels that of Lemma 4.1 and is therefore omitted.∎
5 Proofs of Theorems 1.1-1.5 and Corollary 1.2
Proof of Theorem 1.1. Let $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$. Without loss of generality, we may assume that the number of elements in ${\rm Fix}(\varphi_{H}^{1})$ is finite, otherwise, nothing needs to be proved. The proof of the theorem is divided into two parts.
(1) Since our symplectic manifold $(M,\omega)$ carries the nonzero mixed Gromov-Witten invariant, by Lemma 4.1, there is a class $\zeta\in\oplus_{i=0}^{n-1}H(M;{\mathbb{C}})$ such that
$a_{1}*_{\zeta}\cdots*_{\zeta}a_{l}=\lambda[M]+\beta$ (5.1)
with $\beta\in H_{*<2n}(M;\Lambda_{0})$ and $\nu(\lambda)\geq-\langle[\omega],A\rangle$. Set $\alpha_{0}=[M]$ and
$\alpha_{j}=a_{1}*_{\zeta}\cdots*_{\zeta}a_{j},\quad j=1,\ldots,l.$
It follows from Corollary 3.8 that there exist elements $\tilde{x}_{0}=(x_{0},\overline{x}_{0}),\ldots,\tilde{x}_{l}=(x_{l},\overline{x}_{l})\in\operatorname{Crit}(\mathcal{A}_{H})$ such that
$\rho_{\zeta}(\alpha_{l},H)=\mathcal{A}_{H}(x_{l},\overline{x}_{l})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$ (5.2)
According to Theorem 3.1, we have
$\rho_{\zeta}(\beta,H)=\rho_{\zeta}([M]*_{\zeta}\beta,H)<\rho_{\zeta}([M],H)+I_{\nu}(\beta)$ (5.3)
By our hypothesis, the class $A$ minimizes $\langle[\omega],A\rangle$ among all $B\in H_{2}(M)$ with a nonzero mixed Gromov-Witten invariant of the form
${\rm GW}_{0,l+1+k,B}^{1,\dots,l+1}(a_{1},\ldots,a_{l},c,b_{1},\ldots,b_{k})$
for any $c\in H_{*}(M;{\mathbb{Q}})$ and even-degree classes $b_{j}$. By the definition of $l$-fold products (cf. (2.3)), this implies
$I_{\nu}(\beta)\leq I_{\nu}(\lambda[M])=\nu(\lambda)$
which, combining with (5.3), yields
$\rho_{\zeta}(\beta,H)<\rho_{\zeta}(\lambda[M],H).$
Hence, by Proposition 3.2 (P7), we get
$\rho_{\zeta}(\alpha_{l},H)=\rho_{\zeta}(\lambda[M]+\beta,H)\geq\rho_{\zeta}([M],H)+\nu(\lambda)\geq\rho_{\zeta}([M],H)-\langle[\omega],A\rangle,$
equivalently,
$\rho_{\zeta}(\alpha_{0},H)-\rho_{\zeta}(\alpha_{l},H)\leq\langle[\omega],A\rangle.$ (5.4)
(2) For estimating ${\rm Fix}(\varphi^{1}_{H})$, we set
$\mathcal{E}=\big\{\mathcal{A}_{H}(\widetilde{x}_{i})|i=0,\ldots,l\big\},$
and define an equivalent relation on $\mathcal{E}$ as follows: for $\in\mathcal{E}, aba\sim b$ if and only if aba-b is a symplectic period, i.e., $a-b\in\Gamma_{\omega}$. Denote by $\widehat{\mathcal{E}}$ the set of such equivalence classes. Hence, the cardinality $\widehat{\mathcal{E}}$ gives a lower bound for the number of distinct Hamiltonian periodic orbits $x_{i}$, and hence for ${\rm Fix}(\varphi_{H}^{1})$.
Let $k=\langle[\omega],A\rangle/p$ (recall that $p\in\Gamma_{\omega}$ is the positive generator of $\Gamma_{\omega}$). According to (5.2) and (5.4), there is at most one equivalence class containing $k+1$ elements in $\mathcal{E}$, and each remaining class contains at most $k$ elements. Therefore,
${\rm Fix}(\varphi^{1}_{H})\geq\#\widehat{\mathcal{E}}\geq 1+\bigg\lceil\frac{l+1-k-1}{k}\bigg\rceil=\bigg\lceil\frac{pl}{\langle[\omega],A\rangle}\bigg\rceil.$
This completes the proof of the estimate (1.2). ∎
Proof of Corollary 1.2. If $(M,\omega)$ has a nonzero Gromov-Witten invariant of the form
${\rm GW}_{0,l+1,A}^{\{1,\ldots,l+1\}}(a_{1},\ldots,a_{l},[pt]),$
then it follows from two facts that $\beta\in K^{\rm eff}(M,\omega)$. First, when $\sharp I=3$, the mixed Gromov-Witten invariant ${\rm GW}_{0,k,\beta}^{I}(\beta_{1},\ldots,\beta_{k})$ coincides with the usual Gromov-Witten invariant ${\rm GW}_{0,k,\beta}(\beta_{1},\ldots,\beta_{k})$. Second, the splitting property of mixed Gromov-Witten invariants (cf. [47, Theorem 7.5.10]) applies. Since the class $A$ has the minimal symplectic area among $\beta$ in $K^{\rm eff}(M,\omega)$, the conditions of Theorem 1.1 are met. So we have
${\rm Fix}(\varphi^{1}_{H})\geq\bigg\lceil\frac{pl}{\langle[\omega],A\rangle}\bigg\rceil.$ (5.5)
We now estimate the right hand side of (5.5). The nonvanishing of our Gromov-Witten invariant implies
$\sum_{i=1}^{l}(2n-{\rm deg}(a_{i}))+2n=2n+2\langle c_{1}(TM),A\rangle$
from which we deduce that
$l\cdot\min_{i}\{2n-{\rm deg}(a_{i})\}\leq 2\langle c_{1}(TM),A\rangle\leq l\cdot\max_{i}\{2n-{\rm deg}(a_{i})\}.$ (5.6)
Hence, $\langle c_{1}(TM),A\rangle>0$ because $2n-{\rm deg}(a_{i})\geq 1$ for all $i$ (so the right hand of the inequality (1.3) is positive). Combining (5.5) and (5.6) yields the estimate (1.3). ∎
Proof of Theorem 1.2. Given $\varphi\in\operatorname{Ham}(M,\omega)$, take $H\in{\mathcal{H}}$ such that $\varphi=\varphi^{1}_{H}$. Without loss of generality, we may assume that the number of elements in ${\rm Fix}(\varphi_{H}^{1})$ is finite, otherwise, nothing needs to be proved.
The proof proceeds in two steps.
Step 1. By the definition of cuplength (cf. Section 1) and the Poincaré duality, there exist $u_{i}\in H_{<2n}(M,{\mathbb{C}})$, $i=1\ldots,k$ such that $u_{1}\cap\cdots\cap u_{k}=[pt]$ and $\operatorname{cuplength}(M)=k+1$. Take classes $\alpha_{0},\dots,\alpha_{k}$ in $H_{*}(M;\Lambda)$ given by
$\alpha_{0}=[M]\quad\hbox{and}\quad\alpha_{i}=\alpha_{i-1}*_{\zeta}u_{i},\;i=1\ldots,k.$
Then by Theorem 3.1 we have
$\rho_{\zeta}(\alpha_{i},H)<\rho_{\zeta}(\alpha_{i-1},H)$
for each $i\in\{1,\ldots,k\}$. Note that the big quantum product is a quantum deformation of the homological intersection product in $H_{*}(M;{\mathbb{C}})$. In particular, we have
$\alpha_{k}=[pt]+\hbox{h.o.t.}$
where h.o.t. stands for higher-order terms coming from deformations by $J$-holomorphic spheres. This implies that each $\alpha_{i}$ is nonzero, and that ${\textstyle\prod}(\alpha_{k},[M])\neq 0$. Thus, it follows from Propositions 3.2(P8) that
$\rho_{\zeta}(\alpha_{k},H)\geq-\rho_{\zeta}([M],\overline{H}).$
Consequently, we get
$\rho_{\zeta}(\alpha_{k},H)-\rho_{\zeta}(\alpha_{0},H)\leq\gamma_{\zeta}(H).$ (5.7)
Step 2. The next goal is to prove that
${\rm Fix}(\varphi)\geq\frac{{\rm cuplength}(M;{\mathbb{C}})}{\lfloor\gamma_{\zeta}(H)/p\big\rfloor+1}.$ (5.8)
According to Corollary 3.8, there exist $k+1$ elements $\widetilde{x}_{i}=(x_{i},\overline{x}_{i})\in\operatorname{Crit}(\mathcal{A}_{H})$ such that
$\rho_{\zeta}(\alpha_{k},H)=\mathcal{A}_{H}(\widetilde{x}_{k})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(\widetilde{x}_{0}).$ (5.9)
As in the second part of the proof of Theorem 1.1, we let $\mathcal{E}=\{\mathcal{A}_{H}(x_{i},\overline{x}_{i})\;|\;i=0,\ldots,k\}$, and consider the set $\widehat{\mathcal{E}}$ consisting of equivalence classes. It follows from (5.7) that each equivalence class in $\widehat{\mathcal{E}}$ contains at most
$\bigg\lfloor\frac{\gamma_{\zeta}(H)}{p}\bigg\rfloor+1$
elements in $\mathcal{E}$. So we get
${\rm Fix}(\varphi_{H}^{1})\geq\#\widehat{\mathcal{E}}\geq\frac{k+1}{\big\lfloor(\gamma_{\zeta}(H)/p\big\rfloor+1}.$
Since the above inequality holds for any $H\in{\mathcal{H}}$ with $\varphi=\varphi_{H}^{1}$, we find that
${\rm Fix}(\varphi_{H}^{1})\geq\frac{k+1}{\big\lfloor(\gamma_{\zeta}(\varphi)/p\big\rfloor+1}.$
Proof of Theorem 1.3. Suppose ${{\rm qcl}_{\zeta}}(g)=k+1$ for some $g\in\Gamma_{\omega}$ and $\zeta\in H_{*}(M,\Lambda_{0})$. Let $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$, and assume ${\rm Fix}(\varphi_{H}^{1})$ is finite. Then there exist homology classes $a_{1},\ldots,a_{k}\in H_{<2n}(M;{\mathbb{C}})$ such that $(\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k})_{g}\neq 0$ with $g=\langle[\omega],A\rangle$ for some $A\in H_{*}(M;{\mathbb{Z}})/{\rm torsion}$. Hence,
$\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k}=\sum_{h\in\Gamma_{\omega}}(\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k})_{h}T^{h}\neq 0.$
Set $\beta=\alpha_{1}*_{\zeta}\cdots*_{\zeta}\alpha_{k}$. Then $I_{\nu}(\beta)\geq-g$. By Proposition 3.2 (P1), we obtain that
$I_{\nu}(\beta)+\int^{1}_{0}\min H_{t}dt\leq\rho_{\zeta}(\beta,H)\quad\hbox{and}\quad\rho_{\zeta}([M],H)\leq\int^{1}_{0}\max H_{t}dt.$ (5.10)
According to Corollary 3.8, there exist $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ such that
$\rho_{\zeta}(\beta,H)=\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\rho_{\zeta}([M],H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$
This, together with (5.10), gives
$g+\int^{1}_{0}\min H_{t}dt\leq\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\mathcal{A}_{H}(x_{0},\overline{x}_{0})\leq\int^{1}_{0}\max H_{t}dt.$ (5.11)
As in the proof of Theorem 1.1(2), let $\mathcal{E}=\{\mathcal{A}_{H}(x_{i},\overline{x}_{i})\;|\;i=0,\ldots,k\}$ and consider the set $\widehat{\mathcal{E}}$ consisting of equivalence classes under the relation $a\sim b$ iff $a-b\in\Gamma_{\omega}$. It follows from (5.11) that each equivalence class in $\widehat{\mathcal{E}}$ contains at most
$\bigg\lfloor\frac{\|H\|+g}{p}\bigg\rfloor+1$
elements in $\mathcal{E}$. So we get
${\rm Fix}(\varphi_{H}^{1})\geq\#\widehat{\mathcal{E}}\geq\bigg\lceil\frac{k+1}{\big\lfloor(\|H\|+g)/p\big\rfloor+1}\bigg\rceil\geq\bigg\lceil\frac{p(k+1)}{\|H\|+g+p}\bigg\rceil.$
Therefore, the desired estimate for symplectic fixed points in the theorem follows from the above inequality immediately. ∎
Proof of Theorem 1.4. According to Lemma 4.2, for each integer $l>2n$, there exists a class $\zeta_{l}\in\oplus_{i=0}^{n-1}H_{2i}(M;{\mathbb{C}})$ such that $(\underbrace{a*_{\zeta_{l}},\ldots,*_{\zeta_{l}}a}_{l})_{g_{l}}\neq 0$ for $g_{l}=\langle[\omega],A_{l}\rangle$, and hence ${\rm qcl}_{\zeta_{l}}(g_{l})\geq l+1$. Moreover, for each such $l$, by (4.1), we have the nonzero Gromov-Witten invariant
${\rm GW}_{0,l+1+K,A_{l}}^{\{1,\ldots,l+1\}}(\underbrace{a,\ldots,a}_{l},b^{l},\eta_{1}^{l},\ldots,\eta_{K}^{l})$ (5.12)
for some homogeneous class $b^{l}\in H_{*}(M;{\mathbb{Q}})$ of even degree, and homogeneous classes $\eta_{1}^{l}\ldots,\eta_{K}^{l}\in\oplus_{i=0}^{n-2}H_{2i}(M;{\mathbb{Q}})$ with $0\leq K=K(l)\leq k(l)$ (corresponding to $K=0$ there are no inserted classes $\eta_{i}^{l}$). Then it follows from Theorem 1.3 that for any $\varphi\in\operatorname{Ham}(M,\omega)$,
${\rm Fix}(\varphi)\geq\limsup_{l}\bigg\lceil\frac{p(l+1)}{p+g_{l}+\|\varphi\|}\bigg\rceil.$ (5.13)
We now estimate the right hand side of (5.13). Since the invariant (5.12) does not vanish, according to the dimension formula (2.2), we obtain
$l(2n-\deg(a))+2n-\deg(b^{l})+\sum_{j=1}^{K}(2n-\deg(\eta^{l}_{j}))=2n+2\langle c_{1}(TM),A_{l}\rangle+2K.$ (5.14)
Note that $2n-\deg(\eta^{l}_{j})\geq 4$ for all $l>2n$ and all $j\in\{1,\ldots,K(l)\}$. Then (5.14) yields
  • (a)
    $\langle c_{1}(TM),A_{l}\rangle\to\infty$ as $l\to\infty$,
  • (b)
    $\frac{2}{l}\langle c_{1}(TM),A_{l}\rangle\leq 2n-\deg(a)+\frac{1}{\ell}\sum_{j=1}^{K}(2n-2-\deg(\eta^{l}_{j}))-\frac{1}{l}\deg(b^{l})$ for all $l$.
If the function $l\mapsto k(l)$ is bounded, then $K(l)$ is also bounded. In this case, (b) implies
$\limsup_{l\to\infty}\frac{l}{\langle c_{1}(TM),A_{l}\rangle}\geq\frac{2}{2n-\deg(a)}.$
Since $\omega|_{\pi_{2}(M)}=\lambda c_{1}|_{\pi_{2}(M)}\neq 0$ for some $\lambda>0$, we get that $\omega(A_{l})\to\infty$ as $l\to\infty$, and $p=\lambda N$. Hence,
$\small\limsup_{l\to\infty}\frac{p(l+1)}{p+g_{l}+\|\varphi\|}\geq\limsup_{l}\frac{l}{\langle c_{1}(TM),A_{l}\rangle}\cdot\frac{\lambda N}{(p+\|\varphi\|)/\langle c_{1}(TM),A_{l}\rangle+\lambda}\geq\frac{2N}{2n-\deg(a)}$
from which, together with (5.13), we obtain the estimate as required. ∎
Proof of Theorem 1.5. Without loss of generality, we may assume that $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$ and that $\sharp{\rm Fix}(\varphi_{H}^{1})$ is finite. Set $\alpha_{0}=[M]$ and $\alpha_{i}=\alpha_{i-1}*_{\zeta}u_{i}$ for $i\in\{1,\ldots,k\}$. By Corollary 3.8, there exist $l+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ such that
$\rho_{\zeta}(\alpha_{l},H)=\mathcal{A}_{H}(x_{l},\overline{x}_{l})<\ldots<\rho_{\zeta}([M],H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$ (5.15)
According to our assumption, $\alpha_{l}=zT^{gp}[M]+\beta$ with $\beta\in H_{<2n}(M;\Lambda)$ and $I_{\nu}(\beta)\leq-gp$. It follows from Theorem 3.1 that
$\rho_{\zeta}(\beta,H)=\rho_{\zeta}([M]*_{\zeta}\beta,H)<\rho_{\zeta}([M],H)+I_{\nu}(\beta).$
Hence, using Proposition 3.2 (P4) and (P5), we see that
$\rho_{\zeta}(\beta,H)<\rho_{\zeta}([M],H)-gp=\rho_{\zeta}(zT^{gp}[M],H).$
This, together with Proposition 3.2 (P7), implies that
$\rho_{\zeta}(\alpha_{l},H)=\rho_{\zeta}(zT^{gp}[M],H)=\rho_{\zeta}([M],H)-gp.$
Consequently,
$\rho_{\zeta}([M],H)-\rho_{\zeta}(\alpha_{l},H)=gp.$ (5.16)
Now we let $\mathcal{E}=\{\mathcal{A}_{H}(x_{i},\overline{x}_{i})\;|\;i=0,\ldots,k\}$, and consider the set $\widehat{\mathcal{E}}$ consisting of equivalence classes as in the proof of Theorem 1.1. Then by (5.15) and (5.16) there is at most one equivalence class containing $g+1$ elements in $\mathcal{E}$, and each remaining class contains at most $g$ elements. So we obtain
${\rm Fix}(\varphi^{1}_{H})\geq 1+\bigg\lceil\frac{l+1-g-1}{g}\bigg\rceil=\bigg\lceil\frac{l}{g}\bigg\rceil.$
The proof is complete.∎
Proof of Theorem 1.6. Let $\varphi=\varphi_{H}^{1}$ for some $H\in{\mathcal{H}}$, and assume ${\rm Fix}(\varphi_{H}^{1})$ is finite. By the definition of cuplength and the Poincaré duality, there exist $u_{i}\in H_{<2n}(M,{\mathbb{C}})$, $i=1\ldots,k$ such that $u_{1}\cap\cdots\cap u_{k}=[pt]$ and $\operatorname{cuplength}(M)=k+1$. This implies
$\sum_{i=1}^{k}2n-{\rm deg}(u_{i})=2n.$ (5.17)
In what follows, we will just write $\rho$ instead of $\rho_{0}$ to stand for the spectral invariant associated to $\zeta=0$, and write * instead of $_{0}$ to stand for the undeformed product.
We now proceeds in two steps.
Firstly, we take classes $\alpha_{0},\dots,\alpha_{k}$ in $H_{*}(M;\Lambda)$ such that
$\alpha_{0}=[M]\quad\hbox{and}\quad\alpha_{i}=\alpha_{i-1}*u_{i},\;i=1\ldots,k.$
According to our assumption, $pt]*\beta=T^{gp}[M]$ with $\beta\in H_{<2n}(M;\Lambda_{0})$. Clearly, $I_{\nu}(\beta)\leq 0$. We will prove that either
${\rm Fix}(\varphi^{1}_{H})\geq k.$ (5.18)
or
$\rho(\alpha_{0},H)-\rho(\alpha_{k},H)<gp.$ (5.19)
We notice that
$u_{1}*\cdots*u_{k}=\sum_{A\in H_{2}(M;{\mathbb{Z}})}\sum_{j=1}^{K}{\rm GW}_{0,k+1,A}^{\{1,\ldots,k+1\}}(u_{1}\ldots,u_{k},e_{j})T^{\langle[\omega],A\rangle}e^{j}$
where $\{e_{j}\}_{j=1}^{K}$ is a homogeneous basis for $H_{*}(M;{\mathbb{C}})$ with dual basis $\{e^{j}\}_{j=1}^{K}$. If the Gromov-Witten invariant
${\rm GW}_{0,k+1,A}^{\{1,\ldots,k+1\}}(u_{1}\ldots,u_{k},e_{j})$
is nonzero, then we have the dimension equality
$\sum_{i=1}^{k}2n-{\rm deg}(u_{i})+2n-{\rm deg}(e_{j})=2n+2\langle c_{1}(TM),A\rangle.$
This, combining with (5.17), yields
$2\langle c_{1}(TM),A\rangle=2n-{\rm deg}(e_{j}).$
Since the minimal Chern number of $M$ is assumed to be not less than $n$, we must have: either (i) $A=0$ and ${\rm deg}(e_{j})=2n$, or (ii) $\langle c_{1}(TM),A\rangle=n$ and ${\rm deg}(e_{j})=0$. We now discuss in two cases:
Case (a). (ii) holds for some class $A$. Then, we obtain
${\rm GW}_{0,k+1,A}^{\{1,\ldots,k+1\}}(u_{1}\ldots,u_{k},[pt])\neq 0.$
The assumption that $(M,\omega)$ is monotone implies that the class $A$ achieves the minimal symplectic area, i.e., $p=\langle[\omega],A\rangle$. According to Theorem 1.1, we obtain (5.18).
Case (b). (ii) does not hold for any $A$, then case (i) holds. When $N\geq n+1$, it must belong to this case. Therefore,
$\alpha_{k}=u_{1}*\cdots*u_{k}=[pt].$ (5.20)
According to Theorem 3.1, we see that
$gp+\rho([M],H)=\rho([pt]*\beta)<\rho([pt],H)+I_{\nu}(\beta)\leq\rho([pt],H).$
Hence,
$\rho([M],H)-\rho([pt],H)<gp.$ (5.21)
Combining (5.20) and (5.21) yields (5.19).
Secondly, corresponding to case (b), we will prove
${\rm Fix}(\varphi^{1}_{H})\geq\bigg\lceil\frac{k+1}{g}\bigg\rceil.$ (5.22)
In fact, from (5.20) we see that
$M]*u_{1}*\cdots*u_{k}=[pt]\neq 0$
According to Corollary 3.8, there are $k+1$ critical points $(x_{0},\overline{x}_{0}),\ldots,(x_{k},\overline{x}_{k})$ of $\mathcal{A}_{H}$ such that
$\rho_{\zeta}(\alpha_{k},H)=\mathcal{A}_{H}(x_{k},\overline{x}_{k})<\ldots<\rho_{\zeta}(\alpha_{0},H)=\mathcal{A}_{H}(x_{0},\overline{x}_{0}).$ (5.23)
Let $\mathcal{E}=\{\mathcal{A}_{H}(x_{i},\overline{x}_{i})\;|\;i=0,\ldots,k\}$, and consider the set $\widehat{\mathcal{E}}$ consisting of equivalence classes as in the proof of Theorem 1.1. It follows from (5.19) and (5.23) that each equivalence class in $\widehat{\mathcal{E}}$ contains at most $g$ elements in $\mathcal{E}$. This gives rise to (5.22).
Summing up the two cases concludes the desired estimates. ∎
Appendix A Symplectic toric manifolds
We recall here a definition and some facts on symplectic toric manifolds, and refer the reader to the books [4, 47] for more details on this subject.
Let ${\mathbb{C}}^{n}$ be the standard $2n$-dimensional real vector space equipped with the standard symplectic form $\omega_{0}$, i.e., the imaginary part of $\frac{1}{2}\sum dz_{i}\wedge d\bar{z}_{i}$. Let ${\mathbb{T}}^{k}$ be a $k$-dimensional torus with Lie algebra $\mathfrak{t}$. Denote the integer lattice of $\mathfrak{t}$ by
$\Delta:=\{\xi\in{\mathfrak{t}}\;|\;\exp(\xi)=1\}$
and its dual lattice by
$\Delta^{*}:=\{w\in\mathfrak{t}^{*}|\langle w,\xi\rangle\in{\mathbb{Z}}\;\;\hbox{for any}\;\xi\in\Delta\}.$
Suppose that ${\mathbb{T}}^{k}$ acts diagonally on ${\mathbb{C}}^{n}$ by a homomorphism $\rho=(\rho_{1}\ldots,\rho_{n})$ with actions $\rho_{i}:{\mathbb{T}}^{k}\times{\mathbb{C}}\to{\mathbb{C}}$ given by
$(\exp(\xi),z_{i})\longmapsto\exp(-2\pi\sqrt{-1}\langle{\mathsf{w}}_{i},\xi\rangle)\cdot z_{i},\quad i=1,\ldots,n$
where ${\mathsf{w}}_{i}\in\Delta^{*}$. The moment map of this action is
$\mu(z)=\pi\sum_{i=1}^{n}|z_{i}|^{2}{\mathsf{w}}_{i}$
for any $z\in{\mathbb{C}}^{n}$. Here we require that $\mu$ is proper and the vectors ${\mathsf{w}}_{i}$ span the vector space $\mathfrak{t}$$^{*}$, and hence the linear space of solutions $(\alpha_{1},\ldots,\alpha_{n})\in{\mathbb{R}}^{n}$ for the linear equation $\sum_{i=1}^{n}\alpha_{i}{\mathsf{w}}_{i}=0$ in $\mathfrak{t}$$^{*}$ has dimension nkn-k.
Definition A.1.
A toric symplectic manifold is a symplectic quotient at a regular value $\tau\in{\mathfrak{t}}^{*}$ of the moment map $\mu$ by ${\mathbb{T}}^{k}$ such that ${\mathbb{T}}^{k}$ acts freely on $\mu^{-1}(\tau)$, namely
$M_{\tau}=\mu^{-1}(\tau)/{\mathbb{T}}^{k}$
which has a real dimension 2m=2n2k2m=2n-2k and a symplectic form $\omega_{\tau}\in\Omega^{2}(M_{\tau})$ induced by the symplectic reduction of $({\mathbb{C}}^{n},\omega_{0})$.
We denote by $H_{*}(M_{\tau})$ (resp. $H^{*}(M_{\tau})$) the quotient of the integral (co)homology group by its torsion subgroup. For $j\in I:=\{1,\ldots,n\}$, we denote by
${\rm cone}(I_{j})=\big\{\sum_{i\neq j}\alpha_{i}{\mathsf{w}}_{i}\;\big|\;\alpha_{i}\geq 0\big\}$
the positive cone spanned by the vectors ${\mathsf{w}}_{i},\;i\in I\setminus\{j\}$. Consider now the subspace
$E_{j}=\{z=(z_{1},\ldots,z_{n})\in{\mathbb{C}}^{n}\;|\;z_{j}=0\}.$
Obviously, $\tau\in{\rm cone}(I_{j})$ if and only if $E_{j}\cap\mu^{-1}(\tau)\neq 0$. In this case, one can see that the corresponding submanifold
$X_{j}=E_{j}\cap\mu^{-1}(\tau)/{\mathbb{T}}^{k}=\{[z]\in M_{\tau}\;|\;z_{j}=0\}.$
is a smooth divisor of $M_{\tau}$ and determines a codimension-$2$ class $X_{j}]$ in $H_{2m-2}(M_{\tau})$. It was shown in [28] that $X_{j}]$ ($1\leq j\leq n$) generate the homology ring of $H_{*}(M_{\tau})$ w.r.t. the usual homology intersection product $\cap$ in the following sense:
  • the homology classes $X_{1}],\ldots,[X_{n}]$ span $H_{2m-2}(M_{\tau})$,
  • every relation $\sum_{i=1}^{n}\alpha_{i}{\mathsf{w}}_{i}=0$ implies a corresponding relation $\sum_{i=1}^{n}\alpha_{i}[X_{i}]=0$,
  • for any subset $J=\{i_{1},\ldots,i_{j}\}\subset I$, the relation $\tau\notin\{\sum_{i\in I\setminus J}\alpha_{i}{\mathsf{w}}_{i}\;\big|\;\alpha_{i}\geq 0\}$ implies a corresponding relation $X_{i_{1}}]\cap\cdots\cap[X_{i_{j}}]=0$,
  • the above relations are the only ones for all classes $X_{i}]$ in $H_{*}(M_{\tau})$.
Note that $H_{2}(M_{\tau})$ can be identified with the sublattice of $\Delta\subset{\mathfrak{t}}$:
$\Delta(\tau)=\{\xi\in\Delta\;|\;\langle{\mathsf{w}}_{j},\xi\rangle=0\quad\hbox{for any}\;j\;\hbox{with}\;\tau\notin{\rm cone}(I_{j})\}.$
We denote the bijective map by
$\Delta(\tau)\longrightarrow H_{2}(M_{\tau}),\quad\eta\longmapsto A_{\eta}.$ (A.1)
Then, the dual $H^{2}(M_{\tau})$ corresponds to a quotient of the integer lattice in $\mathfrak{t}$$^{*}$. Furthermore, the first Chern class $c_{1}(TM)$ and the cohomology class of $\omega_{\tau}$ are given by
$c_{1}(TM)=\overline{{\mathsf{w}}}_{1}+\ldots+\overline{{\mathsf{w}}}_{n},\quad[\omega_{\tau}]=\overline{\tau}$
where $\overline{\gamma}$ is the image of $\gamma\in{\mathfrak{t}}^{*}$ under the correspondence
${\mathfrak{t}}^{*}=\Delta^{*}\otimes{\mathbb{R}}\longrightarrow H^{2}(M_{\tau};{\mathbb{R}}),\quad\gamma\longmapsto\overline{\gamma}.$
induced by (A.1). Let $C(\tau)\subset{\mathfrak{t}}^{*}$ be the component of $\tau$ in the set of regular values of the moment map $\mu$ which is called the Kähler cone of $M_{\tau}$. Its dual cone in $\Delta(\tau)$ is given by
$\Delta^{\rm eff}(\tau)=\{\xi\in\Delta(\tau)\;|\;\langle t,\xi\rangle\geq 0\quad\hbox{for any}\;t\in C(\tau)\}.$ (A.2)
It is well known that each nontrivial class in $H_{2}(M_{\tau})$ corresponding to an element in $\Delta^{\rm eff}(\tau)$ can be represented by a holomorphic curve in $M_{\tau}$. Such classes are called effective classes.
Definition A.2.
We say that a toric manifold $M_{\tau}$ is Fano if the vector $c_{1}=\sum_{i=1}^{n}{\mathsf{w}}_{i}$ satisfies $\langle c_{1},\xi\rangle>0$ for all $\xi\in\Delta^{\rm eff}(\tau)\setminus\{0\}$.
Let $\mathcal{D}^{\rm eff}(\tau)\subset{\mathbb{Z}}^{n}$ denote the cone
$\mathcal{D}^{\rm eff}(\tau)=\{(\langle{\mathsf{w}}_{1},\xi\rangle,\ldots,\langle{\mathsf{w}}_{n},\xi\rangle)\;|\;\xi\in\Delta^{\rm eff}(\tau)\}.$ (A.3)
Then we have a bijection $\Delta^{\rm eff}(\tau)\to\mathcal{D}^{\rm eff}(\tau)$, and denote its inverse map by
$\mathcal{D}^{\rm eff}(\tau)\longrightarrow\Delta^{\rm eff}(\tau),\quad d\longmapsto\xi_{d}.$ (A.4)
We emphasize that $\mathcal{D}^{\rm eff}(\tau)$ is not necessarily contained in the positive quadrant of ${\mathbb{Z}}^{n}$.
For $\eta=(d_{1},\ldots,d_{n})\in\Delta(\tau)$, we denote by $A_{\eta}\in H_{2}(M_{\tau})$ the corresponding homology class as in (A.1). Let $X]^{*d}$ denote the $\sum_{k}d_{k}$-fold quantum product * on $H_{*}(M_{\tau};\Lambda)$. as in (1.6). From the Batyrev-Givental formula for the quantum homology of Fano toric manifolds, it is not hard to see the following
Theorem A.1.
[47, Theorem 11.3.3] If $M_{\tau}$ is a Fano toric manifold, then the equality
$X]^{*d^{+}}=T^{\langle[\omega],A_{\xi_{d}}\rangle}[X]^{*d^{-}}$
holds in the small quantum homology $H_{*}(M_{\tau};\Lambda)$ for any $d\in\mathcal{D}^{\rm eff}(\tau)$ with the $n$-tuples $d^{+}$ and $d^{-}$ given by
$d_{i}^{+}:=\max\{d_{i},0\},\quad d_{i}^{-}:=\max\{-d_{i},0\}.$